### v1

# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

Abstract
Despite the outstanding performance of current deep learning models across various domains, their fundamental bottleneck in abstract reasoning remains unresolved. To address this challenge, the academic community has introduced Raven’s Progressive Matrices (RPM) problems as an authoritative benchmark for evaluating the abstract reasoning capabilities of deep learning algorithms, with a focus on core intelligence dimensions such as abstract reasoning, pattern recognition, and complex problem-solving. Therefore, this paper centers on solving RPM problems, aiming to contribute to enhancing the abstract reasoning abilities of machine intelligence. Firstly, this paper adopts a “causal chain modeling” perspective to systematically analyze the complete causal chain in RPM tasks: image $\rightarrow$ abstract attributes $\rightarrow$ progressive attribute patterns $\rightarrow$ pattern consistency $\rightarrow$ correct answer. Based on this analysis, the network architecture of the baseline model DIO is designed. However, experiments reveal that the optimization objective formulated for DIO, namely maximizing the variational lower bound of mutual information between the context and the correct option, fails to enable the model to genuinely acquire the predefined human reasoning logic. This is attributed to two main reasons: the tightness of the lower bound significantly impacts the effectiveness of mutual information maximization, and mutual information, as a statistical measure, does not capture the causal relationship between subjects and objects. To overcome these limitations, this paper progressively proposes three improvement methods:
  1. 1.
    Brando: By generating trainable hypothetical incorrect options, the Brando explicitly tightens the variational lower bound of mutual information.
  2. 2.
    WORLD: Abandoning the generative assumption, the WORLD constructs a Gaussian mixture model for the feature distribution of RPM images, providing an infinite and informative set of incorrect options through weighted sampling to further tighten the lower bound.
  3. 3.
    DIEGO: Introducing metadata supervision, DIEGO directly corrects semantic biases in the “abstract attributes $\rightarrow$ progressive patterns” link of the causal chain, aligning the model’s internal representations with human reasoning frameworks.
Experiments demonstrate that these three methods not only significantly enhance the accuracy of discriminative RPM tasks but also empower DIO to solve open-ended generative RPM problems. The findings of this paper offer causal-driven architectural design principles, learning objective refinement strategies, and novel insights into cross-modal generation for abstract reasoning research.
Index Terms:
Abstract Reasoning, Raven’s Progressive Matrices, Generative RPM, Variational Mutual Information, Causal Chain Modeling, Representation Learning.
I Introduction
Deep learning, a cross-disciplinary technology that draws inspiration from the operational mechanisms of the human brain, has in recent years showcased formidable potential and transformative capabilities across numerous fields, with particularly outstanding accomplishments in generative tasks. Key technologies, which include Generative Adversarial Networks (GANs)[1], Variational Autoencoders (VAEs)[2], Diffusion Model[3], and the Transformer architecture that is based on the attention mechanism[4], empower models to generate high-quality and realistic images, texts, and videos from random noise or given conditions, thereby expanding the application scope of artificial intelligence in the realm of content creation.
In computer vision, deep learning, which makes use of deep convolutional neural networks and their variants, can efficiently extract and accurately classify image features, outperforming traditional methods in tasks such as object detection and image segmentation [5, 6, 7]. In the field of natural language processing, language models that are built on deep learning can automatically learn language grammar, semantics, and contextual associations, performing exceptionally well in tasks such as machine translation and text generation, and propelling the technology towards intelligence and automation [4, 8, 9]. In the multimodal cross-field of visual question answering systems, deep learning, which integrates visual and language information, can achieve joint understanding and reasoning, accurately answering natural language questions that are related to images [11, 12].
Currently, research on Large Language Models (LLMs) is a focal point in the field of deep learning. These models, which possess a large number of parameters and are trained on massive data, demonstrate strong generalization ability in various natural language processing tasks, providing efficient and unified solutions for different tasks and facilitating the progress of artificial intelligence towards generalization and intelligence. However, the academic community has recently found that although LLMs have achieved remarkable results in language understanding and generation, they encounter obvious performance bottlenecks in abstract reasoning tasks[13, 14, 15, 16, 17]. Abstract reasoning ability, which is a vital component of human advanced intelligence and involves the understanding of abstract concepts, the inference of logical relationships, and the formulation of problem-solving strategies, serves as a key indicator for measuring the intelligence level of models. The shortcomings of LLMs not only expose the limitations of existing models in handling complex cognitive tasks but also underscore the significance of enhancing their abstract reasoning ability for achieving more powerful and general artificial intelligence.
Based on the above, this paper centers on the issue of improving the abstract reasoning ability of deep learning models. It explores effective model architecture design, training strategy optimization, and data augmentation methods, which aim to offer theoretical and practical support for breaking through the performance bottlenecks of existing models in abstract reasoning tasks and promoting the application and expansion of deep learning technology in a broader range of cognitive tasks.
II Problem Description
In artificial intelligence, abstract reasoning ability stands as a core hallmark of intelligence and a critical prerequisite for achieving human-level or superhuman AI. Yet current AI technologies are constrained by heavy reliance on large-scale annotated data, performing poorly in abstract tasks requiring causal reasoning, such as medical diagnosis and financial analysis. While academic efforts have produced datasets like RAVEN [19] and PGM [20] for benchmarking reasoning capabilities, designing algorithms to handle symbolic logic, causality, and uncertainty remains challenging. Strengthening this ability demands algorithmic innovation integrated with interdisciplinary insights from cognitive science, neuroscience, and beyond, enabling breakthroughs through cross-domain collaboration.
II-A Raven’s Progressive Matrices
Raven’s Progressive Matrices (RPM) [18] serves as an authoritative benchmark for evaluating the advanced cognitive capabilities of deep learning algorithms, focusing on core intelligence dimensions such as abstract reasoning, pattern recognition, and complex problem-solving. Its standardized testing framework demonstrates high correlation with human intelligence metrics, making it a reliable and stable tool for assessing the cognitive complexity of machine learning models. In current research, the RAVEN dataset [19] and PGM dataset [20] have emerged as the two most representative standardized test paradigms in the RPM task domain.
The RAVEN dataset comprises structured RPM instances, with each test unit containing 16 images: 8 forming a “context matrix” as the reasoning basis, and the remaining 8 constituting a candidate option pool. The task requires selecting appropriate images from the options to construct a complete $\times3$ matrix that adheres to progressive geometric transformation rules and abstract logical principles. Figure 1 (left) illustrates a typical instance from the RAVEN dataset. Similarly, while the PGM dataset also employs 8 images for context and options, its underlying logical rules must satisfy dual constraints in both row and column directions of the matrix. Figure 1 (right) presents a standard instance from the PGM dataset.
Figure 1: RAVEN and PGM Instance
II-B Related work
Image reasoning models such as CoPINet[23], LEN+teacher[22], and DCNet[24] primarily focus on learning differences and rules. Meanwhile, models like NCD[26], SAVIR-T[27], as well as neuro-symbolic systems (including PrAE, NVSA, ALANS[28, 29, 30]), have achieved promising results in both enhancing model interpretability and improving reasoning accuracy. RS-CNN and RS-TRAN[31] demonstrate exceptional capabilities in addressing problems related to RPM, while Triple-CFN[33] stands out among numerous models by implicitly extracting and indexing information about concepts and reasoning units, effectively enhancing reasoning accuracy.
III Methodology
This paper first proposes that to enhance the ability of deep learning models to solve RPM, which are causal reasoning problems, when designing the model’s structure and dividing its modules, one should fully refer to the causal chains within the problems. Based on this methodology, this paper designs a new model for solving RPM problems, named DIO. However, this paper finds that DIO is in satin without substance. Its learning approach is essentially the same as that of the previous RPM solving model, whose purpose is to maximize the mutual information between the context matrix and the correct option in RPM instance. This learning approach leads to a situation where, no matter how well its structure aligns with the causal chains in the problems, its learning and optimization will not progress towards the human predefined reasoning framework. Therefore, this paper proposes three methods, namely Brando, WORLD, and DIEGO, to refine DIO’s learning objectives. As a bonus, when DIO adopts DIEGO method, it demonstrates the capability to solve generative RPM problems.
The images in the RPM instances of this paper have adopted the annotation approach as illustrated in Figure 2, which enables clearer explanation of our model and method.
Figure 2: Annotations of Images Within an RPM Instance.
Spcifically, we denote the 16 RPM images as $\{x_{i}|i\in[1,16]\}$.
IV DIO: A Baseline Model
In this section, we design a high-performance RPM solving model, named DIO (Deep-learning Intelligent-model Organized via Causal Chains). While providing a detailed exposition of the model’s structure, we have also elucidated its underlying design principles.
IV-A The Causal Chains in RPM Problems.
The solving process of RPM problems is as follows: The solving model needs to discern the progressive pattern of abstract image attributes from the context matrices $\{x_{i}|i\in[1,8]\}$ in an RPM instance. Subsequently, it should select the ideal option $x_{\alpha}$ from the eight options $\{x_{i}|i\in[9,16]\}$ in the option pool that will not disrupt this progressive pattern [35]. This implies that there exists a causal chain in RPM problems: RPM images $\rightarrow$ abstract attributes $\rightarrow$ progressive patterns of attributes $\rightarrow$ consistency of progressive patterns $\rightarrow$ correctness of options.
This paper proposes that if the structural design of the RPM solving model is guided by this causal chain, the constructed model will demonstrate high performance. This prompted us to design the DIO model. In accordance with the causal chains in RPM problems, the DIO model is divided into four modules: the image feature extraction module, the progressive pattern induction module, the progressive pattern consistency evaluation module, and the option evaluation module.
IV-B The Image Feature Extraction Module.
RPM problems place significant emphasis on evaluating the abstract reasoning capabilities of deep models. Consequently, when extracting features from RPM images, we opt for a Visual Transformer (ViT) [21] that prioritizes image context. Figure 3 illustrates the process by which ViT extracts tokenized image features $\{z_{ij}|j\in[1,N]\}$ from the input image $x_{i}$, where $j$ denotes the token index of the image features and $N$ denotes the total patch count.
Figure 3: The Image Feature Extraction Module.
This process can be formally expressed as:
$\{z_{ij}|i\in[1,16],j\in[1,N]\}=\text{FEM}(x_{i})|_{i=1}^{16}$ (1)
IV-C The Progressive Pattern Induction Module
The deliberate design of RPM problems necessitates that participants must simultaneously observe two rows and two columns to identify the progressive patterns within the matrix. This represents a critical link in the causal chain of RPM problems. Consequently, the progressive pattern induction module of DIO adopts the structure illustrated in Figure 4.
Figure 4: The Progressive Pattern Induction Module.
Figure 4 indicates that after incorporating the features of option $\{z_{\alpha j}|j\in[1,N]\}$ into the context matrix $\{z_{ij}|i\in[1,8],j\in[1,N]\}$, DIO will summarize the progressive patterns among the features that share the same token index $j$. DIO first utilizes an MLP (Multi-Layer Perceptron) to extract row and column features from the matrix. Subsequently, it randomly selects two row features from the three available row features and two column features from the three column features, combining these four features for further processing. After binding four trainable vectors of the same dimension to the selected four features, DIO employs a Transformer Encoder to process them into eight logical outputs. Among the output vector groups, the four vectors corresponding to the trainable vectors are denoted as $\{P_{jl}|l\in[1,4]\}$.
$\{P_{jl}|l\in[1,4]\}$ is regarded as the partial representation of progressive patterns within the matrix. If all the $C_{3}^{2}\cdot C_{3}^{2}=9$ possible ways of selecting row and column features are traversed (where $C_{n}^{r}$ denotes the number of combinations of $n$ items taken $r$ at a time), then a comprehensive representation of progressive patterns, $\{P_{jkl}|k\in[1,9],l\in[1,4]\}$, can be obtained. This module can be formally expressed as:
$\{P_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$
$=\text{PPIM}(\{z_{ij},z_{\alpha j}|i\in[1,8]\})|_{j=1}^{N}$ (2)
IV-D The Progressive Pattern Consistency Evaluation Module
As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive pattern representations that share the same $j$-index and $l$-index. Specifically, DIO applies a Transformer Encoder to perform dimension-invariant processing on $\{P_{\alpha jkl}|k\in[1,9]\}$, resulting in $\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$. Subsequently, each vector in $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$ is mapped to a scalar via an MLP, serving as a logical score $s_{\alpha jkl}$. This process can be formally expressed as:
$\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$ $=\text{Trans-Encoder}(\{P_{\alpha jkl}|k\in[1,9]\})$ (3)
$s_{\alpha jkl}$ $=\text{MLP}(P^{\prime}_{\alpha jkl})$ (4)
The entire module can be represented as:
$\{s_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$
$=\text{PCEM}(\{P_{\alpha jkl}|k\in[1,9]\}_{i})|_{j=1,l=1}^{N,4}$ (5)
IV-E The Option Evaluation Module.
The DIO’s option evaluation module computes the score $S_{\alpha}$ that reflects the correctness of option $x_{\alpha}$ by summing $\{s_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$ according to the formula:
$S_{\alpha}$ $=\frac{1}{N\cdot 9\cdot 4}\big{(}\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{4}s_{\alpha jkl}\big{)}$ (6)
Use the softmax function to map $\{S_{i}|i\in[8,16]\}$ to a probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$:
$\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{S_{\alpha}}}{\sum_{{\hat{\alpha}}=9}^{16}e^{S_{{\hat{\alpha}}}}}$ (7)
We introduce a constraint mechanism that correlates the DIO-modeled probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$. This constraint enables joint optimization of all trainable parameters across DIO’s modules:
${\ell}_{\text{DIO}}=-\mathbb{E}\left[\log\frac{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\sum_{\hat{\alpha}=9,\hat{\alpha}\neq\alpha}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (8)
Where $\mathbb{E}(\cdot)$ denotes the expectation operation.
V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM Solving Model
This paper proposes that architecture optimization based on causal chains can effectively enhance the performance of deep learning networks in solving RPM problems. However, through further research, we have found that this methodology still cannot fundamentally improve the logical reasoning capabilities of deep learning networks.
V-A The learning objective of the DIO
By abstracting away the specific architecture of DIO (i.e., merging equations (1)-(6)), we obtain a unified formulation that highlights its core inductive mechanism:
$S_{\alpha}$ $=\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})$ (9)
Leveraging the probabilistic framework in Equation (7), DIO can be interpreted as a tunable probabilistic model where $\theta$ represents learnable parameters:
$\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (10)
Notably, the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ can be expressed as:
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}P(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (11)
Where $K$ denotes the number of incorrect options. If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a sufficient approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ , then substituting the $\hat{P}$ defined in Equation (10) into the Equation (12) yields a mutual information lower bound expressed by DIO:
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]$ (12)
When $K\geq 1$, that is, when there exists at least one incorrect option, the lower bound aligns with DIO’s optimization objective, as expressed by the following equation:
$\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]=-\ell_{\text{DIO}}+\log K$ (13)
Equation (13) reveals that DIO’s learning objective (i.e. maximizing $\ell_{\text{DIO}}$) is neither the human-like reasoning logic embedded in RPM problems nor even the suboptimal approach of directly maximizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather optimizing the variational lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. Furthermore, based on the properties of logarithmic functions, the following inequality holds:
${\ell}_{\text{DIO}}$ $<\mathbb{E}\left[\log\frac{\sum_{\hat{\alpha}=9}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8})}\right]$
$=-1\cdot\log\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ (14)
The right-hand side of the Inequality (V-A) embodies the classical formulation for optimizing the probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the cross-entropy (CE) loss function, which is the typical paradigm adopted by most previous end-to-end RPM solving models. From this, it can be inferred that for those end-to-end RPM solving models that have also focused on constructing the probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ in the past, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Equation (13) and Inequality (V-A) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
V-B Introducing Hypothetical Options to Derive a Tight Lower Bound for Mutual Information.
This paper argues that employing DIO to derive a tighter variational lower bound and directing its optimization objective toward the maximization of this new lower bound can more effectively maximize mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, thereby further DIO’s performance.
By analyzing Formula (12), we deduce that increasing the number of incorrect options $K$ encountered by DIO can guide its learning objective toward a tighter variational lower bound. However, RPM problem instances only offer a limited number (seven) of incorrect options. To this end, this paper attempts to introduce additional hypothetical incorrect options for DIO. Specifically, we introduce a set of hypothetical options $\{x_{i}|i\in[17,\beta]\}$ into the training pipeline of DIO. Consequently, the conditional distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ is reparameterized as:
$\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (15)
Correspondingly, the optimization objective of DIO, the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, is updated as follows:
$I\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{\beta-9}\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]$ (16)
Intuitively, the looseness of the updated lower bound has been reduced.
However, In Formula (16), it is not difficult to observe that if these hypothetical options cannot change DIO’s decision-making then this new lower bound, when serving as an optimization objective, will not undergo any fundamental improvement. In other words, when the probabilities assigned by the DIO model to the hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ approach zero, which means that these options do not alter the probability distribution $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16]\}$ that DIO assigns to the original options $\{x_{i}|i\in[9,16]\}$, the variational lower bounds derived from Equation (16) and Equation (12) become equivalent given that both function as loss functions for neural networks. This implies that unconstructive options contribute nothing to tightening the lower bound.
Therefore, this paper proposes introducing extra constructive hypothetical options into DIO—options that reinforce its endorsement of the correct options while simultaneously strengthening its repulsion of the incorrect ones—so as to steer DIO’s learning objective toward a tighter variational lower bound and ultimately further enhance its performance.
VI Brando: Tightening the Variational Lower Bound of Mutual Information by Constructing Hypothetical Options.
This paper has put forward the idea that incorporating extra constructive hypothetical options into DIO’s training process offers a promising avenue for boosting its performance. To this end, this section introduces a training framework specifically designed to furnish DIO with such hypothetical options, thereby enhancing its overall effectiveness.
VI-A Trainable Hypothetical Opions.
The RPM problem expects the training process of the solving model to involve only a limited number of supervised signals, such as the index of correct options. Relying on manual design to create constructive options goes against the original intent of RPM problems. Therefore, this paper proposes training a set of constructive options instead. This methodology is referred to as Brando (Bounded Random Options Mapping Method) in this paper.
This paper establishes a set of trainable hypothetical options $\{x_{i}|i\in[17,\beta]\}$ for the context matrix $\{x_{i}|i\in[1,8]\}$. This paper is not concerned with the specific attributes of these options; rather, it focuses solely on how much these options can alter DIO’s evaluation of the original options $\{x_{i}|i\in[9,16]\}$. In other words, we are solely concerned with the constructiveness of these hypothetical options for DIO, rather than their specific configurations.
Since these options are expected to improve $S_{\alpha}$ while reducing $\{S_{\hat{\alpha}}|{\hat{\alpha}}\in[9,16],{\hat{\alpha}}\neq\alpha\}$, this paper designs a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options and further optimize DIO:
$\ell_{\text{Brando}}=\ell_{\text{up}}+\ell_{down}$ (17)
Where $\ell_{\text{up}}$:
$\ell_{\text{up}}=-\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (18)
and $\ell_{\text{down}}$:
$\ell_{\text{down}}=-\frac{1}{\beta-17}\sum_{{\hat{\beta}}=17}^{\beta}\log\frac{e^{\text{DIO}(\{x_{i},x_{\hat{\beta}}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (19)
In these noval loss function:
  1. 1.
    It can be observed that $\ell_{\text{up}}$ requires the probability $\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ derived by DIO to be greater than all elements in the set $\{\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|{\hat{\alpha}}\in[9,\beta],{\hat{\alpha}}\neq\alpha\}$. In other words, $\ell_{\text{up}}$ demands that DIO be able to identify the correct option amidst the interference from the hypothetical options $\{x_{i}|i\in[17,\beta]\}$.
  2. 2.
    On the other hand, $\ell_{\text{down}}$ constrains that any element in the set $\{\hat{P}(x_{\hat{\beta}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\beta}\in[17,\beta]\}$ must be higher than all elements in the set $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16],\hat{\alpha}\neq\alpha\}$. In other words, $\ell_{\text{down}}$ ensures hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ outscore the original incorrect options $\{x_{i}\mid i\in[9,16],i\neq\alpha\}$ under DIO.
In loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive options. When optimizing the DIO using $\ell_{\text{Brando}}$ as the loss function, it can induce bounded hypothetical options while leveraging a tighter variational lower bound to enable DIO to maximize the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.
VI-B Map Out Hypothetical Options
Formulas (18) and (19) indicate that hypothetical options should demonstrate constructiveness for any given context matrix, meaning that hypothetical options should be customized according to the given context matrix. Therefore, training a set of shared hypothetical options is inappropriate. This implies that we require a mapping with the following form:
$\{x_{i}|i\in[17,\beta]\}=\mathcal{F}(\{x_{i}|i\in[1,8]\})$ (20)
Evidently, the mapping $\mathcal{F}(\cdot)$ is required to derive a variable number ($\beta$) and diverse hypothetical options for the context matrix $\{x_{i}|i\in[1,8]\}$. This indicates that conventional deep neural networks struggle to undertake the task of the mapping $\mathcal{F}(\cdot)$. Firstly, constructing a network with variable and dynamic output specifications is highly challenging. Additionally, from the perspective of neural network optimization, achieving a many-to-many mapping is infeasible. Therefore, this paper proposes a many-to-one mapping approach:
$x_{\hat{i}}=\hat{\mathcal{F}}(\{x_{i},o_{\hat{i}}|i\in[1,8]\})$ (21)
$\text{Where}\quad o_{\hat{i}}\sim N(\mu,\sigma^{2})$
[…]
$\{x_{\hat{i}}|i\in[17,\beta]\}=\hat{\mathcal{F}}(\{x_{i},o_{\hat{i}}|i\in[1,8]\})\mid_{\hat{i}=17}^{\beta}$ (22)
[…]
$z_{\hat{i}j}=\text{Brando}(\{z_{ij},o_{\hat{i}}|i\in[1,8]\})$ (23)
$\text{Where}\quad o_{\hat{i}}\sim N(\mu,\sigma^{2})$
Formula (23) shows that we move away from the option-mapping paradigm and instead employs a representation-mapping approach, easing the learning pressure on Brando. According to Formula (23), we have designed the structure of the Brando network as shown in the Figure 5.
Figure 5: The Brando network.
As indicated in Figure 5, the Brando network is formed by stacking multiple logical layers, with no parameter sharing between each layer. The input to each logical layer consists of $\{z_{ij}|i\in[1,8]\}$ and $o_{\hat{i}}$. Except for the first layer, the input to all subsequent layers comes from the output of the preceding layer. For the first layer, the $\{z_{ij}|i\in[1,8]\}$ is derived from the representations of the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution parameterized by mean and variance $N(\mu,\sigma^{2})$. For the output $o_{\hat{i}}$ from the final logic layer, this paper treats it as the representation $z_{\hat{i}j}$ of a hypothetical option provided by Brando for the context matrix $\{z_{ij}|i\in[1,8]\}$.
For the specific structure of each logical layer, it consists of a set of mapping matrices $\{W_{q},W_{k},W_{v}\}$, a multi-head self-attention mechanism, a feed-forward layer, a memory bank $M_{n}$, a buffer $B_{n}$, and a reparameterization head.
  1. 1.
    When the input set $\{z_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ is fed into a single logical layer, the mapping matrices $W_{q}$, $W_{k}$, and $W_{v}$ transform this input into query ($Q$), key ($K$), and value ($V$) representations, respectively.
  2. 2.
    After these query ($Q$), key ($K$), and value ($V$) representations are appended positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$.
  3. 3.
    As indicated by the computational step 1 in the diagram, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input representations stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the output set $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$. The learning process can be interpreted as:
    $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}=\frac{\phi(Q)M_{n-1}}{\phi(Q)B_{n-1}}$ (24)
    Where basis function $\phi(\cdot)$ is an element-wise ELU+1 operation.
  4. 4.
    In computation steps 2 and 3, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the key ($K$) and value ($V$) mappings derived from the current input representations. The update process can be expressed as:
    $M_{n}$ $=M_{n-1}+\phi(K)^{T}\left(V-\frac{\phi(K)M_{n-1}}{\phi(K)B_{n-1}}\right)$ (25)
    $B_{n}$ $=B_{n-1}+\sum\phi(K)$ (26)
  5. 5.
    After the learnable scalar $\tau$ is compressed into a scaling coefficient via the sigmoid function, a weighted summation is performed on $\tilde{o}_{\hat{i}}$ and $\hat{o}_{\hat{i}}$ using this coefficient to update $\hat{o}_{\hat{i}}$, enabling it to balance its own reasoning with image representations from other instances.
    $\hat{o}_{\hat{i}}\leftarrow[1-\text{Sigmoid}(\tau)]\hat{o}_{\hat{i}}+\text{Sigmoid}(\tau)\tilde{o}_{\hat{i}}$ (27)
  6. 6.
    The attention outputs $\{\hat{z}_{ij}|i\in[1,8]\}$ with the updated $\hat{o}_{\hat{i}}$, which has undergone skip connection with logical inputs, is processed by the feed-forward layer and subjected to another skip connection, yielding the logical output $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}$. This process can be expressed as:
    $Z\leftarrow\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}+\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ (28)
    $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}\leftarrow\text{FFL}(Z)+Z$ (29)
    Here, FFL()($\cdot$) refers to the feed-forward layer.
  7. 7.
    It is noteworthy that only after the $^{o}_{\hat{i}}$ within the logical output $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}$ have been processed by the reparameterization head can they serve as the final output. The process of the reparameterization head can be interpreted as follows:
[…]
    ${o}_{\hat{i}}\leftarrow{o}_{\mu}+e^{\frac{o_{\ln\sigma}}{2}}\cdot\epsilon$ (31)
    The $\epsilon$ represents a random sample drawn from $N(0,1)$ and MLP()($\cdot$) refers to a Multi-Layer Perceptron capable of doubling the dimension of $^{o}_{\hat{i}}$.
It is not hard to find that the memory bank $M_{n}$ and buffer $B_{n}$ in each logical layer will incorporate new memories as a new batch of context matrices is observed. Therefore, the initial form of the memory bank requires a clear definition.
$M_{0}=0,\,B_{0}=1\times 10^{-9}$ (32)
Furthermore, the first update of the memory bank and the buffer can be explained as follows:
$M_{1}=\phi(K)^{T}V,\,B_{1}=\sum\phi(K)$ (33)
From the learnable Gaussian distribution ${N}({\mu},{\sigma}^{2})$, we sample random vectors $^{o}_{\hat{i}}$, which undergo sequential refinement in the Brando network through dynamic interaction with both current context matrix representations and historical image representations accumulated across multi-round training iterations. This iterative conditioning process enables $^{o}_{\hat{i}}$ to progressively assimilate visual patterns observed throughout the training trajectory, ultimately yielding the ultimate hypothesis representation ${z}_{\hat{i}j}$. Moreover, alterations in the initial promoter $^{o}_{\hat{i}}$ are capable of inducing changes in the ultimate option representation $z_{\hat{i}j}$. This satisfies the demand to map diverse hypothetical options for the context matrix.
VI-C The Combination of Brando and DIO.
Prior to integrating the Brando network with DIO, DIO must be pre-trained on the original option pool $\{x_{i}|i\in[9,16]\}$ using the loss function $\ell_{\text{DIO}}$. Once DIO is pre-trained to approach its optimal performance, the Brando network will be integrated with DIO according to the method illustrated in Figure 6.
Figure 6: The Combination Method of DIO and Brando.
Subsequently, the loss function will be switched from $\ell_{\text{DIO}}$ to $\ell_{\text{Brando}}$ for continued training of this integrated system, thereby further optimizing DIO’s performance. Another key point is that during the joint training process, DIO’s parameters excluding those in the image feature extraction module will be frozen every other iteration.
VII WORLD: Tightening the Variational Lower Bound of Mutual Information Through Deconstructing the Problem Space
It is readily apparent that the architecture of the Brando network leaves substantial room for theoretical refinement. The pursuit of a more optimal structure for implementing the mapping $\mathcal{F}(\cdot)$ appears to be an open-ended endeavor, with no discernible asymptote in sight. Moreover, providing a sufficient number of hypothetical options for DIO using Brando entails considerable computational overhead. Therefore, this paper argues that Brando is not a flawless methodology. To this end, this paper proposes a method named WORLD (Weighted Option-Represented Lower-bound Delimitation Method), which breaks away from the paradigm of the mapping $\mathcal{F}(\cdot)$. This method can also tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
VII-A Modeling is Preferable to Estimation.
Formula (12) indicates that the key to guiding the learning objective of solving model towards a tighter lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$ lies in providing additional incorrect options to the model for each context matrix. Moreover, this paper also points out that the introduced additional incorrect options must avoid being unconstructive, that is, after being incorporated into the training process of the solving model, they should be able to enhance the decision-making capability of the solving model.
The Brando network essentially functions as a distribution transformation mechanism, whose task is to map samples $o_{\hat{i}}$ drawn from a Gaussian distribution $N(\mu,\sigma^{2})$ into features $z_{\hat{i}j}$ of incorrect options customized for a given context matrix. This paper argues that instead of implicitly estimating the distribution of incorrect option features for each given context matrix, it is more effective to explicitly estimate the distribution of all RPM image features $z_{ij}$. In this way, when confronted with a specific context matrix, the required incorrect option features can be sampled in a targeted manner from this estimated distribution.
However, accurately estimating the distribution of RPM image features $z_{ij}$ poses significant challenges, and even if such a distribution were successfully estimated, targeted sampling from it would remain highly non-trivial. Therefore, this paper argues that explicitly modeling the distribution of $z_{ij}$ during the process where the DIO’s feature extraction module encodes $x_{i}$ into $z_{ij}$ constitutes an effective approach, since this method eliminates the need for estimation and facilitates targeted sampling through direct distributional control.
This paper has found that the Gaussian Mixture Model (GMM) [36] offers remarkable advantages, as it not only facilitates efficient modeling but also enables targeted sampling. Suppose the features $z_{ij}$ extracted by DIO from RPM images follow a GMM with known numbers of components and parameters. Then, for a given context matrix, after identifying the components corresponding to its correct options within this GMM, we can reasonably regard the components not corresponding to the correct options as the source of incorrect option features. Based on this, we are able to sample comprehensive and diverse incorrect option features from these specific components for DIO. Given these advantages, this section proposes the WORLD method to model the distribution of $z_{ij}$ as a GMM.
VII-B Model the Distribution of RPM Image Features as a Gaussian Mixture Model.
To accurately model the distribution of RPM image features $z_{ij}$ as a GMM, this paper proposes the distribution modeling method WORLD tailored for DIO. Specifically, this paper introduces a trainable vector set $\{U_{m}|m\in[1,M]\}$ and regards it as the mean vectors of a GMM. Using this approach, we construct a predefined GMM with $M$ components. Subsequently, the image features $z_{ij}$ encoded by DIO are projected onto this predefined GMM framework, thereby modeling the distributional characteristics of $z_{ij}$.
The way in which DIO’s related modules extract features from RPM images also only requires minor modifications, so that the distribution of the extracted features $z_{ij}$ can be more easily projected onto the predefined GMM $\{U_{m}|m\in[1,M]\}$. The modifications are shown in Figure 7.
Figure 7: The modified Image Feature Extraction Module.
Specifically, doubling the overall dimension of the ViT within DIO’s image feature extraction module can achieve the desired effect, causing the dimension of the latent features encoded for RPM images to double accordingly. After doubling the dimension, the resulting latent features are evenly split into two parts, denoted as $u_{ij}$ and $b_{ij}$ respectively. Then, the sum of $u_{ij}$ and $b_{ij}$ is taken as the new feature representation $z_{ij}$ for the image $x_{i}$. This process can be expressed by the following formula:
$\{u_{ij},b_{ij}\}=\text{ViT}(x_{i}),\,z_{ij}=u_{ij}+b_{ij}$ (34)
The newly devised derivation method for $z_{ij}$ ensures that, when $b_{ij}$ follows a standard Gaussian distribution $\mathcal{N}(0,1)$, the distribution of $z_{ij}$ explicitly adheres to a Gaussian distribution $\mathcal{N}(u_{ij},1)$. Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the predefined GMM $\{U_{m}|m\in[1,M]\}$ and make the latent image features $b_{ij}$ follow the distribution $N(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM.
The aforementioned method for modeling the distribution of $z_{ij}$ can be implemented by introducing five additional loss function terms.
  1. 1.
    A loss term $\ell_{1}$ aims to encourage the mean of the distribution of $z_{ij}$ to approximate the predefined GMM $\{U_{m}|m\in[1,M]\}$, and its mathematical expression can be formulated as:
    $\ell_{1}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|u_{ij}-\text{gs}[{U}_{\hat{m}}]\|_{2}^{2}$ (35)
    $\text{where}\quad\hat{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{ij}-{U}_{m}\|_{2}^{2}$
    Here, $\text{gs}[\cdot]$ denotes the gradient stopping operation.
  2. 2.
    A loss term $\ell_{2}$ is introduced to enforce $b_{ij}$ to follow a standard normal distribution, expressed as:
    $\ell_{2}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|b_{ij}-\epsilon_{ij}]\|_{2}^{2}$ (36)
    $\text{where}\quad\epsilon_{ij}\sim N(0,1)$
  3. 3.
    A loss term $\ell_{3}$, which is utilized to train the components within the predefined GMM $\{U_{m}|m\in[1,M]\}$, can be expressed as:
    $\ell_{2}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|{U}_{\hat{m}}-\text{gs}[u_{ij}]\|_{2}^{2}$ (37)
    $\text{where}\quad\hat{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{ij}-{U}_{m}\|_{2}^{2}$
  4. 4.
    A mutual information loss term $\ell_{4}$, which facilitates the training of components within the predefined GMM $\{U_{m}|m\in[1,M]\}$, is expressed as:
    $\ell_{4}$ $=\sum_{i=1}^{16}\|x_{i}-\hat{x}_{i}\|_{2}^{2}$ (38)
    where $\hat{x}_{i}=\mathcal{D}\left(\left\{\hat{u}_{ij}+b_{ij}\mid j\in[1,N]\right\}\right),$
    $\hat{u}_{ij}=U_{\hat{m}}-\text{sg}[u_{ij}]+u_{ij},$
    $\hat{m}=\mathop{\mathrm{argmin}}_{m\in[1,M]}\|u_{ij}-U_{m}\|_{2}^{2}.$
    Obviously, the loss term $\ell_{4}$ is formulated to maximize the mutual information between the components within $\{U_{m}|m\in[1,M]\}$ and the latent features $u_{ij}$, via an auxiliary reconstruction task facilitated by the decoder $\mathcal{D}(\cdot)$. The decoder $\mathcal{D}(\cdot)$ is structured as a depth-reversed counterpart of the ViT, transforming latent features into a reconstructed image. The computation of $\hat{u}_{ij}$ ensures the existence of a gradient propagation pathway: $\hat{x}_{i}\rightarrow U_{\hat{m}}\rightarrow\hat{u}_{ij}\rightarrow u_{ij}\rightarrow x_{i}$, thereby guaranteeing the effectiveness of the loss term $\ell_{4}$.
  5. 5.
    A mode collapse prevention loss term $\ell_{5}$ for the latent features $u_{ij}$, defined as:
    $\ell_{5}$ $=\sum_{i=1}^{16}\|x_{i}-\tilde{x}_{i}\|_{2}^{2}$ (39)
    where $\tilde{x}_{i}=\mathcal{D}\left(\left\{\hat{u}_{ij}+\epsilon_{ij}\mid j\in[1,N]\right\}\right),$
    $\epsilon_{ij}\sim N(0,1)$
    $\hat{u}_{ij}=U_{\hat{m}}-\text{sg}[u_{ij}]+u_{ij},$
    $\hat{m}=\mathop{\mathrm{argmin}}_{m\in[1,M]}\|u_{ij}-U_{m}\|_{2}^{2}.$
    It is not hard to notice that during the construction of the loss term $\ell_{4}$, there exists a variable calculation pathway, which can also be referred to as an information pathway, namely: $x_{i}\rightarrow b_{ij}\rightarrow\hat{x}_{i}$. The presence of this pathway may potentially lead to mode collapse in the latent features $u_{ij}$, thereby making it impossible to model the distribution of $z_{ij}$ as a GMM. Therefore, the loss term $\ell_{5}$ comes into being, ensuring that the latent feature $u_{ij}$ cannot be bypassed during the reconstruction process of the RPM image $x_{i}$.
After performing a weighted summation of these five loss terms, we obtain the complete loss function $\ell_{\text{WORLD}}$:
$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+0.25\cdot\ell_{3}+\ell_{4}+\ell_{5}$ (40)
By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss functions, we can train a DIO model where the distribution of the latent variable $z_{ij}$ is modeled as a GMM. Notably, when training the DIO using $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, the feedforward mechanism undergoes the transformation illustrated in Figure 8.
Figure 8: Forward Feed Process of DIO Model Incorporating $\ell_{\text{WORLD}}$ as an Additional Loss Term.
In figure 8, the latent variable fed into the progressive pattern induction module within DIO is no longer $z_{ij}$, but rather $u_{ij}+\epsilon$. In other words, the forward feed process indicated by Equation (IV-C) is updated as follows:
$\{P_{jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}_{i=\alpha}$
$=\text{PPIM}(\{u_{ij}+\epsilon_{ij},u_{\alpha j}+\epsilon_{\alpha j}|i\in[1,8]\})|_{j=1}^{N}$ (41)
Where $u_{ij}+\epsilon_{ij}$ can be regarded as a sample drawn from the GMM component $U_{\hat{m}}$ that $z_{ij}$ follows. This design enables DIO to interact with a fully modeled GMM rather than deterministic image features, thereby training the DIO to make decisions based on the complete probabilistic structure of the GMM instead of specific feature instantiations.
VII-C Leveraging the Modeled GMM to Provide Incorrect Options for Context Matrices
To steer DIO’s learning objective towards a tighter variational lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, we also need to sample sufficient and diverse incorrect option features from the modeled GMM $\{U_{m}|m\in[1,M]\}$ for each given context matrix. Thanks to the inherent properties of GMM, targeted sampling within the set $\{U_{m}|m\in[1,M]\}$ is relatively straightforward. For a given context matrix $\{x_{i}|i\in[1,8]\}$, after identifying the components in the set $\{U_{m}|m\in[1,M]\}$ that correspond to the provided correct option $x_{\alpha}$, samples randomly drawn from the remaining components can be considered, to a certain extent, as features of incorrect options.
Identifying the components corresponding to the correct option, targeted sampling of incorrect option features, and steering the learning objective of DIO can all be achieved by introducing an additional loss function term:
${\ell_{\text{Weighted}}}=$ $-\sum_{j=1}^{N}\log\frac{{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{\tilde{m}}+\epsilon_{\tilde{m}}|i\in[1,8]\})}{|\tau|}}}}}{\sum_{m=1,m\neq\tilde{m}}^{M}{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{m}+\epsilon_{m}|i\in[1,8]\})}{|\tau|}}}}}$ (42)
where $\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2}$
$\epsilon_{m}\sim N(0,1)$
$\tau$ is a trainable smoothing parameter. Given that RPM problems often feature a diverse range of correct options, the rigid approach of designating samples drawn from components that do not correspond to the provided correct option as features of incorrect options carries a potential risk of overgeneralization. To mitigate this sample classification bias, we introduce a smoothing parameter $\tau$. $\tilde{\text{DIO}}$ refers to the DIO without the image feature extraction module, which is capable of processing features into scores. The calculation of $\tilde{m}$ serves to identify the component index in the modeled GMM $\{U_{m}|m\in[1,M]\}$ that corresponds to the feature $z_{\alpha j}$ of the correct option $x_{\alpha}$.
The role of $\ell_{\text{Weighted}}$ is to fine-tune the DIO model. Specifically, when DIO approaches optimal performance under the constraint of the loss function $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, incorporating the additional loss term $\ell_{\text{Weighted}}$ to update the loss function to $\ell_{\text{WORLD}}+\ell_{\text{DIO}}+\ell_{\text{Weighted}}$ creates additional learning capacity and performance improvement potential. This adjustment enables the introduction of supplementary incorrect options during DIO’s late-stage training, thereby steering its learning objective toward a tighter variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, ultimately leading to further performance enhancement.
VIII DIEGO: Methods for Rectifying the Causal Chain in DIO
This study has made two attempts to refine the learning objectives of the DIO model, resulting in the development of two distinct methodology: Brando and WORLD. In this section, we devise a methodology named DIEGO (Direct Intervention for Enforcing Guided Optimal Reasoning), which is designed to rectifying the causal chains modeled by DIO, thereby advancing its reasoning performance.
VIII-A Mutual Information Excels at Problem-Solving Rather Than Reasoning
This paper realizes that setting the maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal. While mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ directly correlates with accuracy in solving RPM problems, it fails to capture causal relationships inherent in the task design. This creates a critical discrepancy: DIO is optimized for high-accuracy RPM solution generation rather than learning the human-predefined reasoning logic. Both the Brando and WORLD methodologies aim to assist DIO in better maximizing mutual information, yet they fail to address this discrepancy.
Therefore, this paper argues that the causal chain underlying RPM problems should be further modeled. The network architecture of DIO strongly drew reference from the causal chain within RPM problems, with the hope that DIO can autonomously model the causal chain during the training process. However, this can only ensure that DIO is structurally aligned with the causal chain in RPM problems. Whether its latent variables are semantically aligned cannot be guaranteed and remains to be verified.
In order to verify semantic alignment, this paper selects the “Distribute 9” database under the RAVEN problems[19] as the research object. We conduct training on this database using DIO, as well as DIO supplemented with Brando and DIO supplemented with WORLD, respectively. Subsequently, we select 30 instances with strictly identical progressive patterns from this database and verify the pairwise similarity of the progressive pattern representations encoded by DIO for these 30 instances. The calculated $30\times 30$ similarity matrices have been plotted into heatmaps as shown in Figure 9.
Figure 9: Heatmap of Progressive Pattern Representation Extracted from the Distribute 9 under RAVEN Problems.
If DIO accurately models the causal chain embedded in RAVEN problems, then after these 30 instances with the same progressive patterns are encoded into progressive pattern representations by DIO, they should demonstrate high cosine similarity among themselves. However, the three similarity matrices in Figure 9 do not corroborate this expectation. This indicates that, even with auxiliary support from the Brando or WORLD methods, the causal chain modeled by DIO remains systematically misaligned with the ground-truth chains inherent in RPM problems. In addition, this further points out the limitations of using mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ maximization as the learning objective. Given this, this paper proposes that rectifying the causal chain modeled by DIO during its training process will substantially improve performance in solving RPM problems.
VIII-B Rectifying the Causal Chain Modeled by DIO.
In preliminary experiments, we found that DIO exhibits deviations in modeling the segment of the causal chain that involves “abstract attributes $\rightarrow$ progressive patterns of attributes”. This corresponds to a misalign between the results $P_{\alpha{jkl}}$ derived from Formula (IV-C) and human semantics. Therefore, this paper proposes that refining the progressive pattern representations $P_{\alpha{jkl}}$ encoded by DIO would be an effective solution.
This paper has taken note of the metadata provided for each instance in RPM problems, which describes the corresponding instance’s progressive pattern. We argue that incorporating this metadata as a supervisory signal into DIO’s training process to supervise the derivation of $P_{\alpha{jkl}}$ can achieve the objective of rectifying the causal chain modeled by DIO. Thus, we have designed a method named DIEGO. DIEGO involves introducing an additional set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$ for DIO, along with an auxiliary loss function term $\ell_{\text{DIEGO}}$.
Specifically, the DIEGO method first establishes a set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$, where each component represents a distinct metadata modality. Consequently, the reference vector set $\{V_{f}|f\in[1,F]\}$ can serve as a comprehensive dictionary for metadata. The loss term $\ell_{\text{DIEGO}}$ serves to align each progressive pattern representation $P_{\alpha{jkl}}$ extracted by DIO with its corresponding reference vector within $\{V_{f}|f\in[1,F]\}$ according to the accompanying metadata. The loss function $\ell_{\text{DIEGO}}$ can be expressed as follows:
${\ell_{\text{DIEGO}}}=-\sum_{\hat{f}=1}^{F}\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{2}\text{Meta}_{{\hat{f}l}}\cdot\log\frac{{{e^{({P_{\alpha jkl}}\cdot{V_{\hat{f}}})/|\tau|}}}}{\sum_{f=1}^{F}{{e^{({P_{\alpha jkl}}\cdot{V_{f}})/|\tau|}}}}$ (43)
Here, $\text{Meta}_{\hat{f}l}$ denotes the indication status of the $l$-th metadata entry with respect to the $\hat{f}-$th type of progressive pattern. Meanwhile, $\tau$ represents a trainable positive scaling coefficient. The detailed expression of the loss term $\ell_{\text{DIEGO}}$ demonstrates our methodology for achieving alignment. This method utilizes a cross-entropy loss function with a scaling coefficient to control the cosine similarities between $P_{\alpha_{jkl}}$ and all components in the metadata dictionary $\{V_{f}|f\in[1,F]\}$ according to the indication information from the metadata.
When training the DIO model using $\ell_{\text{DIO}}+\ell_{\text{DIEGO}}$ as the loss function, a solving model that accurately captures the causal chains in RPM problem will be obtained. Previously, some studies have also attempted to treat metadata as additional supervision signals and introduce extra progressive pattern learning tasks into their solving models, aiming to simultaneously enhance the model’s interpretability and reasoning accuracy. However, research led by MRNet[25, 28, 27, 31] indicates that this approach may backfire. Specifically, if metadata is directly incorporated into the training process of the solving models they designed, the model’s reasoning performance will instead decline. This may be logically flawed, but it is indeed an established fact. Nevertheless, this paper argues that the DIO model constructed based on causal chains, coupled with the DIEGO method capable of strategically correcting key segments of the constructed causal chains, holds the key to breaking through this impasse.
IX Solving Generative RPM Problems
Generative RPM problems represent an advanced variant of RPM tasks, where the solving model is required to generate an ideal answer for the context matrix $\{x_{i}|i\in[1,8]\}$ instead of selecting from a predefined bounded option pool.
IX-A DIO And WORLD Solve Generative RPM Problems
The WORLD method not only enhances the performance of DIO but also equips it with the capability to solve generative RPM problems. Specifically, the WORLD method introduces an external GMM $\{U_{m}|m\in[1,M]\}$ and a decoder $\mathcal{D}(\cdot)$ for DIO. Meanwhile, under the WORLD framework, the loss function term $\ell_{\text{Weighted}}$ defined in Formula (42) not only tightens the variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$ for DIO but also endows DIO with the ability to assign weights to each component in the GMM $\{U_{m}|m\in[1,M]\}$ with respect to the given context matrix features $\{z_{ij}|i\in[1,8],j\in[1,N]\}$. The weights assigned to the component $U_{m}$ for the contextual features $\{z_{ij}|i\in[1,8]\}$ can be calculated as follows:
$S_{{m}{j}}=\tilde{\text{DIO}}(\{z_{i{j}},U_{{m}}|i\in[1,8]\})$ (44)
Based on all computed weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$, we employ the Gumbel-Softmax trick to draw $N$ components from the GMM $\{U_{m}|m\in[1,M]\}$. This process can be formulated as:
$G_{mj}$ $=-\log(-\log(\text{U}(0,1)))$ (45)
$\tilde{S}_{mj}$ $=S_{mj}+G_{mj}$ (46)
$\hat{m}_{j}$ $=\underset{m}{\text{argmax}}\,\{\tilde{S}_{mj}|m\in[1,M]\}$ (47)
$\hat{U}_{j}$ $=U_{\hat{m}_{j}}$ (48)
These $N$ components $\{\hat{U}_{j}|j\in[1,N]\}$ can be regarded as the distribution of solutions estimated by DIO for the context matrix $\{x_{i}|i\in[1,8]\}$. The decoder $\mathcal{D}$ is capable of transforming the samples from these $N$ components into the generated answer $\hat{x}$:
$\hat{x}=\mathcal{D}(\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\})$ (49)
$\text{Where}\quad\epsilon_{ij}\sim N(0,1)$
Formulas (34), as well as formulas (44) - (49), indicate that the DIO trained by the WORLD method has the capability to estimate the distribution of solution for the context matrix $\{x_{i}\mid i\in[1,8]\}$ and generate RPM images that can serve as solutions.
IX-B DIEGO Validates the Rationality of Mutual Information Supervision
Verifying the solution accuracy of generative RPM problems has long been challenging. RPM problems often have multiple valid solutions, which means that the generative solving models cannot be evaluated simply by comparing each generated solution with the given correct option. This also makes it challenging for RPM experts to devise a perfect verification algorithm. Most prior work on generative RPM problems has used self-verification, i.e., using their own discriminative models to validate their own generative models.
However, this paper argues that this paradigm of self-verification is inadequate. The cause of the multiple-solution characteristic in RPM problems lies in the fact that solution identification is driven by a progressive pattern. Any RPM image that adheres to the progressive pattern exhibited in the contextual matrices can be considered a valid solution. This implies that algorithms or models used to verify the solution accuracy of generative RPM problems should also be progressive-pattern-driven. The DIEGO approach meets this expectation, as the DIO model trained solely with the loss term $\ell_{\text{DIEGO}}$ will be a progressive-pattern-driven solving model.
Additionally, DIO+WORLD is a strong mutual-information-based inductive model. Using the progressive-pattern-driven model DIEGO to verify DIO+WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also tests the deeper potential of mutual-information-driven solution models.
X Experiment
All experiments conducted in this paper were programmed in Python, utilizing the PyTorch[37] framework. This paper introduces a comprehensive framework for solving RPM problems through three distinct configurations that integrate a purpose-built deep learning model with specialized training methodologies. The proposed configurations leverage:
  1. 1.
    Solo DIO, which solves RPM problems through a structured causal chain modeling approach without additional optimization strategies. In the DIO model, the hyperparameter $N$ is set to 16, indicating that the ViT within its image feature extraction module employs a $4\times 4$ patch division strategy;
  2. 2.
    Brando-enhanced DIO, which employs trainable hypothetical incorrect options to tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the Brando framework, the hyperparameter $\beta$ is set to 24, which means the Brando network provides 88 ($\beta-17=8$) constructive hypothetical incorrect options to the DIO in each iteration;
  3. 3.
    WORLD-augmented DIO, utilizing a Gaussian Mixture Model for feature distribution modeling to provide an infinite set of weighted incorrect options for tighting the lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the WORLD framework, the hyperparameter $M$ is set to 256, equipping the GMM employed for DIO with 256 components. This configuration enables the generation of $M^{N}=256^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options.
  4. 4.
    DIEGO-guided DIO, incorporating metadata supervision to directly rectify semantic biases in the causal chain of abstract attribute patterns. In the DIEGO framework, the parameter $F$ is determined by the number of metadata modality in the targeted RPM problem, rather than being arbitrarily set.
This paper conducted experiments on the RAVEN [19], I-RAVEN[34] and PGM [20] datasets, centered around these four configurations. To ensure fair comparison, we uniformly align the optimizer hyperparameters, dataset scales, and batch sizes for optimizing the aforementioned four configurations with those adopted by the current state-of-the-art RS-Tran model [31].
X-A Experiment on RAVEN and I-RAVEN Problems
The performance of our four configurations alongside previous models on the RAVEN and I-RAVEN datasets is summarized in Table I.
TABLE I: Reasoning Accuracies on RAVEN and I-RAVEN.
Test Accuracy(%)
Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
CoPINet[23] 52.96/22.84 49.45/24.50 61.55/31.10 52.15/25.35 68.10/20.60 65.40/19.85 39.55/19.00 34.55/19.45
PrAE Learner[28] 65.03/77.02 76.50/90.45 78.60/85.35 28.55/45.60 90.05/96.25 90.85/97.35 48.05/63.45 42.60/60.70
SAVIR-T [27] 94.0/98.1 97.8/99.5 94.7/98.1 83.8/93.8 97.8/99.6 98.2/99.1 97.6/99.5 88.0/97.2
MRNet [25] 96.6/- -/- -/- -/- -/- -/- -/- -/-
RS-TRAN[31] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
Triple-CFN[33] 98.9/99.1 100.0/100.0 99.7/99.8 96.2/97.5 99.8/99.9 99.8/99.9 99.9/99.9 97.0/97.3
DIO 98.8/99.3 100.0/100.0 99.6/99.7 96.1/97.4 99.8/99.9 99.8/99.9 99.8/99.8 96.7/97.9
DIO+Brando 99.4/99.5 100.0/100.0 99.8/99.8 98.4/99.0 100.0/100.0 99.9/100.0 100.0/100.0 98.6/98.9
DIO+WORLD 99.5/99.7 100.0/100.0 99.8/99.8 98.8/99.5 99.9/100.0 100.0/100.0 100.0/100.0 99.0/99.5
DIO+DIEGO 99.9/99.9 100.0/100.0 99.9/99.9 99.7/99.8 100.0/100.00 100.0/100.0 100.0/100.0 99.8/99.8
Experimental results recorded in Table I indicate that all four model configurations proposed in this study outperform all existing baselines on both the RAVEN and I-RAVEN datasets. Notably, the DIO+DIEGO configuration achieves a peak accuracy of 99.9%, delivering significant performance gains over prior state-of-the-art models (Triple-CFN[33] and RS-TRAN[31]). Particularly in complex reasoning tasks such as $\times3$ Grid and O-IG, the DIO series models demonstrate distinct advantages, validating the generalizability and robustness of their architectural design for abstract visual reasoning challenges.
More critically, the performance achieved by the DIEGO method has overturned a key experimental conclusion in prior studies[33, 25, 31], which posited that the progressive pattern supervision directly implemented from metadata would inevitably undermine the model’s reasoning capabilities. This logically flawed assertion is obsolete. Furthermore, our experiments demonstrate that when no additional supervision signals are incorporated, the strength of mutual information inferred by the solving model serves as the primary determinant of its problem-solving performance, offering critical architectural insights for future abstract reasoning systems. When progressive pattern descriptions are employed as supervisory signals, the supervision utilization strategy combining causal chain modeling with correction of key chain segments proves highly effective.
X-B Experiment on PGM Problem
The performance comparison among the four model configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset is summarized in Table II.
TABLE II: Reasoning accuracies of models on PGM-neutral.
Model Test Accuracy(%)
CoPINet[23] 56.4
SAVIR-T [27] 91.2
MRNet [25] 94.5
RS-CNN[31] 82.8
RS-TRAN[31] 97.5
Triple-CFN[33] 97.8
DIO 98.1
DIO+Brando 99.0
DIO+WORLD 99.2
DIO+DIEGO 99.6
As shown in Table II, DIO achieves state-of-the-art performance on the PGM-neutral subset of the i.i.d. RPM dataset, a benchmark where prior models exhibited relatively low accuracy. Notably, the integration of mutual information lower bound tightening methods (Brando and WORLD) yields substantial performance improvements for DIO. Furthermore, the DIO+DIEGO attains a peak accuracy of 99.6%, thereby challenging the prevailing hypothesis that metadata supervision necessarily impairs relational reasoning capabilities.
The paper also conducted experiments on the o.o.d. sub-task of PGM, with the results documented in Table III.
TABLE III: Generalization Results on PGM.
Dataset
Model/
Task
Interpolation Extrapolation
Held-out
Attribute
shape-colour
Held-out
Attribute
line-type
Held-out
Triples
Held-out
Pairs of
Triples
Held-out
Attribute
Pairs
MRNet[25]
68.1 19.2 16.9 30.1 25.9 55.3 38.4
RS-Tran[31]
77.2 19.2 12.9 24.7 22.2 43.6 28.4
Triple-CFN[33]
80.4 18.4 12.6 25.2 22.0 44.5 29.2
DIO
82.0 18.8 13.6 26.5 23.3 45.5 29.9
DIO+Brando
83.2 18.2 14.2 26.7 23.6 45.8 30.5
DIO+WORLD
87.8 18.9 15.0 27.6 23.8 46.2 30.9
DIO+DIEGO
92.5 13.8 14.4 27.2 28.8 98.1 98.1
The results recorded in Table III reveal that while mutual information tightening and causal chain modeling yield comparable performance gains for DIO on i.i.d. RPM tasks, the latter demonstrates significantly superior generalization capabilities under o.o.d. RPM conditions. To a certain extent, these results suggest that mutual information tightening remains a highly effective approach for i.i.d. reasoning tasks, despite mutual information being a statistical measure that does not inherently encode causal relationships crucial for reasoning. However, when addressing o.o.d. reasoning challenges, adopting causal chain modeling as the methodological cornerstone proves to be a more robust choice.
X-C Experiment on Generative RAVEN and PGM Problems
While DIO+WORLD can solve discriminative RPM problems with high precision, it is also capable of solving generative RPM problems. Specifically, DIO+WORLD can generate solutions for the context matrices instead of making a selection from a predefined bound option pool. This paper proposes that it is reasonable to use a progressive-pattern-driven discriminative model to verify the accuracy of the generative model in solving generative RPM problems. Therefore, we employ DIO+DIEGO to validate the generative accuracy of DIO+WORLD. Prior to this, we must report DIEGO’s discriminative accuracy for progressive patterns under the supervision of RPM metadata to confirm that DIEGO can effectively verify the solution accuracy for generative RPM problems. DIEGO’s reasoning accuracy for progressive patterns on the RAVEN, I-RAVEN, and PGM datasets is recorded in Table IV.
TABLE IV: DIEGO’s Reasoning Accuracy for Progressive Patterns on the RAVEN, I-RAVEN, and PGM Datasets.
Dataset Sub-task Accuracy(%)
RAVEN center entire:99.9
$\times2$ Grid entire:99.9
$\times3$ Grid entire:99.6
L-R left:99.9 right:99.9
U-D left:99.9 right:99.9
OI-C in:99.9 out:99.9
OIG in:99.8 out:99.9
I-RAVEN center entire:99.9
$\times2$ Grid entire:99.9
$\times3$ Grid entire:99.6
L-R left:99.9 right:99.9
U-D left:99.9 right:99.9
OI-C in:99.9 out:99.9
OIG in:99.8 out:99.9
PGM neutral shape:99.7 line:99.9
Table IV indicates that the DIO+DIEGO demonstrates remarkably high accuracy in progressive pattern reasoning for three categories of RPM problems, making it a suitable validation model for generative RPM tasks. Consequently, we evaluate the DIO+WORLD model using DIO+DIEGO and report its accuracy on generative RAVEN, I-RAVEN and PGM-neutral problems in Table V.
TABLE V: Generation Accuracies of DIO+WORLD
Dataset Sub-task Accuracy (%)
RAVEN/I-RAVEN
Center 91.0/91.2
$\times2$ Grid 85.2/85.0
$\times3$ Grid 81.5/81.8
L-R 87.9/88.5
U-D 87.9/88.4
O-IC 89.0/89.3
O-IG 83.0/84.2
PGM Neutral 84.5
Table V clearly demonstrates that, without relying on any additional supervisory signals, DIO+WORLD has achieved highly competitive results in open-ended generative RPM problems, which strongly confirms the significant research value of the DIO model. In addition, these results also demonstrate that although mutual information does not directly indicate causal relationships, it remains highly valuable in constructing reasoning capabilities and can withstand the tests conducted by models driven by causal chains.
XI Ablation Study
In this section, this paper conducts a series of ablation experiments. The ablation experiments will be conducted focusing on the key parameter settings.
XI-A Ablation Study on DIO
The ablation study on DIO focuses on the design of the network’s feed-forward structure. Within DIO’s progressive pattern induction module, a permutation and combination strategy ($C_{n}^{r}$) is employed to extract progressive patterns from row and column features. This approach aims to model the causal chain details in RPM problems. In this ablation experiment, we will remove the $C_{n}^{r}$ operation. Instead, we will directly input the total of six row-column features for progressive pattern extraction. By doing so, we aim to obscure the structure in DIO that aligns with the causal chain, thereby validating the significance of causal chain modeling as a methodology for network construction. The obscured DIO is denoted as DIO. The reasoning accuracy achieved by DIO on key sub-tasks of RAVEN, I-RAVEN, and PGM is recorded in Table VI.
TABLE VI: Reasoning Accuracies of DIO.
Dataset Sub-task Accuracy (%)
RAVEN/I-RAVEN
$\times3$ Grid 95.6/96.8
O-IG 95.5/95.6
PGM Neutral 97.2
Table VI indicates that constructing RPM solving models based on causal chains is a critically important condition for achieving high reasoning performance.
XI-B Ablation Study on Brando
The ablation study on Brando focuses on the hyperparameter $\beta$. This paper demonstrates that the number of constructive incorrect options provided to DIO directly influences the tightness of the variational lower bound on mutual information derived by DIO. Therefore, it is essential to conduct ablation experiments focusing on the number ($\beta$-17) of hypothetical options provided by the Brando network. The reasoning accuracies of DIO+Brando on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $\beta$ values are recorded in Table VII.
TABLE VII: Reasoning Accuracy of DIO+Brando under Different $\beta$ Values.
Dataset Sub-task Accuracy(%)
$\beta$=24 $\beta$=22 $\beta$=20 $\beta$=18 No $\beta$
RAVEN/
I-RAVEN
$\times3$ Grid 98.4 / 99.0 98.0 / 98.5 97.2 / 98.0 96.8 / 97.8 96.1 / 97.4
O-IG 98.6 / 98.9 98.2 / 98.6 97.6 / 98.3 97.5 / 98.3 96.7 / 97.9
PGM Neutral 99.0 98.8 98.8 98.5 98.1
Table VII indicates that even when only a small number of constructive hypothetical options are introduced into DIO, significant performance improvements can already be achieved for DIO.
XI-C Ablation Study on WORLD
The ablation study on WORLD focuses on the hyperparameter $M$. $M$ indicates the number of components in the GMM constructed by World for RPM image features. The value of $M$ determines the maximum diversity of solution distributions that WORLD can generate for DIO, expressed as $M^{N}$. Therefore, ablation experiments centered around the value of $M$ are warranted. The reasoning accuracies of DIO+WORLD on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $M$ values are recorded in Table VIII.
TABLE VIII: Reasoning Accuracy of DIO+WORLD under Different $M$ Values.
Dataset Sub-task Accuracy(%)
MM=256 MM=128 MM=64 MM=32 No $M$
RAVEN/
I-RAVEN
$\times3$ Grid 98.8 / 99.5 98.5 / 99.1 97.9 / 98.5 96.5 / 97.9 96.1 / 97.4
O-IG 99.0 / 99.5 98.6 / 99.0 98.0 / 98.7 97.1 / 98.0 96.7 / 97.9
PGM Neutral 99.2 99.0 98.5 98.0 98.1
The number of components directly affects the complexity of the GMM. As shown in Table VIII, when the number of GMM components is relatively small ($M=32$), meaning that WORLD models a distribution with lower complexity, it will have a negative impact on the performance of DIO in the PGM task. This study posits that the effectiveness of the WORLD method depends on whether the GMM it constructs can effectively represent the inherent complexity of the natural distribution of features encoded by DIO for RPM images. Specifically, when $M$ is set to 32, this GMM with a low number of components cannot adequately capture the complexity of the natural distribution of PGM image features. Forcing the distribution of PGM image features with higher complexity into this simplified GMM will inevitably lead to a performance degradation of DIO.
WORLD introduces an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM. As a highly versatile distribution model, the GMM theoretically ensures that the WORLD method has a considerable scope of effectiveness. However, the phenomenon where WORLD’s modeling of a GMM with a low number of components causes performance degradation in DIO imposes a crucial prerequisite for the validity of WORLD: the GMM modeled by WORLD must be capable of accommodating the natural complexity of the target problem.
Certainly, in traditional deep learning tasks, underfitting is a prevalent and expected phenomenon that arises when neural networks lack sufficient capacity to handle the complexity of problems. However, the WORLD method makes this phenomenon manageable through its $M$-dependent capacity control. Therefore, particular attention must be paid to the parameter $M$ when applying and adjusting this method.
XII Conclusion
This study has constructed a complete abstract reasoning solution through systematic innovations, with four core innovations forming a progressive technical system that achieves significant progress in both theoretical breakthroughs and empirical performance:
  1. 1.
    Causal Chain Modeling
    • First established a complete causal chain analysis framework of “Image $\rightarrow$ Abstract Attributes $\rightarrow$ Progressive Patterns $\rightarrow$ Pattern Consistency $\rightarrow$ Answer”
    • Provided an explainable causal relationship modeling paradigm for deep learning models.
  2. 2.
    Brando Error Mapping Mechanism
    • Developed a trainable hypothetical incorrect option mapping mechanism to explicitly tighten the variational lower bound of mutual information.
    • Creatively introduced mutual information tightening concepts into abstract reasoning frameworks, thereby opening novel avenues for variational inference methodologies.
  3. 3.
    WORLD Feature Modeling
    • Constructed a Gaussian Mixture Model (GMM) for RPM image feature distributions, enabling infinite and semantically-rich incorrect option sampling
    • Empowering DIO to solve open-ended generative RPM problems.
  4. 4.
    DIEGO Causal Correction Innovation
    • Introduced external metadata to construct supervision signals, precisely correcting semantic biases in the causal chain modeled by DIO
    • Established mappings between model reasoning chains and ground-truth causal chains, aligning internal representations with human cognitive frameworks
[…]
This study’s multi-dimensional innovations not only achieve technical breakthroughs but, more importantly, provide a new paradigm for causal explainability research in deep learning models, holding significant theoretical value and practical implications for advancing artificial intelligence toward higher-order cognitive intelligence.

### v3

@@
 Therefore, this paper proposes three methods, namely Brando, WORLD, and DIEGO, to refine DIO’s learning objectives.
-As a bonus, when DIO adopts DIEGO method, it demonstrates the capability to solve generative RPM problems.
-The images in the RPM instances of this paper have adopted the annotation approach as illustrated in Figure 2, which enables clearer explanation of our model and method.
+As a bonus, when DIO adopts the WORLD method, it demonstrates the capability to solve generative RPM problems.
+The RPM instances in this paper use the annotation approach illustrated in Figure 2, which allows for a clearer explanation of our model and methods.
 Figure 2: Annotations of Images Within an RPM Instance.
-Spcifically, we denote the 16 RPM images as $\{x_{i}|i\in[1,16]\}$.
+Specifically, we denote the 16 RPM images as $\{x_{i}|i\in[1,16]\}$.
 IV DIO: A Baseline Model
@@
 IV-A The Causal Chains in RPM Problems.
-The solving process of RPM problems is as follows: The solving model needs to discern the progressive pattern of abstract image attributes from the context matrices $\{x_{i}|i\in[1,8]\}$ in an RPM instance.
-Subsequently, it should select the ideal option $x_{\alpha}$ from the eight options $\{x_{i}|i\in[9,16]\}$ in the option pool that will not disrupt this progressive pattern [35].
+The process of solving RPM problems proceeds as follows: the model must first identify the progressive patterns governing abstract image attributes across the context matrices $\{x_{i}|i\in[1,8]\}$, and then select the correct option $x_{\alpha}$, which comprehensively continues these patterns [35], from the option pool $\{x_{i}|i\in[9,16]\}$.
 This implies that there exists a causal chain in RPM problems: RPM images $\rightarrow$ abstract attributes $\rightarrow$ progressive patterns of attributes $\rightarrow$ consistency of progressive patterns $\rightarrow$ correctness of options.
-This paper proposes that if the structural design of the RPM solving model is guided by this causal chain, the constructed model will demonstrate high performance.
-This prompted us to design the DIO model.
-In accordance with the causal chains in RPM problems, the DIO model is divided into four modules: the image feature extraction module, the progressive pattern induction module, the progressive pattern consistency evaluation module, and the option evaluation module.
+This paper proposes that designing the structure of an RPM solving model in accordance with this causal chain will yield high performance, which prompted us to design the DIO model.
+Accordingly, the DIO comprises four modules: the image feature extraction module, the progressive pattern induction module, the progressive pattern consistency evaluation module, and the option evaluation module.
 IV-B The Image Feature Extraction Module.
@@
 Consequently, when extracting features from RPM images, we opt for a Visual Transformer (ViT) [21] that prioritizes image context.
-Figure 3 illustrates the process by which ViT extracts tokenized image features $\{z_{ij}|j\in[1,N]\}$ from the input image $x_{i}$, where $j$ denotes the token index of the image features and $N$ denotes the total patch count.
+Figure 3 illustrates the process by which ViT extracts the tokenized image features $\{z_{ij}|j\in[1,N]\}$ from the input image $x_{i}$, where $j$ denotes the token index of the image features and $N$ denotes the total patch count.
 Figure 3: The Image Feature Extraction Module.
@@
 The deliberate design of RPM problems necessitates that participants must simultaneously observe two rows and two columns to identify the progressive patterns within the matrix.
-This represents a critical link in the causal chain of RPM problems.
+This represents a critical segment in the causal chain of RPM problems.
 Consequently, the progressive pattern induction module of DIO adopts the structure illustrated in Figure 4.
 Figure 4: The Progressive Pattern Induction Module.
-Figure 4 indicates that after incorporating the features of option $\{z_{\alpha j}|j\in[1,N]\}$ into the context matrix $\{z_{ij}|i\in[1,8],j\in[1,N]\}$, DIO will summarize the progressive patterns among the features that share the same token index $j$.
-DIO first utilizes an MLP (Multi-Layer Perceptron) to extract row and column features from the matrix.
+Figure 4 indicates that after appending the option features $\{z_{\alpha j}|j\in[1,N]\}$ into the context matrix $\{z_{ij}|i\in[1,8],j\in[1,N]\}$, DIO summarizes the progressive patterns among the features that share the same token index $j$.
+First, DIO utilizes a Multi-Layer Perceptron (MLP) to extract row and column features from the matrix.
 Subsequently, it randomly selects two row features from the three available row features and two column features from the three column features, combining these four features for further processing.
 After binding four trainable vectors of the same dimension to the selected four features, DIO employs a Transformer Encoder to process them into eight logical outputs.
-Among the output vector groups, the four vectors corresponding to the trainable vectors are denoted as $\{P_{jl}|l\in[1,4]\}$.
-$\{P_{jl}|l\in[1,4]\}$ is regarded as the partial representation of progressive patterns within the matrix.
-If all the $C_{3}^{2}\cdot C_{3}^{2}=9$ possible ways of selecting row and column features are traversed (where $C_{n}^{r}$ denotes the number of combinations of $n$ items taken $r$ at a time), then a comprehensive representation of progressive patterns, $\{P_{jkl}|k\in[1,9],l\in[1,4]\}$, can be obtained.
+Among the output vector groups, the four vectors corresponding to the trainable vectors are denoted as $\{P_{\alpha jl}|l\in[1,4]\}$.
+$\{P_{\alpha jl}|l\in[1,4]\}$ are regarded as the partial representations of progressive patterns within the matrix $\{z_{ij},z_{\alpha j}|i\in[1,8]\}$.
+If all $C_{3}^{2}\cdot C_{3}^{2}=9$ possible row–column selections are considered (where $C_{n}^{r}$ denotes the number of combinations of $n$ items taken $r$ at a time), we obtain comprehensive representations $\{P_{\alpha jkl}|k\in[1,9],l\in[1,4]\}$ of these patterns.
 This module can be formally expressed as:
@@
 IV-D The Progressive Pattern Consistency Evaluation Module
-As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive pattern representations that share the same $j$-index and $l$-index.
-Specifically, DIO applies a Transformer Encoder to perform dimension-invariant processing on $\{P_{\alpha jkl}|k\in[1,9]\}$, resulting in $\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$.
-Subsequently, each vector in $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$ is mapped to a scalar via an MLP, serving as a logical score $s_{\alpha jkl}$.
+As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive pattern representations $P_{\alpha jkl}$ that share the same $j$-index and $l$-index.
+Specifically, the Transformer Encoder performs a dimension-preserving mapping on $\{P_{\alpha jkl}|k\in[1,9]\}$ to obtain $\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$.
+Repeating this for all $j\in[1,N]$ and $j\in[1,4]$ yields the complete representation set $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$.
+Finally, each vector in this full set is mapped to a scalar via an MLP, producing the consistency score $s_{\alpha jkl}$.
 This process can be formally expressed as:
@@
 IV-E The Option Evaluation Module.
-The DIO’s option evaluation module computes the score $S_{\alpha}$ that reflects the correctness of option $x_{\alpha}$ by summing $\{s_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$ according to the formula:
+DIO’s option evaluation module computes the score $S_{\alpha}$ that reflects the correctness of option $x_{\alpha}$ by averaging the consistency scores $\{s_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$:
 $S_{\alpha}$ $=\frac{1}{N\cdot 9\cdot 4}\big{(}\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{4}s_{\alpha jkl}\big{)}$ (6)
-Use the softmax function to map $\{S_{i}|i\in[8,16]\}$ to a probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$:
+The softmax function then maps $\{S_{i}|i\in[8,16]\}$ to a probability distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ parameterized by DIO:
 $\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{S_{\alpha}}}{\sum_{{\hat{\alpha}}=9}^{16}e^{S_{{\hat{\alpha}}}}}$ (7)
-We introduce a constraint mechanism that correlates the DIO-modeled probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$.
-This constraint enables joint optimization of all trainable parameters across DIO’s modules:
-${\ell}_{\text{DIO}}=-\mathbb{E}\left[\log\frac{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\sum_{\hat{\alpha}=9,\hat{\alpha}\neq\alpha}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (8)
-Where $\mathbb{E}(\cdot)$ denotes the expectation operation.
+We introduce a loss function ${\ell}_{\text{DIO}}$ to refine the DIO-parameterized distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$, thereby enabling end-to-end training of all parameters across DIO’s modules:
+${\ell}_{\text{DIO}}=-\mathbb{E}\left[\log\frac{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)}{\sum_{\hat{\alpha}=9,\hat{\alpha}\neq\alpha}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)}\right]$ (8)
+where $\mathbb{E}[\cdot]$ denotes the expectation over the data distribution.
 V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM Solving Model
-This paper proposes that architecture optimization based on causal chains can effectively enhance the performance of deep learning networks in solving RPM problems.
-However, through further research, we have found that this methodology still cannot fundamentally improve the logical reasoning capabilities of deep learning networks.
+This paper proposes that causal-chain-based architecture optimization can effectively enhance the performance of deep learning networks in solving RPM problems.
+However, further investigation reveals that this methodology still fails to fundamentally improve the logical reasoning capabilities of deep learning networks.
 V-A The learning objective of the DIO
-By abstracting away the specific architecture of DIO (i.e., merging equations (1)-(6)), we obtain a unified formulation that highlights its core inductive mechanism:
+By abstracting the specific architecture of DIO (i.e., by merging equations (1)–(6)), we obtain a unified formulation that highlights its core inductive mechanism:
 $S_{\alpha}$ $=\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})$ (9)
-Leveraging the probabilistic framework in Equation (7), DIO can be interpreted as a tunable probabilistic model where $\theta$ represents learnable parameters:
+By substituting Equation (7), DIO can instantiate the parameterized probability distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$:
 $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (10)
-Notably, the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ can be expressed as:
+Notably, $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, which denotes the mutual information between the context matrix $\{x_{i}|i\in[1,8]\}$ and the correct option $x_{\alpha}$, admits the following variational lower bound:
 $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}P(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (11)
-Where $K$ denotes the number of incorrect options.
-If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a sufficient approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ , then substituting the $\hat{P}$ defined in Equation (10) into the Equation (12) yields a mutual information lower bound expressed by DIO:
-$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]$ (12)
-When $K\geq 1$, that is, when there exists at least one incorrect option, the lower bound aligns with DIO’s optimization objective, as expressed by the following equation:
+where $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ represents the true probability that $x_{\alpha}$ can serve as the correct option for the given context matrix $\{x_{i}|i\in[1,8]\}$ and $K$ denotes the number of all potential incorrect options.
+If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a sufficient approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$, then substituting Equation (10) into Equation (11) yields a mutual information lower bound parameterized by DIO:
+$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right],$ (12)
+$\quad\quad\quad\quad\quad\quad\quad\text{where}\quad K=7$
+It can be observed that the parameterized lower bound within Equation (12) aligns with DIO’s optimization objective, as expressed by the following equation:
 $\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]=-\ell_{\text{DIO}}+\log K$ (13)
 Equation (13) reveals that DIO’s learning objective (i.e.
-maximizing $\ell_{\text{DIO}}$) is neither the human-like reasoning logic embedded in RPM problems nor even the suboptimal approach of directly maximizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather optimizing the variational lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
+maximizing $\ell_{\text{DIO}}$) is neither the human-like reasoning logic embedded in RPM problems nor even the suboptimal approach of directly maximizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather optimizing the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
 Furthermore, based on the properties of logarithmic functions, the following inequality holds:
-${\ell}_{\text{DIO}}$ $<\mathbb{E}\left[\log\frac{\sum_{\hat{\alpha}=9}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8})}\right]$
-$=-1\cdot\log\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ (14)
-The right-hand side of the Inequality (V-A) embodies the classical formulation for optimizing the probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the cross-entropy (CE) loss function, which is the typical paradigm adopted by most previous end-to-end RPM solving models.
-From this, it can be inferred that for those end-to-end RPM solving models that have also focused on constructing the probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ in the past, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
-Equation (13) and Inequality (V-A) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
+$-{\ell}_{\text{DIO}}>\mathbb{E}\left[\log\frac{\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\sum_{\hat{\alpha}=9}^{16}\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (14)
+The right-hand side of Inequality (14) is precisely the classical formulation for optimizing the probabilistic distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the Cross-Entropy (CE) loss function.
+From this, it can be inferred that for those past end-to-end RPM solving models that parameterize $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ and employ the CE loss function, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
+Equation (13) and Inequality (14) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
 In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
 V-B Introducing Hypothetical Options to Derive a Tight Lower Bound for Mutual Information.
-This paper argues that employing DIO to derive a tighter variational lower bound and directing its optimization objective toward the maximization of this new lower bound can more effectively maximize mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, thereby further DIO’s performance.
-By analyzing Formula (12), we deduce that increasing the number of incorrect options $K$ encountered by DIO can guide its learning objective toward a tighter variational lower bound.
+This paper contends that by utilizing DIO to parameterize a tighter variational lower bound and adjusting its optimization objective to maximize this new lower bound, we can achieve a more effective maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, ultimately enhancing DIO’s performance.
+By analyzing Formula (12), we deduce that increasing the number $K$ of incorrect options encountered by DIO can guide its learning objective toward a tighter variational lower bound.
 However, RPM problem instances only offer a limited number (seven) of incorrect options.
-To this end, this paper attempts to introduce additional hypothetical incorrect options for DIO.
-Specifically, we introduce a set of hypothetical options $\{x_{i}|i\in[17,\beta]\}$ into the training pipeline of DIO.
+To this end, this paper proposes to introduce additional hypothetical incorrect options for DIO.
+Specifically, we attempt to introduce a set of hypothetical options $\{x_{i}|i\in[17,\beta]\}$ into the training pipeline of DIO.
 Consequently, the conditional distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ is reparameterized as:
@@
 Intuitively, the looseness of the updated lower bound has been reduced.
-However, In Formula (16), it is not difficult to observe that if these hypothetical options cannot change DIO’s decision-making then this new lower bound, when serving as an optimization objective, will not undergo any fundamental improvement.
-In other words, when the probabilities assigned by the DIO model to the hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ approach zero, which means that these options do not alter the probability distribution $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16]\}$ that DIO assigns to the original options $\{x_{i}|i\in[9,16]\}$, the variational lower bounds derived from Equation (16) and Equation (12) become equivalent given that both function as loss functions for neural networks.
-This implies that unconstructive options contribute nothing to tightening the lower bound.
-Therefore, this paper proposes introducing extra constructive hypothetical options into DIO—options that reinforce its endorsement of the correct options while simultaneously strengthening its repulsion of the incorrect ones—so as to steer DIO’s learning objective toward a tighter variational lower bound and ultimately further enhance its performance.
+However, as shown in Formula (16), if these introduced hypothetical options fail to influence DIO’s decisions, this new lower bound, when used as an optimization objective, will not bring DIO any significant improvement.
+Specifically, when the probabilities that DIO assigns to the hypothetical options $\{x_{i}|i\in[17,\beta]\}$ approach zero, whereby these options cease to perturb the original distribution $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16]\}$, the variational lower bounds derived from Equation (16) and Equation (12) become equivalent when employed as loss functions.
+This demonstrates that unconstructive options, which are readily assigned low weights and low probabilities by DIO, contribute nothing to tightening the lower bound.
+Therefore, this paper proposes introducing extra constructive hypothetical options into DIO, options which can reinforce acceptance of the correct options and simultaneously strengthen repulsion of the incorrect ones, so as to steer its learning objective toward a tighter variational lower bound and, in turn, ultimately further enhance its performance.
 VI Brando: Tightening the Variational Lower Bound of Mutual Information by Constructing Hypothetical Options.
@@
 The RPM problem expects the training process of the solving model to involve only a limited number of supervised signals, such as the index of correct options.
-Relying on manual design to create constructive options goes against the original intent of RPM problems.
+Relying on manual design to create constructive options goes against this original intent of RPM problems.
 Therefore, this paper proposes training a set of constructive options instead.
-This methodology is referred to as Brando (Bounded Random Options Mapping Method) in this paper.
+This methodology is referred to as Brando (Bounded Random Options Mapping Method).
 This paper establishes a set of trainable hypothetical options $\{x_{i}|i\in[17,\beta]\}$ for the context matrix $\{x_{i}|i\in[1,8]\}$.
 This paper is not concerned with the specific attributes of these options; rather, it focuses solely on how much these options can alter DIO’s evaluation of the original options $\{x_{i}|i\in[9,16]\}$.
-In other words, we are solely concerned with the constructiveness of these hypothetical options for DIO, rather than their specific configurations.
-Since these options are expected to improve $S_{\alpha}$ while reducing $\{S_{\hat{\alpha}}|{\hat{\alpha}}\in[9,16],{\hat{\alpha}}\neq\alpha\}$, this paper designs a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options and further optimize DIO:
+In other words, we are solely concerned with the constructiveness of these hypothetical options with respect to the context matrix $\{x_{i}|i\in[1,8]\}$ in DIO, rather than their specific configurations.
+To this end, this paper designs a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options for their constructiveness:
 $\ell_{\text{Brando}}=\ell_{\text{up}}+\ell_{down}$ (17)
@@
 $\ell_{\text{down}}=-\frac{1}{\beta-17}\sum_{{\hat{\beta}}=17}^{\beta}\log\frac{e^{\text{DIO}(\{x_{i},x_{\hat{\beta}}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (19)
-In these noval loss function:
+In this novel loss function:
   1. 1.
@@
 In other words, $\ell_{\text{down}}$ ensures hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ outscore the original incorrect options $\{x_{i}\mid i\in[9,16],i\neq\alpha\}$ under DIO.
-In loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive options.
+In the loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive options.
 When optimizing the DIO using $\ell_{\text{Brando}}$ as the loss function, it can induce bounded hypothetical options while leveraging a tighter variational lower bound to enable DIO to maximize the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.
@@
 $\{x_{i}|i\in[17,\beta]\}=\mathcal{F}(\{x_{i}|i\in[1,8]\})$ (20)
-Evidently, the mapping $\mathcal{F}(\cdot)$ is required to derive a variable number ($\beta$) and diverse hypothetical options for the context matrix $\{x_{i}|i\in[1,8]\}$.
+Evidently, the mapping $\mathcal{F}(\cdot)$ is required to derive a variable number ($\beta-17$) and diverse hypothetical options for the context matrix $\{x_{i}|i\in[1,8]\}$.
 This indicates that conventional deep neural networks struggle to undertake the task of the mapping $\mathcal{F}(\cdot)$.
@@
 Additionally, from the perspective of neural network optimization, achieving a many-to-many mapping is infeasible.
+Given these challenges, we need to explore an alternative approach.
 Therefore, this paper proposes a many-to-one mapping approach:
@@
 $\text{Where}\quad o_{\hat{i}}\sim N(\mu,\sigma^{2})$
-Formula (23) shows that we move away from the option-mapping paradigm and instead employs a representation-mapping approach, easing the learning pressure on Brando.
+Formula (23) shows that we move away from the option-mapping paradigm and instead adopt a representation-mapping approach, easing the learning pressure on Brando.
 According to Formula (23), we have designed the structure of the Brando network as shown in the Figure 5.
 Figure 5: The Brando network.
-As indicated in Figure 5, the Brando network is formed by stacking multiple logical layers, with no parameter sharing between each layer.
+As indicated in Figure 5, the Brando network is formed by stacking multiple logical layers, with no parameter sharing among each layer.
 The input to each logical layer consists of $\{z_{ij}|i\in[1,8]\}$ and $o_{\hat{i}}$.
 Except for the first layer, the input to all subsequent layers comes from the output of the preceding layer.
-For the first layer, the $\{z_{ij}|i\in[1,8]\}$ is derived from the representations of the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution parameterized by mean and variance $N(\mu,\sigma^{2})$.
-For the output $o_{\hat{i}}$ from the final logic layer, this paper treats it as the representation $z_{\hat{i}j}$ of a hypothetical option provided by Brando for the context matrix $\{z_{ij}|i\in[1,8]\}$.
-For the specific structure of each logical layer, it consists of a set of mapping matrices $\{W_{q},W_{k},W_{v}\}$, a multi-head self-attention mechanism, a feed-forward layer, a memory bank $M_{n}$, a buffer $B_{n}$, and a reparameterization head.
+For the first layer, $\{z_{ij}|i\in[1,8]\}$ are the features extracted by DIO from the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution $N(\mu,\sigma^{2})$ parameterized by mean $\mu$ and variance $\sigma^{2}$.
+For the final layer, this paper treats its output $o_{\hat{i}}$ as the representation $z_{\hat{i}j}$ of a hypothetical option $x_{\hat{i}}$ provided by Brando with respect to the context matrix $\{x_{i}|i\in[1,8]\}$.
+For the specific structure of each logical layer, it consists of a set of mapping matrices $\{W_{q},W_{k},W_{v}\}$, a multi-head self-attention mechanism, a feedforward layer (FFL), a dynamic memory bank $M_{n}$, a global buffer $B_{n}$, a learnable scalar $\tau$, and a reparameterization head.
+The feedforward process of each logical layer is as follows:
   1. 1.
-    When the input set $\{z_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ is fed into a single logical layer, the mapping matrices $W_{q}$, $W_{k}$, and $W_{v}$ transform this input into query ($Q$), key ($K$), and value ($V$) representations, respectively.
+    When the features $\{z_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ are input into a single logical layer, the mapping matrices $W_{q}$, $W_{k}$, and $W_{v}$ transform them into query ($Q$), key ($K$), and value ($V$) representations, respectively.
   2. 2.
-    After these query ($Q$), key ($K$), and value ($V$) representations are appended positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$.
+    After these $Q$, $K$, and $V$ representations are appended positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$:
   3. 3.
-    As indicated by the computational step 1 in the diagram, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input representations stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the output set $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$.
+    As indicated by the computation step 1 in the Figure 5, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input feature information stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the memory outputs $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$.
 The learning process can be interpreted as:
@@
   4. 4.
-    In computation steps 2 and 3, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the key ($K$) and value ($V$) mappings derived from the current input representations.
+    As indicated by the computation steps 2 and 3 in the Figure 5, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the $K$ and $V$ representations derived from the current input features.
 The update process can be expressed as:
@@
   5. 5.
-    After the learnable scalar $\tau$ is compressed into a scaling coefficient via the sigmoid function, a weighted summation is performed on $\tilde{o}_{\hat{i}}$ and $\hat{o}_{\hat{i}}$ using this coefficient to update $\hat{o}_{\hat{i}}$, enabling it to balance its own reasoning with image representations from other instances.
+    After the learnable scalar $\tau$ is compressed into a scaling coefficient via the sigmoid function, a weighted summation is performed on the memory output $\tilde{o}_{\hat{i}}$ and attention output $\hat{o}_{\hat{i}}$ using this coefficient.
+The summation result is utilized to update $\hat{o}_{\hat{i}}$, enabling it to balance current reasoning with historical feature information.
+The update process of $\hat{o}_{\hat{i}}$ can be expressed as:
     $\hat{o}_{\hat{i}}\leftarrow[1-\text{Sigmoid}(\tau)]\hat{o}_{\hat{i}}+\text{Sigmoid}(\tau)\tilde{o}_{\hat{i}}$ (27)
   6. 6.
-    The attention outputs $\{\hat{z}_{ij}|i\in[1,8]\}$ with the updated $\hat{o}_{\hat{i}}$, which has undergone skip connection with logical inputs, is processed by the feed-forward layer and subjected to another skip connection, yielding the logical output $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}$.
-This process can be expressed as:
+    The attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ (where $\hat{o}_{\hat{i}}$ has been updated), after undergoing residual connection with the input features, are processed by the feedforward layer and then subjected to another residual connection with their unprocessed state, ultimately yielding the logical outputs $\{z_{ij},o_{\hat{i}}|i\in[1,8]\}$.
+These two residual connections can be formalized as:
     $Z\leftarrow\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}+\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ (28)
     $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}\leftarrow\text{FFL}(Z)+Z$ (29)
-    Here, FFL()($\cdot$) refers to the feed-forward layer.
   7. 7.
-    It is noteworthy that only after the $^{o}_{\hat{i}}$ within the logical output $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}$ have been processed by the reparameterization head can they serve as the final output.
-The process of the reparameterization head can be interpreted as follows:
+    Importantly, the component $^{o}_{\hat{i}}$ in the logical output $\{z_{ij},o_{\hat{i}}|i\in[1,8]\}$ must be processed by the reparameterization head before serving as the final output.
+The reparameterization head operates as follows:
 […]
@@
     The $\epsilon$ represents a random sample drawn from $N(0,1)$ and MLP()($\cdot$) refers to a Multi-Layer Perceptron capable of doubling the dimension of $^{o}_{\hat{i}}$.
-It is not hard to find that the memory bank $M_{n}$ and buffer $B_{n}$ in each logical layer will incorporate new memories as a new batch of context matrices is observed.
-Therefore, the initial form of the memory bank requires a clear definition.
+As indicated in Formulas (25)-(26), in each logical layer, the memory bank $M_{n}$ and the buffer $B_{n}$ are dynamically updated as each new batch of context matrices is input.
+Therefore, their initial values must be explicitly defined:
 $M_{0}=0,\,B_{0}=1\times 10^{-9}$ (32)
 Furthermore, the first update of the memory bank and the buffer can be explained as follows:
-$M_{1}=\phi(K)^{T}V,\,B_{1}=\sum\phi(K)$ (33)
-From the learnable Gaussian distribution ${N}({\mu},{\sigma}^{2})$, we sample random vectors $^{o}_{\hat{i}}$, which undergo sequential refinement in the Brando network through dynamic interaction with both current context matrix representations and historical image representations accumulated across multi-round training iterations.
+$M_{1}=\phi(K)^{T}V,\,B_{1}=\sum\phi(K)+10^{-9}$ (33)
+From the learnable Gaussian distribution ${N}({\mu},{\sigma}^{2})$, we sample random vectors $^{o}_{\hat{i}}$, which undergo sequential refinement in the Brando network through dynamic interaction with both current context matrix representations and historical image representations accumulated over multiple rounds of training iterations.
 This iterative conditioning process enables $^{o}_{\hat{i}}$ to progressively assimilate visual patterns observed throughout the training trajectory, ultimately yielding the ultimate hypothesis representation ${z}_{\hat{i}j}$.
 Moreover, alterations in the initial promoter $^{o}_{\hat{i}}$ are capable of inducing changes in the ultimate option representation $z_{\hat{i}j}$.
-This satisfies the demand to map diverse hypothetical options for the context matrix.
+This satisfies the requirement to map diverse hypothetical options for the context matrix.
 VI-C The Combination of Brando and DIO.
 Prior to integrating the Brando network with DIO, DIO must be pre-trained on the original option pool $\{x_{i}|i\in[9,16]\}$ using the loss function $\ell_{\text{DIO}}$.
-Once DIO is pre-trained to approach its optimal performance, the Brando network will be integrated with DIO according to the method illustrated in Figure 6.
+Once DIO is pre-trained to reach a state close to its optimal performance, the Brando network will be integrated with DIO according to the method illustrated in Figure 6.
 Figure 6: The Combination Method of DIO and Brando.
-Subsequently, the loss function will be switched from $\ell_{\text{DIO}}$ to $\ell_{\text{Brando}}$ for continued training of this integrated system, thereby further optimizing DIO’s performance.
+Subsequently, the loss function will be switched from $\ell_{\text{DIO}}$ to $\ell_{\text{Brando}}$ to continue training this integrated system, thereby further optimizing DIO’s performance.
 Another key point is that during the joint training process, DIO’s parameters excluding those in the image feature extraction module will be frozen every other iteration.
@@
 Moreover, providing a sufficient number of hypothetical options for DIO using Brando entails considerable computational overhead.
-Therefore, this paper argues that Brando is not a flawless methodology.
+Therefore, this paper argues that Brando is not a flawless method.
 To this end, this paper proposes a method named WORLD (Weighted Option-Represented Lower-bound Delimitation Method), which breaks away from the paradigm of the mapping $\mathcal{F}(\cdot)$.
 This method can also tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
-VII-A Modeling is Preferable to Estimation.
+VII-A Local Estimation << Global Estimation << Global Modeling.
 Formula (12) indicates that the key to guiding the learning objective of solving model towards a tighter lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$ lies in providing additional incorrect options to the model for each context matrix.
-Moreover, this paper also points out that the introduced additional incorrect options must avoid being unconstructive, that is, after being incorporated into the training process of the solving model, they should be able to enhance the decision-making capability of the solving model.
+Moreover, this paper also points out that the introduced additional incorrect options must avoid being unconstructive.
 The Brando network essentially functions as a distribution transformation mechanism, whose task is to map samples $o_{\hat{i}}$ drawn from a Gaussian distribution $N(\mu,\sigma^{2})$ into features $z_{\hat{i}j}$ of incorrect options customized for a given context matrix.
@@
 However, accurately estimating the distribution of RPM image features $z_{ij}$ poses significant challenges, and even if such a distribution were successfully estimated, targeted sampling from it would remain highly non-trivial.
-Therefore, this paper argues that explicitly modeling the distribution of $z_{ij}$ during the process where the DIO’s feature extraction module encodes $x_{i}$ into $z_{ij}$ constitutes an effective approach, since this method eliminates the need for estimation and facilitates targeted sampling through direct distributional control.
+Therefore, this paper argues that explicitly modeling the distribution of $z_{ij}$ during the process where the DIO’s feature extraction module encodes $x_{i}$ into $z_{ij}$ constitutes an effective approach.
+Because this method can obtain the distribution of $z_{ij}$ without estimation and facilitate targeted sampling by directly controlling the distribution.
 This paper has found that the Gaussian Mixture Model (GMM) [36] offers remarkable advantages, as it not only facilitates efficient modeling but also enables targeted sampling.
@@
 Then, for a given context matrix, after identifying the components corresponding to its correct options within this GMM, we can reasonably regard the components not corresponding to the correct options as the source of incorrect option features.
-Based on this, we are able to sample comprehensive and diverse incorrect option features from these specific components for DIO.
-Given these advantages, this section proposes the WORLD method to model the distribution of $z_{ij}$ as a GMM.
+Based on this, we are able to sample comprehensive and diverse incorrect option features from these specific components for DIO to tighten its learning objective.
+Given these advantages, we shift our focus to transforming the underlying distribution of RPM image features into a GMM.
 VII-B Model the Distribution of RPM Image Features as a Gaussian Mixture Model.
-To accurately model the distribution of RPM image features $z_{ij}$ as a GMM, this paper proposes the distribution modeling method WORLD tailored for DIO.
-Specifically, this paper introduces a trainable vector set $\{U_{m}|m\in[1,M]\}$ and regards it as the mean vectors of a GMM.
-Using this approach, we construct a predefined GMM with $M$ components.
-Subsequently, the image features $z_{ij}$ encoded by DIO are projected onto this predefined GMM framework, thereby modeling the distributional characteristics of $z_{ij}$.
-The way in which DIO’s related modules extract features from RPM images also only requires minor modifications, so that the distribution of the extracted features $z_{ij}$ can be more easily projected onto the predefined GMM $\{U_{m}|m\in[1,M]\}$.
+To model the distribution of features encoded by DIO for RPM images as a GMM, this paper proposes the WORLD method.
+Specifically, this paper introduces a trainable vector set $\{U_{m}|m\in[1,M]\}$ and considers it as the mean vectors of a GMM.
+Using this approach, we define a GMM comprising $M$ components, where the means of all components are trainable while the variances remain unknown.
+Subsequently, by projecting the image features $z_{ij}$ encoded by DIO onto this constructed GMM framework, we can model the distributional characteristics of $z_{ij}$.
+In order to enable the features $z_{ij}$ to be more easily projected onto the constructed GMM $\{U_{m}|m\in[1,M]\}$, the way DIO encodes features for RPM images requires minor modifications.
 The modifications are shown in Figure 7.
@@
 $\{u_{ij},b_{ij}\}=\text{ViT}(x_{i}),\,z_{ij}=u_{ij}+b_{ij}$ (34)
-The newly devised derivation method for $z_{ij}$ ensures that, when $b_{ij}$ follows a standard Gaussian distribution $\mathcal{N}(0,1)$, the distribution of $z_{ij}$ explicitly adheres to a Gaussian distribution $\mathcal{N}(u_{ij},1)$.
-Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the predefined GMM $\{U_{m}|m\in[1,M]\}$ and make the latent image features $b_{ij}$ follow the distribution $N(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM.
-The aforementioned method for modeling the distribution of $z_{ij}$ can be implemented by introducing five additional loss function terms.
+The new derivation method for $z_{ij}$ ensures that, when $b_{ij}$ follows a standard Gaussian distribution $\mathcal{N}(0,1)$, the distribution of $z_{ij}$ explicitly adheres to a Gaussian distribution $\mathcal{N}(u_{ij},1)$.
+Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the constructed GMM $\{U_{m}|m\in[1,M]\}$ and make the latent features $b_{ij}$ follow the distribution $N(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM with trainable means and known variances.
+The aforementioned method for modeling the distribution of $z_{ij}$ can be implemented by introducing five additional loss function terms:
   1. 1.
-    A loss term $\ell_{1}$ aims to encourage the mean of the distribution of $z_{ij}$ to approximate the predefined GMM $\{U_{m}|m\in[1,M]\}$, and its mathematical expression can be formulated as:
+    A loss term $\ell_{1}$ aims to encourage the mean of the distribution of $z_{ij}$ to approximate the constructed GMM $\{U_{m}|m\in[1,M]\}$, and its mathematical expression can be formulated as:
     $\ell_{1}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|u_{ij}-\text{gs}[{U}_{\hat{m}}]\|_{2}^{2}$ (35)
@@
   3. 3.
-    A loss term $\ell_{3}$, which is utilized to train the components within the predefined GMM $\{U_{m}|m\in[1,M]\}$, can be expressed as:
+    A loss term $\ell_{3}$, which is utilized to train the components within the constructed GMM $\{U_{m}|m\in[1,M]\}$, can be expressed as:
     $\ell_{2}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|{U}_{\hat{m}}-\text{gs}[u_{ij}]\|_{2}^{2}$ (37)
@@
   4. 4.
-    A mutual information loss term $\ell_{4}$, which facilitates the training of components within the predefined GMM $\{U_{m}|m\in[1,M]\}$, is expressed as:
+    A mutual information loss term $\ell_{4}$, which facilitates the training of components within the constructed GMM $\{U_{m}|m\in[1,M]\}$, is expressed as:
     $\ell_{4}$ $=\sum_{i=1}^{16}\|x_{i}-\hat{x}_{i}\|_{2}^{2}$ (38)
@@
 After performing a weighted summation of these five loss terms, we obtain the complete loss function $\ell_{\text{WORLD}}$:
-$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+0.25\cdot\ell_{3}+\ell_{4}+\ell_{5}$ (40)
-By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss functions, we can train a DIO model where the distribution of the latent variable $z_{ij}$ is modeled as a GMM.
+$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+\lambda\cdot\ell_{3}+\ell_{4}+\ell_{5}$ (40)
+The coefficient $\lambda$ for $\ell_{3}$ is set to 0.25 in order to balance its relative magnitude compared to $\ell_{1}$.
+By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss functions, we can train a DIO model where the distribution of image features $z_{ij}$ is modeled as a GMM.
 Notably, when training the DIO using $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, the feedforward mechanism undergoes the transformation illustrated in Figure 8.
-Figure 8: Forward Feed Process of DIO Model Incorporating $\ell_{\text{WORLD}}$ as an Additional Loss Term.
+Figure 8: Feedforward Process of DIO Model Incorporating $\ell_{\text{WORLD}}$ as an Additional Loss Term.
 In Figure 8, the latent variable fed into the progressive pattern induction module within DIO is no longer $z_{ij}$, but rather $u_{ij}+\epsilon$.
-In other words, the forward feed process indicated by Equation (IV-C) is updated as follows:
+In other words, the feedforward process indicated by Equation (IV-C) is updated as follows:
 $\{P_{jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}_{i=\alpha}$
 $=\text{PPIM}(\{u_{ij}+\epsilon_{ij},u_{\alpha j}+\epsilon_{\alpha j}|i\in[1,8]\})|_{j=1}^{N}$ (41)
-Where $u_{ij}+\epsilon_{ij}$ can be regarded as a sample drawn from the GMM component $U_{\hat{m}}$ that $z_{ij}$ follows.
+Where $u_{ij}+\epsilon_{ij}$ can be regarded as a sample drawn from the GMM component $U_{\hat{m}}$ that feature $z_{ij}$ follows.
 This design enables DIO to interact with a fully modeled GMM rather than deterministic image features, thereby training the DIO to make decisions based on the complete probabilistic structure of the GMM instead of specific feature instantiations.
@@
 ${\ell_{\text{Weighted}}}=$ $-\sum_{j=1}^{N}\log\frac{{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{\tilde{m}}+\epsilon_{\tilde{m}}|i\in[1,8]\})}{|\tau|}}}}}{\sum_{m=1,m\neq\tilde{m}}^{M}{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{m}+\epsilon_{m}|i\in[1,8]\})}{|\tau|}}}}}$ (42)
-where $\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2}$
-$\epsilon_{m}\sim N(0,1)$
-$\tau$ is a trainable smoothing parameter.
-Given that RPM problems often feature a diverse range of correct options, the rigid approach of designating samples drawn from components that do not correspond to the provided correct option as features of incorrect options carries a potential risk of overgeneralization.
-To mitigate this sample classification bias, we introduce a smoothing parameter $\tau$.
-$\tilde{\text{DIO}}$ refers to the DIO without the image feature extraction module, which is capable of processing features into scores.
-The calculation of $\tilde{m}$ serves to identify the component index in the modeled GMM $\{U_{m}|m\in[1,M]\}$ that corresponds to the feature $z_{\alpha j}$ of the correct option $x_{\alpha}$.
+where $\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2},\,\epsilon_{m}\sim N(0,1)$
+$\tilde{\text{DIO}}$ refers to the DIO without the image feature extraction module, which is capable of processing image features into scores.
+In RPM problems, any RPM image that conforms to the progressive patterns exhibited in the context matrices can be considered a correct option.
+This indicates that RPM problems inherently possess a multi-solution characteristic.
+Consequently, the rigid approach of designating samples drawn from components that do not correspond to the provided correct option $x_{\alpha}$ as features of incorrect options carries a potential risk of overgeneralization.
+To mitigate this sample classification bias, we introduce a trainable smoothing parameter $\tau$.
+$\tilde{m}$ identifies the index of the GMM component that best matches the feature $z_{\alpha j}$ (or $u_{\alpha j}$) of the correct option $x_{\alpha}$.
 The role of $\ell_{\text{Weighted}}$ is to fine-tune the DIO model.
@@
 VIII DIEGO: Methods for Rectifying the Causal Chain in DIO
-This study has made two attempts to refine the learning objectives of the DIO model, resulting in the development of two distinct methodology: Brando and WORLD.
-In this section, we devise a methodology named DIEGO (Direct Intervention for Enforcing Guided Optimal Reasoning), which is designed to rectifying the causal chains modeled by DIO, thereby advancing its reasoning performance.
+This study has made two attempts to refine the learning objectives of the DIO model, resulting in the development of two distinct methodologies: Brando and WORLD.
+In this section, we devise a methodology named DIEGO (Direct Intervention for Enforcing Guided Optimal Reasoning), which is designed to rectify the causal chains modeled by DIO, thereby advancing its reasoning performance.
 VIII-A Mutual Information Excels at Problem-Solving Rather Than Reasoning
-This paper realizes that setting the maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal.
-While mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ directly correlates with accuracy in solving RPM problems, it fails to capture causal relationships inherent in the task design.
-This creates a critical discrepancy: DIO is optimized for high-accuracy RPM solution generation rather than learning the human-predefined reasoning logic.
-Both the Brando and WORLD methodologies aim to assist DIO in better maximizing mutual information, yet they fail to address this discrepancy.
+This paper realizes that setting maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal.
+Although this mutual information measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic, yielding a model optimized for high accuracy rather than for faithful causal reasoning.
+Both the Brando and WORLD methodologies aim to assist DIO in better maximizing mutual information, yet neither resolves this limitation.
+In a nutshell, mutual-information-driven approaches remain effective but must be complemented by an explicit causal perspective.
 Therefore, this paper argues that the causal chain underlying RPM problems should be further modeled.
-The network architecture of DIO strongly drew reference from the causal chain within RPM problems, with the hope that DIO can autonomously model the causal chain during the training process.
-However, this can only ensure that DIO is structurally aligned with the causal chain in RPM problems.
-Whether its latent variables are semantically aligned cannot be guaranteed and remains to be verified.
-In order to verify semantic alignment, this paper selects the “Distribute 9” database under the RAVEN problems[19] as the research object.
-We conduct training on this database using DIO, as well as DIO supplemented with Brando and DIO supplemented with WORLD, respectively.
-Subsequently, we select 30 instances with strictly identical progressive patterns from this database and verify the pairwise similarity of the progressive pattern representations encoded by DIO for these 30 instances.
-The calculated $30\times 30$ similarity matrices have been plotted into heatmaps as shown in Figure 9.
+The network architecture of DIO is designed to mirror this causal chain, aiming for DIO to automatically capture the chain during training.
+Nevertheless, while this design guarantees structural alignment, it cannot ensure that the latent variables derived by DIO are also semantically aligned.
+Given that mutual information fails to account for causality, the semantic alignment awaits empirical verification.
+In order to verify semantic alignment, this paper selects the “Distribute 9” dataset from the RPM-style RAVEN benchmark [19] as the research object.
+We conduct training on this dataset using DIO, as well as DIO supplemented with Brando and WORLD.
+Subsequently, we select 30 instances with strictly identical progressive patterns from this dataset and verify the pairwise similarity of the progressive pattern representations encoded by DIO for these 30 instances.
+The calculated $30\times 30$ similarity matrices are visualized as heatmaps in Figure 9.
 Figure 9: Heatmap of Progressive Pattern Representation Extracted from the Distribute 9 under RAVEN Problems.
-If DIO accurately models the causal chain embedded in RAVEN problems, then after these 30 instances with the same progressive patterns are encoded into progressive pattern representations by DIO, they should demonstrate high cosine similarity among themselves.
-However, the three similarity matrices in Figure 9 do not corroborate this expectation.
-This indicates that, even with auxiliary support from the Brando or WORLD methods, the causal chain modeled by DIO remains systematically misaligned with the ground-truth chains inherent in RPM problems.
-In addition, this further points out the limitations of using mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ maximization as the learning objective.
-Given this, this paper proposes that rectifying the causal chain modeled by DIO during its training process will substantially improve performance in solving RPM problems.
+If DIO accurately models the causal chain embedded in RAVEN problems, then once DIO encodes 30 instances sharing the same progressive patterns into pattern representations, those representations should exhibit high cosine similarity among themselves.
+Contrary to expectations, the three similarity matrices in Figure 9 fail to support this anticipated outcome, indicating that, even with auxiliary support from the Brando or WORLD methods, the causal chain modeled by DIO systematically deviates from the ground-truth causal chains inherent in RPM problems.
+This underscores the limitations of adopting mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective.
+Given these findings, we propose that further modeling of the causal chain could start with semantic rectification in DIO, which would substantially enhance its performance in solving RPM problems.
 VIII-B Rectifying the Causal Chain Modeled by DIO.
 In preliminary experiments, we found that DIO exhibits deviations in modeling the segment of the causal chain that involves “abstract attributes $\rightarrow$ progressive patterns of attributes”.
-This corresponds to a misalign between the results $P_{\alpha{jkl}}$ derived from Formula (IV-C) and human semantics.
-Therefore, this paper proposes that refining the progressive pattern representations $P_{\alpha{jkl}}$ encoded by DIO would be an effective solution.
-This paper has taken note of the metadata provided for each instance in RPM problems, which describes the corresponding instance’s progressive pattern.
-We argue that incorporating this metadata as a supervisory signal into DIO’s training process to supervise the derivation of $P_{\alpha{jkl}}$ can achieve the objective of rectifying the causal chain modeled by DIO.
-Thus, we have designed a method named DIEGO.
-DIEGO involves introducing an additional set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$ for DIO, along with an auxiliary loss function term $\ell_{\text{DIEGO}}$.
-Specifically, the DIEGO method first establishes a set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$, where each component represents a distinct metadata modality.
-Consequently, the reference vector set $\{V_{f}|f\in[1,F]\}$ can serve as a comprehensive dictionary for metadata.
-The loss term $\ell_{\text{DIEGO}}$ serves to align each progressive pattern representation $P_{\alpha{jkl}}$ extracted by DIO with its corresponding reference vector within $\{V_{f}|f\in[1,F]\}$ according to the accompanying metadata.
-The loss function $\ell_{\text{DIEGO}}$ can be expressed as follows:
-${\ell_{\text{DIEGO}}}=-\sum_{\hat{f}=1}^{F}\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{2}\text{Meta}_{{\hat{f}l}}\cdot\log\frac{{{e^{({P_{\alpha jkl}}\cdot{V_{\hat{f}}})/|\tau|}}}}{\sum_{f=1}^{F}{{e^{({P_{\alpha jkl}}\cdot{V_{f}})/|\tau|}}}}$ (43)
-Here, $\text{Meta}_{\hat{f}l}$ denotes the indication status of the $l$-th metadata entry with respect to the $\hat{f}-$th type of progressive pattern.
-Meanwhile, $\tau$ represents a trainable positive scaling coefficient.
-The detailed expression of the loss term $\ell_{\text{DIEGO}}$ demonstrates our methodology for achieving alignment.
-This method utilizes a cross-entropy loss function with a scaling coefficient to control the cosine similarities between $P_{\alpha_{jkl}}$ and all components in the metadata dictionary $\{V_{f}|f\in[1,F]\}$ according to the indication information from the metadata.
-When training the DIO model using $\ell_{\text{DIO}}+\ell_{\text{DIEGO}}$ as the loss function, a solving model that accurately captures the causal chains in RPM problem will be obtained.
-Previously, some studies have also attempted to treat metadata as additional supervision signals and introduce extra progressive pattern learning tasks into their solving models, aiming to simultaneously enhance the model’s interpretability and reasoning accuracy.
-However, research led by MRNet[25, 28, 27, 31] indicates that this approach may backfire.
-Specifically, if metadata is directly incorporated into the training process of the solving models they designed, the model’s reasoning performance will instead decline.
-This may be logically flawed, but it is indeed an established fact.
-Nevertheless, this paper argues that the DIO model constructed based on causal chains, coupled with the DIEGO method capable of strategically correcting key segments of the constructed causal chains, holds the key to breaking through this impasse.
+This corresponds to a misalignment between the outputs $P_{\alpha{jkl}}$ produced by Formula (IV-C) and human semantics.
+Therefore, this paper proposes rectifying the progressive pattern representations $P_{\alpha{jkl}}$ encoded by DIO as an effective remedy.
+This paper notes that each RPM instance is accompanied by metadata describing its progressive patterns that precisely match the semantics predefined for $P_{\alpha{jkl}}$ in the DIO architecture.
+We argue that incorporating this metadata as a supervisory signal into DIO’s training process to guide the derivation of $P_{\alpha{jkl}}$ can rectify the causal chain modeled by DIO.
+To this end, we propose the DIEGO method.
+DIEGO constructs a set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$.
+By assigning each vector a distinct metadata modality, this entire set can be considered as a comprehensive metadata dictionary.
+In addition, DIEGO introduces an auxiliary loss term $\ell_{\text{DIEGO}}$ that serves to align each progressive pattern representation $P_{\alpha{jkl}}$ extracted by the DIO with its corresponding reference vector in $\{V_{f}|f\in[1,F]\}$ based on the accompanying metadata.
+This term is formulated as follows:
+${\ell_{\text{DIEGO}}}=-\sum_{\hat{f}=1}^{F}\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{2}\text{Meta}_{{\hat{f}l}}\cdot\log\frac{{{e^{\frac{\left({P_{\alpha jkl}}\cdot{V_{\hat{f}}}\right)}{|\tau|}}}}}{\sum_{f=1}^{F}{{e^{\frac{\left({P_{\alpha jkl}}\cdot{V_{f}}\right)}{|\tau|}}}}}$ (43)
+Here, $\text{Meta}_{\hat{f}l}$ is a binary indicator that flags whether the $l$-th metadata entry matches the modality associated with the $\hat{f}-$th vector within the reference dictionary $\{V_{f}|f\in[1,F]\}$.
+Meanwhile $\tau$ is a trainable scaling coefficient.
+The detailed formulation of $\ell_{\text{DIEGO}}$ elucidates our alignment mechanism.
+Guided by the metadata-derived indicators, this loss term employs logarithmic smoothing and exponential scaling to regulate the cosine similarities between each $P_{\alpha_{jkl}}$ and all vectors in the reference dictionary.
+Evidently, our objective has shifted from statistical quantity control to vector similarity control, distinguishing DIEGO from Brando and WORLD: it is no longer mutual-information-driven but progressive-pattern-driven.
+When the DIO model is trained with the composite loss $\ell_{\text{DIO}}+\ell_{\text{DIEGO}}$ and the trainable reference dictionary $\{V_{f}|f\in[1,F]\}$, the forward pass of the relevant modules proceeds as illustrated in Figure 10.
+Figure 10: Feedforward Process of DIO Model Incorporating $\ell_{\text{DIEGO}}$ as an Additional Loss Term.
+After training, a solving model that captures causal chains in RPM problems is obtained.
+It is widely accepted that introducing more supervisory signals should improve model performance.
+Indeed, previous studies have attempted to leverage metadata as additional supervision by designing progressive pattern-learning tasks, aiming to boost both interpretability and inference accuracy simultaneously.
+Recent work [25, 28, 27, 31] has collectively shown, however, that this practice can backfire: directly embedding metadata into proposed models actually degrades reasoning performance.
+This paper argues that the DIO model, constructed upon causal chains, together with the DIEGO method that strategically rectifies key segments of those chains, offers a viable way to overcome this limitation.
 IX Solving Generative RPM Problems
-Generative RPM problems represent an advanced variant of RPM tasks, where the solving model is required to generate an ideal answer for the context matrix $\{x_{i}|i\in[1,8]\}$ instead of selecting from a predefined bounded option pool.
-IX-A DIO And WORLD Solve Generative RPM Problems
-The WORLD method not only enhances the performance of DIO but also equips it with the capability to solve generative RPM problems.
-Specifically, the WORLD method introduces an external GMM $\{U_{m}|m\in[1,M]\}$ and a decoder $\mathcal{D}(\cdot)$ for DIO.
-Meanwhile, under the WORLD framework, the loss function term $\ell_{\text{Weighted}}$ defined in Formula (42) not only tightens the variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$ for DIO but also endows DIO with the ability to assign weights to each component in the GMM $\{U_{m}|m\in[1,M]\}$ with respect to the given context matrix features $\{z_{ij}|i\in[1,8],j\in[1,N]\}$.
-The weights assigned to the component $U_{m}$ for the contextual features $\{z_{ij}|i\in[1,8]\}$ can be calculated as follows:
-$S_{{m}{j}}=\tilde{\text{DIO}}(\{z_{i{j}},U_{{m}}|i\in[1,8]\})$ (44)
-Based on all computed weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$, we employ the Gumbel-Softmax trick to draw $N$ components from the GMM $\{U_{m}|m\in[1,M]\}$. This process can be formulated as:
+Generative RPM problems represent an advanced variant of RPM tasks.
+In these problems, the solving models are required to generate at least one solution that conforms to the progressive patterns within the given context matrix, rather than selecting from a predefined bounded option pool, thereby more accurately reflecting the model’s comprehension of the reasoning logic.
+As solution identification remains progressive-pattern-driven, generative RPM problems inherit the multi-solution characteristic from conventional RPM problems.
+Consequently, models are expected to produce multiple solutions that are structurally diverse, which poses significant challenges for most traditional neural networks.
+IX-A WORLD Guides DIO on Generative RPM problems
+Beyond enabling DIO to solve discriminative RPM tasks, the WORLD method also equips it with the capability to tackle generative RPM problems.
+This method constructs a complete generation pipeline for DIO: “estimating solution distribution $\rightarrow$ sampling solution distribution $\rightarrow$ decoding samples”.
+The WORLD method extends DIO by incorporating an external GMM $\{U_{m}|m\in[1,M]\}$, an additional loss term $\ell_{\text{Weighted}}$, and a decoder $\mathcal{D}(\cdot)$.
+Once trained in this framework, DIO can not only discriminate among the options for each context matrix but also directly generate the corresponding solutions.
+Specifically, as shown in Formula (42), the loss term $\ell_{\text{Weighted}}$ prompts DIO to assign a context-dependent weight to each GMM component $U_{m}$, where the weight reflects the likelihood that $U_{m}$ contains the correct options for the given context matrix.
+Within this loss term, the weight for each $U_{m}$ given the context matrix features $\{z_{ij}|i\in[1,8]\}$ is calculated as:
+$S_{mj}=\tilde{\text{DIO}}(\{z_{ij},U_{m}|i\in[1,8]\}).$ (44)
+Since these weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$ indicate the likelihood that the components cover the correct options, we not only gain the chance to sample diverse incorrect options from low-weight components to tighten the lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, but also have the opportunity to combine high-weight components into a solution distribution.
+To this end, we adopt the Gumbel-Max trick [37] to sample $N$ high-weight components as the solution distribution that DIO estimates for the context matrix $\{x_{i}|i\in[1,8]\}$.
+Specifically, this trick is applied as follows:
 $G_{mj}$ $=-\log(-\log(\text{U}(0,1)))$ (45)
@@
 $\hat{m}_{j}$ $=\underset{m}{\text{argmax}}\,\{\tilde{S}_{mj}|m\in[1,M]\}$ (47)
-$\hat{U}_{j}$ $=U_{\hat{m}_{j}}$ (48)
-These $N$ components $\{\hat{U}_{j}|j\in[1,N]\}$ can be regarded as the distribution of solutions estimated by DIO for the context matrix $\{x_{i}|i\in[1,8]\}$.
-The decoder $\mathcal{D}$ is capable of transforming the samples from these $N$ components into the generated answer $\hat{x}$:
+In this case, the sampled indices $\{\hat{m}_{j}|j\in[1,N]\}$ can point out the solution distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$.
+Subsequently, Formula (39) shows that the well-trained decoder $\mathcal{D}(\cdot)$ is capable of transforming samples from this estimated distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$ into the generated answer $\hat{x}$:
 $\hat{x}=\mathcal{D}(\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\})$ (48)
 $\text{where}\quad\epsilon_{ij}\sim N(0,1)$
-Formulas (34), as well as formulas (44) - (49), indicate that the DIO trained by the WORLD method has the capability to estimate the distribution of solution for the context matrix $\{x_{i}\mid i\in[1,8]\}$ and generate RPM images that can serve as solutions.
-IX-B DIEGO Validates the Rationality of Mutual Information Supervision
+In essence, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve the generative RPM problems.
+Notably, WORLD is well-aligned with the multi-solution characteristic of generative RPM problems.
+Formulas (45)–(47) indicate that the Gumbel-Max trick ensures the estimated solution distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$ exhibits diversity.
+Formula (48) further reveals that the solutions decoded from samples $\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\}$ of the estimated distribution have the potential to incorporate the sampling diversity $\epsilon_{ij}$.
+By iterating the “estimating $\rightarrow$ sampling $\rightarrow$ decoding” pipeline, which can generate diverse solutions, WORLD systematically explores multiple potential solutions and fully embraces the multi-solution logic of generative RPM tasks.
+IX-B Evaluating the Generated Solutions
 Verifying the solution accuracy of generative RPM problems has long been challenging.
-RPM problems often have multiple valid solutions, which means that the generative solving models cannot be evaluated simply by comparing each generated solution with the given correct option.
+Generative RPM problems possess the multi-solution characteristic, which means that the generative solving models cannot be evaluated simply by comparing each generated solution with the provided correct option.
 This also makes it challenging for RPM experts to devise a perfect verification algorithm.
-Most prior work on generative RPM problems has used self-verification, i.e., using their own discriminative models to validate their own generative models.
-However, this paper argues that this paradigm of self-verification is inadequate.
-The cause of the multiple-solution characteristic in RPM problems lies in the fact that solution identification is driven by a progressive pattern.
-Any RPM image that adheres to the progressive pattern exhibited in the contextual matrices can be considered a valid solution.
-This implies that algorithms or models used to verify the solution accuracy of generative RPM problems should also be progressive-pattern-driven.
+Most prior work on generative RPM problems has used self-verification, i.e., using their own discriminative models to evaluate their own generative models.
+However, this paper argues that this self-verification paradigm is inadequate.
+In generative RPM problems, the multi-solution characteristic stems from the progressive-pattern-driven solution evaluation approach.
+This implies that the algorithms or models used to validate the accuracy of solutions for such problems should also be progressive-pattern-driven rather than mutual-information-driven.
 The DIEGO approach meets this expectation, as the DIO model trained solely with the loss term $\ell_{\text{DIEGO}}$ will be a progressive-pattern-driven solving model.
-Additionally, DIO+WORLD is a strong mutual-information-based inductive model.
-Using the progressive-pattern-driven model DIEGO to verify DIO+WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also tests the deeper potential of mutual-information-driven solution models.
+More importantly, WORLD is a strong mutual-information-driven inductive model.
+Using the progressive-pattern-driven model DIEGO to verify WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also examines the rationality and potential of mutual information as a reasoning cue.
 X Experiment
-All experiments conducted in this paper were programmed in Python, utilizing the PyTorch[37] framework.
-This paper introduces a comprehensive framework for solving RPM problems through three distinct configurations that integrate a purpose-built deep learning model with specialized training methodologies.
-The proposed configurations leverage:
-  1. 1.
-    Solo DIO, which solves RPM problems through a structured causal chain modeling approach without additional optimization strategies.
+All experiments conducted in this paper were programmed in Python based on the PyTorch[38] framework. This paper proposes four distinct configurations for RPM problems:
+X-1 Solo DIO
+which solves RPM problems through a structured causal chain modeling approach without additional optimization strategies.
 In the DIO model, the hyperparameter $N$ is set to 16, indicating that the ViT within its image feature extraction module employs a $4\times 4$ patch division strategy;
-  2. 2.
-    Brando-enhanced DIO, which employs trainable hypothetical incorrect options to tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
-In the Brando framework, the hyperparameter $\beta$ is set to 24, which means the Brando network provides 88 ($\beta-17=8$) constructive hypothetical incorrect options to the DIO in each iteration;
-  3. 3.
-    WORLD-augmented DIO, utilizing a Gaussian Mixture Model for feature distribution modeling to provide an infinite set of weighted incorrect options for tighting the lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
+X-2 Brando-enhanced DIO
+which maps out constructive hypothetical incorrect options to tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
+In the Brando framework, the hyperparameter $\beta$ is set to 24, which means the Brando network provides 88 ($\beta-16=8$) constructive hypothetical incorrect options to the DIO in each iteration;
+X-3 WORLD-augmented DIO
+utilizing a Gaussian Mixture Model for feature distribution modeling to provide an infinite set of weighted incorrect options for tightening the lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
 In the WORLD framework, the hyperparameter $M$ is set to 256, equipping the GMM employed for DIO with 256 components.
-This configuration enables the generation of $M^{N}=256^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options.
-  4. 4.
-    DIEGO-guided DIO, incorporating metadata supervision to directly rectify semantic biases in the causal chain of abstract attribute patterns.
-In the DIEGO framework, the parameter $F$ is determined by the number of metadata modality in the targeted RPM problem, rather than being arbitrarily set.
+This configuration enables the generation of $M^{N}=256^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options;
+X-4 DIEGO-guided DIO
+incorporating metadata supervision to directly rectify semantic biases in the causal chain of abstract attribute patterns.
+In the DIEGO framework, the parameter $F$ is determined by the number of metadata modality in the targeted RPM problem.
 This paper conducted experiments on the RAVEN [19], I-RAVEN[34] and PGM [20] datasets, centered around these four configurations.
@@
 DIO+DIEGO 99.9/99.9 100.0/100.0 99.9/99.9 99.7/99.8 100.0/100.00 100.0/100.0 100.0/100.0 99.8/99.8
-Experimental results recorded in Table I indicate that all four model configurations proposed in this study outperform all existing baselines on both the RAVEN and I-RAVEN datasets.
+These results indicate that all four model configurations proposed in this study outperform existing baselines on both the RAVEN and I-RAVEN datasets.
 Notably, the DIO+DIEGO configuration achieves a peak accuracy of 99.9%, delivering significant performance gains over prior state-of-the-art models (Triple-CFN[33] and RS-TRAN[31]).
@@
 X-B Experiment on PGM Problem
-The performance comparison among the four model configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset is summarized in Table II.
-TABLE II: Reasoning accuracies of models on PGM-neutral.
+The performance comparison among the four DIO configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset is summarized in Table II.
+TABLE II: Reasoning Accuracies on PGM-neutral.
 Model Test Accuracy(%)
@@
 The paper also conducted experiments on the o.o.d. sub-task of PGM, with the results documented in Table III.
-TABLE III: Generalization Results on PGM.
+TABLE III: Reasoning Accuracies on PGM Generalization Tasks.
 Dataset
@@
 DIO+DIEGO
-92.5 13.8 14.4 27.2 28.8 98.1 98.1
+93.2 13.8 14.4 27.2 28.8 98.3 98.3
 The results recorded in Table III reveal that while mutual information tightening and causal chain modeling yield comparable performance gains for DIO on i.i.d.
@@
 $\times2$ Grid entire:99.9
-$\times3$ Grid entire:99.6
+$\times3$ Grid entire:99.8
 L-R left:99.9 right:99.9
@@
 $\times2$ Grid entire:99.9
-$\times3$ Grid entire:99.6
+$\times3$ Grid entire:99.8
 L-R left:99.9 right:99.9
@@
 PGM neutral shape:99.7 line:99.9
-Table IV indicates that the DIO+DIEGO demonstrates remarkably high accuracy in progressive pattern reasoning for three categories of RPM problems, making it a suitable validation model for generative RPM tasks.
+Table IV indicates that the DIO+DIEGO demonstrates remarkably high accuracy in progressive pattern reasoning, making it a suitable validation model for generative RPM tasks.
 Consequently, we evaluate the DIO+WORLD model using DIO+DIEGO and report its accuracy on generative RAVEN, I-RAVEN and PGM-neutral problems in Table V.
 TABLE V: Generation Accuracies of DIO+WORLD
-Dataset Sub-task Accuracy (%)
+Dataset Sub-task Accuracy (%) Average (%)
 RAVEN/I-RAVEN
-Center 91.0/91.2
-$\times2$ Grid 85.2/85.0
+Center 98.0/98.2
+91.3/92.6
+$\times2$ Grid 82.2/82.0
 $\times3$ Grid 81.5/81.8
-L-R 87.9/88.5
-U-D 87.9/88.4
-O-IC 89.0/89.3
-O-IG 83.0/84.2
-PGM Neutral 84.5
-Table V clearly demonstrates that, without relying on any additional supervisory signals, DIO+WORLD has achieved highly competitive results in open-ended generative RPM problems, which strongly confirms the significant research value of the DIO model.
-In addition, these results also demonstrate that although mutual information does not directly indicate causal relationships, it remains highly valuable in constructing reasoning capabilities and can withstand the tests conducted by models driven by causal chains.
+L-R 96.5/96.9
+U-D 95.7/96.9
+O-IC 94.9/96.5
+O-IG 90.5/96.2
+PGM Neutral 80.5 80.5
+Table V shows that without extra supervisory signals, DIO+WORLD has attained highly competitive results in open-ended generative RPM problems, strongly confirming the significant research value of the WORLD method.
+Moreover, these results indicate that though mutual information doesn’t directly reveal causal relationships, it’s still highly valuable for building reasoning capabilities and can withstand the tests by progressive-pattern-driven models.
+Notably, the results in Table V still have the potential to improve with further training.
+In the WORLD method, the ability to tighten mutual information emerges in the early training epochs, whereas the generation capability matures only after the loss term $\ell_{\text{Weighted}}$ stabilizes at a low value, necessitating additional training epochs.
+To avoid excessive resource consumption, training was halted once the results adequately demonstrated WORLD’s effectiveness.
 XI Ablation Study
-In this section, this paper conducts a series of ablation experiments. The ablation experiments will be conducted focusing on the key parameter settings.
+This section presents a series of ablation experiments focusing on key parameter settings and network architecture.
 XI-A Ablation Study on DIO
-The ablation study on DIO focuses on the design of the network’s feed-forward structure.
-Within DIO’s progressive pattern induction module, a permutation and combination strategy ($C_{n}^{r}$) is employed to extract progressive patterns from row and column features.
-This approach aims to model the causal chain details in RPM problems.
-In this ablation experiment, we will remove the $C_{n}^{r}$ operation.
-Instead, we will directly input the total of six row-column features for progressive pattern extraction.
-By doing so, we aim to obscure the structure in DIO that aligns with the causal chain, thereby validating the significance of causal chain modeling as a methodology for network construction.
-The obscured DIO is denoted as DIO.
-The reasoning accuracy achieved by DIO on key sub-tasks of RAVEN, I-RAVEN, and PGM is recorded in Table VI.
-TABLE VI: Reasoning Accuracies of DIO.
+The ablation study on DIO centers on the design of the network’s feed-forward structure.
+Within its progressive-pattern induction module, DIO employs a permutation and combination strategy ($C_{n}^{r}$) to extract progressive patterns from row and column features, aiming to model the causal-chain details inherent in RPM problems.
+For ablation, we remove the ($C_{n}^{r}$) operation and instead feed all six row-column features directly into the progressive-pattern extractor.
+By masking the structure that aligns with the causal chain, we assess the importance of causal-chain modeling in network design.
+The resulting accuracies on key sub-tasks of RAVEN, I-RAVEN and PGM are reported in Table VI.
+TABLE VI: Reasoning Accuracies of masked DIO.
 Dataset Sub-task Accuracy (%)
@@
 PGM Neutral 97.2
-Table VI indicates that constructing RPM solving models based on causal chains is a critically important condition for achieving high reasoning performance.
+Table VI shows that building RPM-solving models that explicitly incorporate causal chains is crucial for superior reasoning performance.
 XI-B Ablation Study on Brando
 The ablation study on Brando focuses on the hyperparameter $\beta$.
-This paper demonstrates that the number of constructive incorrect options provided to DIO directly influences the tightness of the variational lower bound on mutual information derived by DIO.
-Therefore, it is essential to conduct ablation experiments focusing on the number ($\beta$-17) of hypothetical options provided by the Brando network.
-The reasoning accuracies of DIO+Brando on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $\beta$ values are recorded in Table VII.
-TABLE VII: Reasoning Accuracy of DIO+Brando under Different $\beta$ Values.
+$\beta$ denotes the number ($\beta-17$) of hypothetical options the Brando network provides to DIO.
+As this paper shows that the count of constructive incorrect options for DIO directly influences the tightness of its derived mutual information variational lower bound, which in turn directly impacts DIO’s reasoning potential.
+Therefore, this paper conducts ablation experiments on the DIO + Brando model with different $\beta$ values across key sub-tasks of RAVEN, I-RAVEN, and PGM, and records the results in the Table VII.
+TABLE VII: Reasoning Accuracies of DIO+Brando under Different $\beta$ Values.
 Dataset Sub-task Accuracy(%)
-$\beta$=24 $\beta$=22 $\beta$=20 $\beta$=18 No $\beta$
+$\beta$=24(8) $\beta$=22(6) $\beta$=20(4) $\beta$=18(2) Solo DIO
 RAVEN/
@@
 O-IG 98.6 / 98.9 98.2 / 98.6 97.6 / 98.3 97.5 / 98.3 96.7 / 97.9
+$\times3$ Grid 97.9 / 98.4 97.6 / 98.0 96.9 / 97.8 96.4 / 97.7 96.1 / 97.4
+O-IG 98.0 / 98.6 97.5 / 98.3 97.2 / 98.2 97.0 / 98.2 96.7 / 97.9
 PGM Neutral 99.0 98.8 98.8 98.5 98.1
-Table VII indicates that even when only a small number of constructive hypothetical options are introduced into DIO, significant performance improvements can already be achieved for DIO.
+Neutral 98.7 98.4 98.4 98.3 98.1
+Table VII indicates that even when only two constructive hypothetical options are introduced into DIO, significant performance improvements can be achieved.
+The asterisk (*) on a sub-task indicates that, when the Brando network is applied to this task, the process of learning historical image features defined by Equation (27) is deactivated, facilitating ablation and analysis of this mechanism.
+The asterisk-marked results demonstrate that learning historical image features is effective in mapping out constructive hypothetical options.
 XI-C Ablation Study on WORLD
@@
 The reasoning accuracies of DIO+WORLD on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $M$ values are recorded in Table VIII.
-TABLE VIII: Reasoning Accuracy of DIO+WORLD under Different $M$ Values.
+TABLE VIII: Reasoning Accuracies of DIO+WORLD under Different $M$ Values.
 Dataset Sub-task Accuracy(%)
-MM=256 MM=128 MM=64 MM=32 No $M$
+MM=256 MM=128 MM=64 MM=32 Solo DIO
 RAVEN/
@@
 The number of components directly affects the complexity of the GMM.
-As shown in Table VIII, when the number of GMM components is relatively small ($M=32$), meaning that WORLD models a distribution with lower complexity, it will have a negative impact on the performance of DIO in the PGM task.
-This study posits that the effectiveness of the WORLD method depends on whether the GMM it constructs can effectively represent the inherent complexity of the natural distribution of features encoded by DIO for RPM images.
-Specifically, when $M$ is set to 32, this GMM with a low number of components cannot adequately capture the complexity of the natural distribution of PGM image features.
-Forcing the distribution of PGM image features with higher complexity into this simplified GMM will inevitably lead to a performance degradation of DIO.
-WORLD introduces an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM.
-As a highly versatile distribution model, the GMM theoretically ensures that the WORLD method has a considerable scope of effectiveness.
-However, the phenomenon where WORLD’s modeling of a GMM with a low number of components causes performance degradation in DIO imposes a crucial prerequisite for the validity of WORLD: the GMM modeled by WORLD must be capable of accommodating the natural complexity of the target problem.
-Certainly, in traditional deep learning tasks, underfitting is a prevalent and expected phenomenon that arises when neural networks lack sufficient capacity to handle the complexity of problems.
-However, the WORLD method makes this phenomenon manageable through its $M$-dependent capacity control.
-Therefore, particular attention must be paid to the parameter $M$ when applying and adjusting this method.
+WORLD proposes an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM.
+As a highly versatile distribution model, the GMM can theoretically ensure WORLD’s broad applicability.
+However, as shown in Table VIII, when the number of components is relatively small ($M=32$), WORLD constructs a less complex GMM, which negatively impacts DIO’s performance in the PGM task.
+This paper contends that this phenomenon arises because a 32-component GMM fails to capture the complexity of the underlying distribution of PGM image features.
+If DIO attempts to fit these more complex features into such a simplified GMM, a performance decline can be anticipated.
+Therefore, this study establishes a critical prerequisite for WORLD’s applicability: the GMM modeled by WORLD must be able to adapt to the inherent complexity of the target problem.
+In traditional deep learning tasks, underfitting is a prevalent phenomenon arising when neural networks lack sufficient capacity.
+However, WORLD makes this phenomenon manageable through its $M$-dependent capacity control.
+Therefore, particular attention is required to the parameter $M$ when applying this method.
 XII Conclusion
 This study has constructed a complete abstract reasoning solution through systematic innovations, with four core innovations forming a progressive technical system that achieves significant progress in both theoretical breakthroughs and empirical performance:
-  1. 1.
-    Causal Chain Modeling
-    • First established a complete causal chain analysis framework of “Image $\rightarrow$ Abstract Attributes $\rightarrow$ Progressive Patterns $\rightarrow$ Pattern Consistency $\rightarrow$ Answer”
-    • Provided an explainable causal relationship modeling paradigm for deep learning models.
-  2. 2.
-    Brando Error Mapping Mechanism
-    • Developed a trainable hypothetical incorrect option mapping mechanism to explicitly tighten the variational lower bound of mutual information.
-    • Creatively introduced mutual information tightening concepts into abstract reasoning frameworks, thereby opening novel avenues for variational inference methodologies.
-  3. 3.
-    WORLD Feature Modeling
-    • Constructed a Gaussian Mixture Model (GMM) for RPM image feature distributions, enabling infinite and semantically-rich incorrect option sampling
-    • Empowering DIO to solve open-ended generative RPM problems.
-  4. 4.
-    DIEGO Causal Correction Innovation
-    • Introduced external metadata to construct supervision signals, precisely correcting semantic biases in the causal chain modeled by DIO
-    • Established mappings between model reasoning chains and ground-truth causal chains, aligning internal representations with human cognitive frameworks
-[…]
-[…]
-This study’s multi-dimensional innovations not only achieve technical breakthroughs but, more importantly, provide a new paradigm for causal explainability research in deep learning models, holding significant theoretical value and practical implications for advancing artificial intelligence toward higher-order cognitive intelligence.
+XII-1 DIO
+We propose that structuring neural network architectures around the causal chains inherent in reasoning problems can lead to enhanced reasoning performance.
+We also suggest that the learning objectives of end-to-end solving models, including DIO, are essentially variational lower bounds of the mutual information between the context matrix and the correct option.
+Consequently, both performance and learning efficacy are directly impacted by the looseness of these lower bounds.
+XII-2 Brando
+We have integrated the concept of mutual information tightening into abstract reasoning frameworks, showing that training with more constructive incorrect options serves as a key mechanism for bounding mutual information.
+This explores novel pathways for variational inference methodologies.
+Concurrently, we have developed a constructive hypothetical incorrect option mapping mechanism to analytically constrain the variational lower bound of mutual information.
+XII-3 WORLD
+We constructed a GMM for the distribution of RPM image features, thereby enabling targeted sampling of infinite and semantically rich options.
+This approach not only assists the DIO model in tightening mutual information but also equips it with the capability to solve open-ended generative RPM problems.
+XII-4 DIEGO
+We have introduced external metadata to construct supervision signals, precisely correcting semantic biases in the causal chain modeled by DIO.
+We have established mappings between model reasoning chains and ground-truth causal chains, aligning internal representations with human cognitive frameworks.
+Combining these four innovations yields a complete technical closed loop of “Architecture Design $\rightarrow$ Training Optimization $\rightarrow$ Data Augmentation $\rightarrow$ Causal Rectification”.

### v4

@@
 In the field of natural language processing, language models that are built on deep learning can automatically learn language grammar, semantics, and contextual associations, performing exceptionally well in tasks such as machine translation and text generation, and propelling the technology towards intelligence and automation [4, 8, 9].
-In the multimodal cross-field of visual question answering systems, deep learning, which integrates visual and language information, can achieve joint understanding and reasoning, accurately answering natural language questions that are related to images [11, 12].
+In the multimodal cross-field of visual question answering systems, deep learning, which integrates visual and language information, can achieve joint understanding and reasoning, accurately answering natural language questions that are related to images [11].
 Currently, research on Large Language Models (LLMs) is a focal point in the field of deep learning.
@@
 II-B Related work
-Image reasoning models such as CoPINet[23], LEN+teacher[22], and DCNet[24] primarily focus on learning differences and rules.
+Image reasoning models such as CoPINet[21] and DCNet[22] primarily focus on learning differences and rules.
 Meanwhile, models like NCD[24], SAVIR-T[25], as well as neuro-symbolic systems (including PrAE, NVSA, ALANS[26, 27, 28]), have achieved promising results in both enhancing model interpretability and improving reasoning accuracy.
@@
 However, this paper finds that DIO is in satin without substance.
-Its learning approach is essentially the same as that of the previous RPM solving model, whose purpose is to maximize the mutual information between the context matrix and the correct option in RPM instance.
+Its learning approach is essentially the same as that of the previous RPM-solving model, whose purpose is to maximize the mutual information between the context matrix and the correct option in the RPM instance.
 This learning approach leads to a situation where, no matter how well its structure aligns with the causal chains in the problems, its learning and optimization will not progress towards the human predefined reasoning framework.
@@
 IV DIO: A Baseline Model
-In this section, we design a high-performance RPM solving model, named DIO (Deep-learning Intelligent-model Organized via Causal Chains).
+In this section, we design a high-performance RPM-solving model, named DIO (Deep-learning Intelligent-model Organized via Causal Chains).
 While providing a detailed exposition of the model’s structure, we have also elucidated its underlying design principles.
@@
 This implies that there exists a causal chain in RPM problems: RPM images $\rightarrow$ abstract attributes $\rightarrow$ progressive patterns of attributes $\rightarrow$ consistency of progressive patterns $\rightarrow$ correctness of options.
-This paper proposes that designing the structure of an RPM solving model in accordance with this causal chain will yield high performance, which prompted us to design the DIO model.
+This paper proposes that designing the structure of an RPM-solving model in accordance with this causal chain will yield high performance, which prompted us to design the DIO model.
 Accordingly, the DIO comprises four modules: the image feature extraction module, the progressive pattern induction module, the progressive pattern consistency evaluation module, and the option evaluation module.
@@
 where $\mathbb{E}[\cdot]$ denotes the expectation over the data distribution.
-V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM Solving Model
+V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM-solving Model
 This paper proposes that causal-chain-based architecture optimization can effectively enhance the performance of deep learning networks in solving RPM problems.
@@
 The right-hand side of Inequality (14) is precisely the classical formulation for optimizing the probabilistic distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the Cross-Entropy (CE) loss function.
-From this, it can be inferred that for those past end-to-end RPM solving models that parameterize $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ and employ the CE loss function, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
-Equation (13) and Inequality (14) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
+From this, it can be inferred that for those past end-to-end RPM-solving models that parameterize $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ and employ the CE loss function, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
+Equation (13) and Inequality (14) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM-solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
 In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
@@
 In other words, $\ell_{\text{down}}$ ensures hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ outscore the original incorrect options $\{x_{i}\mid i\in[9,16],i\neq\alpha\}$ under DIO.
-In the loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive options.
-When optimizing the DIO using $\ell_{\text{Brando}}$ as the loss function, it can induce bounded hypothetical options while leveraging a tighter variational lower bound to enable DIO to maximize the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.
+In the loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive ones.
+By adopting $\ell_{\text{Brando}}$ as the training loss function, DIO not only obtains constructive hypothetical options but also parameterizes a tighter variational lower bound, thereby maximizing the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.
 VI-B Map Out Hypothetical Options
-Formulas (18) and (19) indicate that hypothetical options should demonstrate constructiveness for any given context matrix, meaning that hypothetical options should be customized according to the given context matrix.
-Therefore, training a set of shared hypothetical options is inappropriate.
-This implies that we require a mapping with the following form:
+Formulas (18) and (19) indicate that hypothetical options should demonstrate constructiveness for any given context matrix, meaning that such options should be customized according to the provided context matrix.
+This implies that naively training a set of shared hypothetical options across all context matrices is suboptimal.
+Given this, we require a mapping with the following form:
 $\{x_{i}|i\in[17,\beta]\}=\mathcal{F}(\{x_{i}|i\in[1,8]\})$ (20)
@@
 This adjustment enables the introduction of supplementary incorrect options during DIO’s late-stage training, thereby steering its learning objective toward a tighter variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, ultimately leading to further performance enhancement.
-VIII DIEGO: Methods for Rectifying the Causal Chain in DIO
+VIII DIEGO: Methods for Rectifying the Causal Chain modeled by DIO
 This study has made two attempts to refine the learning objectives of the DIO model, resulting in the development of two distinct methodologies: Brando and WORLD.
 In this section, we devise a methodology named DIEGO (Direct Intervention for Enforcing Guided Optimal Reasoning), which is designed to rectify the causal chains modeled by DIO, thereby advancing its reasoning performance.
-VIII-A Mutual Information Excels at Problem-Solving Rather Than Reasoning
+VIII-A Target Reasoning Rather Than Solving Problems
 This paper realizes that setting maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal.
-Although this mutual information measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic, yielding a model optimized for high accuracy rather than for faithful causal reasoning.
-Both the Brando and WORLD methodologies aim to assist DIO in better maximizing mutual information, yet neither resolves this limitation.
-In a nutshell, mutual-information-driven approaches remain effective but must be complemented by an explicit causal perspective.
+Although this statistical measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic [35], yielding a model optimized for high accuracy rather than for faithful causal reasoning.
+Thus, the performance gains brought by mutual information may be superficial.
+Both the Brando and WORLD methodologies aim to assist DIO in maximizing mutual information, yet neither resolves this limitation.
+In a nutshell, mutual-information-driven approaches are effective but need an explicit causal perspective.
 Therefore, this paper argues that the causal chain underlying RPM problems should be further modeled.
@@
 Nevertheless, while this design guarantees structural alignment, it cannot ensure that the latent variables derived by DIO are also semantically aligned.
-Given that mutual information fails to account for causality, the semantic alignment awaits empirical verification.
-In order to verify semantic alignment, this paper selects the “Distribute 9” dataset from the RPM-style RAVEN benchmark [19] as the research object.
+Given that mutual information fails to account for causality [35], the semantic alignment awaits empirical verification.
+In order to verify semantic alignment, this paper takes the “Distribute 9” dataset from the RPM-style RAVEN benchmark [18] as the research object.
 We conduct training on this dataset using DIO, as well as DIO supplemented with Brando and WORLD.
-Subsequently, we select 30 instances with strictly identical progressive patterns from this dataset and verify the pairwise similarity of the progressive pattern representations encoded by DIO for these 30 instances.
+Subsequently, we extract 30 instances that share strictly identical progressive patterns and compute the pairwise similarities among the progressive pattern representations that DIO encodes for these 30 instances.
 The calculated $30\times 30$ similarity matrices are visualized as heatmaps in Figure 9.
@@
 If DIO accurately models the causal chain embedded in RAVEN problems, then once DIO encodes 30 instances sharing the same progressive patterns into pattern representations, those representations should exhibit high cosine similarity among themselves.
-Contrary to expectations, the three similarity matrices in Figure 9 fail to support this anticipated outcome, indicating that, even with auxiliary support from the Brando or WORLD methods, the causal chain modeled by DIO systematically deviates from the ground-truth causal chains inherent in RPM problems.
-This underscores the limitations of adopting mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective.
-Given these findings, we propose that further modeling of the causal chain could start with semantic rectification in DIO, which would substantially enhance its performance in solving RPM problems.
+Contrary to expectations, the three similarity matrices in Figure 9 fail to support this anticipated outcome.
+This finding indicates that, even with auxiliary supervision from Brando or WORLD, the causal chain learned by DIO systematically diverges from the ground-truth chain inherent in RPM problems.
+This further underscores the limitations of adopting mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective.
+Therefore, we propose that further modeling of the causal chain could prioritize the rectification of DIO’s semantic representations, which would substantially enhance its performance on RPM problems.
 VIII-B Rectifying the Causal Chain Modeled by DIO.
@@
 It is widely accepted that introducing more supervisory signals should improve model performance.
-Indeed, previous studies have attempted to leverage metadata as additional supervision by designing progressive pattern-learning tasks, aiming to boost both interpretability and inference accuracy simultaneously.
+Indeed, previous studies have attempted to leverage metadata as additional supervision, aiming to boost both interpretability and reasoning accuracy simultaneously.
 Recent work [23, 26, 25, 29] has collectively shown, however, that this practice can backfire: directly embedding metadata into proposed models actually degrades reasoning performance.
@@
 $\text{where}\quad\epsilon_{ij}\sim N(0,1)$
-In essence, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve the generative RPM problems.
+In a nutshell, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve the generative RPM problems.
 Notably, WORLD is well-aligned with the multi-solution characteristic of generative RPM problems.
@@
 The DIEGO approach meets this expectation, as the DIO model trained solely with the loss term $\ell_{\text{DIEGO}}$ will be a progressive-pattern-driven solving model.
-More importantly, WORLD is a strong mutual-information-driven inductive model.
-Using the progressive-pattern-driven model DIEGO to verify WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also examines the rationality and potential of mutual information as a reasoning cue.
+More importantly, WORLD is a strong mutual-information-driven inductive model, yet it cannot escape the inherent drawback of neglecting reasoning logic.
+Using the progressive-pattern-driven model DIEGO to verify WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also examines the extent to which WORLD’s performance gains are merely superficial.
 X Experiment
@@
 X-1 Solo DIO
-which solves RPM problems through a structured causal chain modeling approach without additional optimization strategies.
 In the DIO model, the hyperparameter $N$ is set to 16, indicating that the ViT within its image feature extraction module employs a $4\times 4$ patch division strategy;
 X-2 Brando-enhanced DIO
-which maps out constructive hypothetical incorrect options to tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
 In the Brando framework, the hyperparameter $\beta$ is set to 24, which means the Brando network provides 88 ($\beta-16=8$) constructive hypothetical incorrect options to the DIO in each iteration;
 X-3 WORLD-augmented DIO
-utilizing a Gaussian Mixture Model for feature distribution modeling to provide an infinite set of weighted incorrect options for tightening the lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
-In the WORLD framework, the hyperparameter $M$ is set to 256, equipping the GMM employed for DIO with 256 components.
-This configuration enables the generation of $M^{N}=256^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options;
+In the WORLD framework, the hyperparameter $M$ is set to 512, equipping the GMM employed for DIO with 512 components.
+This configuration enables the generation of $M^{N}=512^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options.
+It is worth noting that when WORLD is used to assist DIO in solving generative RPM problems, $M$ is set to 2048;
 X-4 DIEGO-guided DIO
-incorporating metadata supervision to directly rectify semantic biases in the causal chain of abstract attribute patterns.
-In the DIEGO framework, the parameter $F$ is determined by the number of metadata modality in the targeted RPM problem.
+In the DIEGO framework, the parameter $F$ is determined by the number of metadata modalities in the targeted RPM problem.
 This paper conducted experiments on the RAVEN [18], I-RAVEN[32] and PGM [19] datasets, centered around these four configurations.
@@
 DIO+Brando 99.0
-DIO+WORLD 99.2
+DIO+WORLD 99.3
 DIO+DIEGO 99.6
@@
 The paper also conducted experiments on the o.o.d. sub-task of PGM, with the results documented in Table III.
-TABLE III: Reasoning Accuracies on PGM Generalization Tasks.
+TABLE III: Accuracies on PGM Generalization Tasks.
 Dataset
@@
 Table VI shows that building RPM-solving models that explicitly incorporate causal chains is crucial for superior reasoning performance.
+Moreover, across all RAVEN and PGM sub-tasks, DIO uniformly splits every image into $4\times 4$ patches.
+This design deliberately bypasses the grid-like entity layouts in RAVEN’s $2\times 2$ Grid and $3\times 3$ Grid sub-tasks, preventing unintended inductive biases and prior injection while embracing the causal chain rationale of RPM problems: multiple entities must be analyzed holistically.
+We therefore conducted ablation experiments on the hyperparameter $N$ to examine how aligning the patch grid with RAVEN’s $2\times 2$ Grid and $3\times 3$ Grid sub-tasks affects performance, and the results are shown in Table VII.
+The results indicate that favoring priors over the causal chain may undermine model performance.
+TABLE VII: The reasoning accuracy of DIO when its patch grid is aligned with entity positions in the “2×2 Grid” and “3×3 Grid” sub-tasks of RAVEN and I-RAVEN.
+Accuracy(%)
+Patch grid 2 $\times$ 2 Grid 3 $\times$ 3 Grid
+Aligned with entity 95.0/98.3 92.8/94.2
+$4\times 4$ Grid 99.6/99.7 96.1/97.4
 XI-B Ablation Study on Brando
@@
 As this paper shows that the count of constructive incorrect options for DIO directly influences the tightness of its derived mutual information variational lower bound, which in turn directly impacts DIO’s reasoning potential.
-Therefore, this paper conducts ablation experiments on the DIO + Brando model with different $\beta$ values across key sub-tasks of RAVEN, I-RAVEN, and PGM, and records the results in the Table VII.
-TABLE VII: Reasoning Accuracies of DIO+Brando under Different $\beta$ Values.
+Therefore, this paper conducts ablation experiments on the DIO + Brando model with different $\beta$ values across key sub-tasks of RAVEN, I-RAVEN, and PGM, and records the results in the Table VIII.
+TABLE VIII: Accuracies of DIO+Brando under Different $\beta$.
 Dataset Sub-task Accuracy(%)
@@
 Neutral 98.7 98.4 98.4 98.3 98.1
-Table VII indicates that even when only two constructive hypothetical options are introduced into DIO, significant performance improvements can be achieved.
+Table VIII indicates that even when only two constructive hypothetical options are introduced into DIO, significant performance improvements can be achieved.
 The asterisk (*) on a sub-task indicates that, when the Brando network is applied to this task, the process of learning historical image features defined by Equation (27) is deactivated, facilitating ablation and analysis of this mechanism.
@@
 Therefore, ablation experiments centered around the value of $M$ are warranted.
-The reasoning accuracies of DIO+WORLD on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $M$ values are recorded in Table VIII.
-TABLE VIII: Reasoning Accuracies of DIO+WORLD under Different $M$ Values.
+The reasoning accuracies of DIO+WORLD on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $M$ values are recorded in Table IX.
+TABLE IX: Accuracies of DIO+WORLD under Different $M$.
 Dataset Sub-task Accuracy(%)
-MM=256 MM=128 MM=64 MM=32 Solo DIO
+MM=512 MM=256 MM=128 MM=32 Solo DIO
 RAVEN/
@@
 O-IG 99.0 / 99.5 98.6 / 99.0 98.0 / 98.7 97.1 / 98.0 96.7 / 97.9
-PGM Neutral 99.2 99.0 98.5 98.0 98.1
+PGM Neutral 99.3 99.0 98.5 98.0 98.1
 The number of components directly affects the complexity of the GMM.
@@
 As a highly versatile distribution model, the GMM can theoretically ensure WORLD’s broad applicability.
-However, as shown in Table VIII, when the number of components is relatively small ($M=32$), WORLD constructs a less complex GMM, which negatively impacts DIO’s performance in the PGM task.
+However, as shown in Table IX, when the number of components is relatively small ($M=32$), WORLD constructs a less complex GMM, which negatively impacts DIO’s performance in the PGM task.
 This paper contends that this phenomenon arises because a 32-component GMM fails to capture the complexity of the underlying distribution of PGM image features.
@@
 Therefore, this study establishes a critical prerequisite for WORLD’s applicability: the GMM modeled by WORLD must be able to adapt to the inherent complexity of the target problem.
-In traditional deep learning tasks, underfitting is a prevalent phenomenon arising when neural networks lack sufficient capacity.
-However, WORLD makes this phenomenon manageable through its $M$-dependent capacity control.
-Therefore, particular attention is required to the parameter $M$ when applying this method.
 XII Conclusion
@@
 XII-1 DIO
-We propose that structuring neural network architectures around the causal chains inherent in reasoning problems can lead to enhanced reasoning performance.
-We also suggest that the learning objectives of end-to-end solving models, including DIO, are essentially variational lower bounds of the mutual information between the context matrix and the correct option.
+We propose that structuring neural network architectures around the causal chains can lead to enhanced reasoning performance.
+We also suggest that the learning objectives of end-to-end reasoning models, including DIO, are essentially variational lower bounds of the mutual information between the contexts and the correct options.
 Consequently, both performance and learning efficacy are directly impacted by the looseness of these lower bounds.

### v7

@@
 Abstract
-Despite the outstanding performance of current deep learning models across various domains, their fundamental bottleneck in abstract reasoning remains unresolved.
-To address this challenge, the academic community has introduced Raven’s Progressive Matrices (RPM) problems as an authoritative benchmark for evaluating the abstract reasoning capabilities of deep learning algorithms, with a focus on core intelligence dimensions such as abstract reasoning, pattern recognition, and complex problem-solving.
-Therefore, this paper centers on solving RPM problems, aiming to contribute to enhancing the abstract reasoning abilities of machine intelligence.
-Firstly, this paper adopts a “causal chain modeling” perspective to systematically analyze the complete causal chain in RPM tasks: image $\rightarrow$ abstract attributes $\rightarrow$ progressive attribute patterns $\rightarrow$ pattern consistency $\rightarrow$ correct answer.
-Based on this analysis, the network architecture of the baseline model DIO is designed.
-However, experiments reveal that the optimization objective formulated for DIO, namely maximizing the variational lower bound of mutual information between the context and the correct option, fails to enable the model to genuinely acquire the predefined human reasoning logic.
-This is attributed to two main reasons: the tightness of the lower bound significantly impacts the effectiveness of mutual information maximization, and mutual information, as a statistical measure, does not capture the causal relationship between subjects and objects.
-To overcome these limitations, this paper progressively proposes three improvement methods:
+Despite the outstanding performance of deep learning across various domains, their fundamental bottleneck in abstract reasoning remains unresolved.
+To tackle this challenge, the community has introduced Raven’s Progressive Matrices (RPM) as an authoritative benchmark for evaluating abstract reasoning, pattern recognition, and complex problem-solving capabilities.
+Centering on RPM, this paper aims to enhance the abstract reasoning ability of machine intelligence.
+This paper proposes a unified “causal-information” framework that systematically addresses three pivotal questions: (i) how to embed a human-like reasoning chain into network architecture; (ii) how to tighten the variational lower bound on mutual information, which limits existing models; and (iii) how to reduce semantic misalignment between learned representations and human-like causal chains.
+We deliver four incremental contributions:
   1. 1.
-    Brando: By generating trainable hypothetical incorrect options, the Brando explicitly tightens the variational lower bound of mutual information.
+    DIO – a baseline that explicitly encodes the full causal chain “image $\rightarrow$ attributes $\rightarrow$ patterns $\rightarrow$ consistency $\rightarrow$ answer” and reveals that its learning objective, like that of previous end-to-end RPM solvers, is essentially a variational lower bound on the mutual information between contexts and solutions, thereby exposing how the tightness of this bound governs the performance of such models;
   2. 2.
-    WORLD: Abandoning the generative assumption, the WORLD constructs a Gaussian mixture model for the feature distribution of RPM images, providing an infinite and informative set of incorrect options through weighted sampling to further tighten the lower bound.
+    Brando – identifies that one key opportunity for tightening this bound is to supply additional “constructive incorrect options,” and designs a learnable mapper to provide such options, offering a novel strategy for tightening the variational lower bound on mutual information in reasoning models;
   3. 3.
-    DIEGO: Introducing metadata supervision, DIEGO directly corrects semantic biases in the “abstract attributes $\rightarrow$ progressive patterns” link of the causal chain, aligning the model’s internal representations with human reasoning frameworks.
-Experiments demonstrate that these three methods not only significantly enhance the accuracy of discriminative RPM tasks but also empower DIO to solve open-ended generative RPM problems.
-The findings of this paper offer causal-driven architectural design principles, learning objective refinement strategies, and novel insights into cross-modal generation for abstract reasoning research.
-Index Terms:
-Abstract Reasoning, Raven’s Progressive Matrices, Generative RPM, Variational Mutual Information, Causal Chain Modeling, Representation Learning.
+    WORLD – models RPM image features with a Gaussian mixture, enabling unlimited, diverse, and target-sampled incorrect options that push DIO toward the theoretical bound and empower open-ended generative RPM solving;
+  4. 4.
+    DIEGO – rectifies semantic biases in the “attributes $\rightarrow$ patterns” link, aligning internal representations with human causal frameworks and boosting out-of-distribution (o.o.d.) generalization.
+Experiments show that our complete model achieves significant performance gains on RAVEN, I-RAVEN, and PGM benchmarks, exhibits stronger generalization on several o.o.d.
+sub-tasks, and is capable of generating structurally diverse RPM solutions.
+This study not only achieves remarkable performance in abstract visual reasoning but also validates “causal-chain-modeling plus mutual-information-tightening” as a core design principle for AI.
 I Introduction
-Deep learning, a cross-disciplinary technology that draws inspiration from the operational mechanisms of the human brain, has in recent years showcased formidable potential and transformative capabilities across numerous fields, with particularly outstanding accomplishments in generative tasks.
-Key technologies, which include Generative Adversarial Networks (GANs)[1], Variational Autoencoders (VAEs)[2], Diffusion Model[3], and the Transformer architecture that is based on the attention mechanism[4], empower models to generate high-quality and realistic images, texts, and videos from random noise or given conditions, thereby expanding the application scope of artificial intelligence in the realm of content creation.
-In computer vision, deep learning, which makes use of deep convolutional neural networks and their variants, can efficiently extract and accurately classify image features, outperforming traditional methods in tasks such as object detection and image segmentation [5, 6, 7].
-In the field of natural language processing, language models that are built on deep learning can automatically learn language grammar, semantics, and contextual associations, performing exceptionally well in tasks such as machine translation and text generation, and propelling the technology towards intelligence and automation [4, 8, 9].
-In the multimodal cross-field of visual question answering systems, deep learning, which integrates visual and language information, can achieve joint understanding and reasoning, accurately answering natural language questions that are related to images [11].
-Currently, research on Large Language Models (LLMs) is a focal point in the field of deep learning.
-These models, which possess a large number of parameters and are trained on massive data, demonstrate strong generalization ability in various natural language processing tasks, providing efficient and unified solutions for different tasks and facilitating the progress of artificial intelligence towards generalization and intelligence.
-However, the academic community has recently found that although LLMs have achieved remarkable results in language understanding and generation, they encounter obvious performance bottlenecks in abstract reasoning tasks[12, 13, 14, 15, 16].
-Abstract reasoning ability, which is a vital component of human advanced intelligence and involves the understanding of abstract concepts, the inference of logical relationships, and the formulation of problem-solving strategies, serves as a key indicator for measuring the intelligence level of models.
-The shortcomings of LLMs not only expose the limitations of existing models in handling complex cognitive tasks but also underscore the significance of enhancing their abstract reasoning ability for achieving more powerful and general artificial intelligence.
-Based on the above, this paper centers on the issue of improving the abstract reasoning ability of deep learning models.
-It explores effective model architecture design, training strategy optimization, and data augmentation methods, which aim to offer theoretical and practical support for breaking through the performance bottlenecks of existing models in abstract reasoning tasks and promoting the application and expansion of deep learning technology in a broader range of cognitive tasks.
+Deep learning, a cross-disciplinary technology inspired by the brain’s operational mechanisms, has demonstrated formidable potential across numerous fields and achieved outstanding accomplishments in generative tasks.
+Key technologies, including Generative Adversarial Networks (GANs)[1], Variational Autoencoders (VAEs)[2] and the attention-based Transformer architecture[3], enable models to generate high-quality images, texts and videos from random noise or given conditions, thereby expanding AI’s scope in content creation.
+In CV, deep CNNs and their variants extract and classify features efficiently, surpassing traditional methods in detection and segmentation [4, 5, 6].
+In NLP, deep language models learn grammar, semantics, and context, excelling in translation and generation [3, 7, 8].
+In multimodal VQA, fused visual-linguistic representations enable joint reasoning to answer natural-language questions about images [10].
+Large Language Models (LLMs), with massive parameters and data, now dominate deep learning by delivering unified, high-performance NLP solutions and driving AI toward generality.
+Yet recent studies [11, 12, 13] consistently show that LLMs encounter clear bottlenecks in abstract reasoning tasks that demand concept understanding, logical reasoning, and strategy formation, which are core hallmarks of human intelligence and key yardsticks for model capability.
+These shortcomings reveal current limits in tackling complex cognition and underscore the critical need to boost LLMs’ abstract reasoning capacity for stronger, more general artificial intelligence.
+This paper focuses on enhancing deep learning’s abstract reasoning ability through effective model architecture design, training strategy optimization, and data augmentation methods that aim to provide both theory and practice to break current bottlenecks and extend deep learning to a broader perspective and cognitive tasks.
 II Problem Description
-In artificial intelligence, abstract reasoning ability stands as a core hallmark of intelligence and a critical prerequisite for achieving human-level or superhuman AI.
-Yet current AI technologies are constrained by heavy reliance on large-scale annotated data, performing poorly in abstract tasks requiring causal reasoning, such as medical diagnosis and financial analysis.
-While academic efforts have produced datasets like RAVEN [18] and PGM [19] for benchmarking reasoning capabilities, designing algorithms to handle symbolic logic, causality, and uncertainty remains challenging.
-Strengthening this ability demands algorithmic innovation integrated with interdisciplinary insights from cognitive science, neuroscience, and beyond, enabling breakthroughs through cross-domain collaboration.
+In AI, abstract reasoning is regarded as a core hallmark of intelligence and a key prerequisite for achieving human-level or superhuman AI.
+The research community has proposed Raven’s Progressive Matrices (RPM)[14] as an evaluation benchmark, thereby presenting the challenge of designing algorithms that simultaneously handle symbolic logic, causal modeling, and uncertainty.
 II-A Raven’s Progressive Matrices
-Raven’s Progressive Matrices (RPM) [17] serves as an authoritative benchmark for evaluating the advanced cognitive capabilities of deep learning algorithms, focusing on core intelligence dimensions such as abstract reasoning, pattern recognition, and complex problem-solving.
-Its standardized testing framework demonstrates high correlation with human intelligence metrics, making it a reliable and stable tool for assessing the cognitive complexity of machine learning models.
-In current research, the RAVEN dataset [18] and PGM dataset [19] have emerged as the two most representative standardized test paradigms in the RPM task domain.
+Raven’s Progressive Matrices (RPM) [14] constitute an authoritative benchmark for assessing higher-order cognitive abilities in deep-learning systems, targeting core facets of intelligence such as abstract reasoning, pattern recognition and complex problem solving.
+Their standardized protocol correlates strongly with human intelligence metrics, offering a reliable gauge of model cognitive complexity.
+Today, the RAVEN [15] and PGM [16] datasets stand as the two most representative RPM testbeds.
 The RAVEN dataset comprises structured RPM instances, with each test unit containing 16 images: 8 forming a “context matrix” as the reasoning basis, and the remaining 8 constituting a candidate option pool.
@@
 Meanwhile, models like NCD[21], SAVIR-T[22], as well as neuro-symbolic systems (including PrAE, NVSA, ALANS[23, 24, 25]), have achieved promising results in both enhancing model interpretability and improving reasoning accuracy.
-RS-CNN and RS-TRAN[29] demonstrate exceptional capabilities in addressing problems related to RPM, while Triple-CFN[31] stands out among numerous models by implicitly extracting and indexing information about concepts and reasoning units, effectively enhancing reasoning accuracy.
+RS-CNN and RS-TRAN[26] demonstrate exceptional capabilities in addressing problems related to RPM.
 III Methodology
@@
 IV-D The Progressive Pattern Consistency Evaluation Module
-As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive pattern representations $P_{\alpha jkl}$ that share the same $j$-index and $l$-index.
+As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive-pattern representations $P_{\alpha jkl}$ that share the same $j$-index and $l$-index.
 Specifically, the Transformer Encoder performs a dimension-preserving mapping on $\{P_{\alpha jkl}|k\in[1,9]\}$ to obtain $\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$.
-Repeating this for all $j\in[1,N]$ and $j\in[1,4]$ yields the complete representation set $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$.
+Repeating this for all $j\in[1,N]$ and $l\in[1,4]$ yields the complete representation set $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$.
 Finally, each vector in this full set is mapped to a scalar via an MLP, producing the consistency score $s_{\alpha jkl}$.
-This process can be formally expressed as:
-$\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$ $=\text{Trans-Encoder}(\{P_{\alpha jkl}|k\in[1,9]\})$ (3)
-$s_{\alpha jkl}$ $=\text{MLP}(P^{\prime}_{\alpha jkl})$ (4)
-The entire module can be represented as:
+This entire module can be represented as:
 $\{s_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$
-$=\text{PCEM}(\{P_{\alpha jkl}|k\in[1,9]\}_{i})|_{j=1,l=1}^{N,4}$ (5)
+$=\text{PCEM}(\{P_{\alpha jkl}|k\in[1,9]\})|_{j=1,l=1}^{N,4}$ (3)
 IV-E The Option Evaluation Module.
@@
 $S_{\alpha}$ $=\frac{1}{N\cdot 9\cdot 4}\big(\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{4}s_{\alpha jkl}\big)$ (4)
-The softmax function then maps $\{S_{i}|i\in[8,16]\}$ to a probability distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ parameterized by DIO:
+Subsequently, the softmax function maps the scores $\{S_{i}\mid i\in[8,16]\}$ derived by DIO to a probability distribution $\hat{P}(x\mid\{x_{i}\}_{i=1}^{8},\theta)$ parameterized by DIO:
 $\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{S_{\alpha}}}{\sum_{{\hat{\alpha}}=9}^{16}e^{S_{{\hat{\alpha}}}}}$ (5)
-We introduce a loss function ${\ell}_{\text{DIO}}$ to refine the DIO-parameterized distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$, thereby enabling end-to-end training of all parameters across DIO’s modules:
-${\ell}_{\text{DIO}}=-\mathbb{E}\left[\log\frac{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)}{\sum_{\hat{\alpha}=9,\hat{\alpha}\neq\alpha}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)}\right]$ (8)
-where $\mathbb{E}[\cdot]$ denotes the expectation over the data distribution.
+We introduce a loss function ${\ell}_{\text{DIO}}$ to refine this parameterized distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$, thereby enabling end-to-end training of all parameters across DIO’s modules:
+${\ell}_{\text{DIO}}=\mathbb{E}\left[\log(\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)}{\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)}+\delta)\right]-\log\delta$ (6)
+where $\mathbb{E}[\cdot]$ denotes the expectation over the data distribution; $\delta$ is a small safeguard against numerical overflow, initialized at 1 and annealed to $1\times 10^{-5}$ as training progresses.
 V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM-solving Model
-This paper proposes that causal-chain-based architecture optimization can effectively enhance the performance of deep learning networks in solving RPM problems.
-However, further investigation reveals that this methodology still fails to fundamentally improve the logical reasoning capabilities of deep learning networks.
-V-A The learning objective of the DIO
-By abstracting the specific architecture of DIO (i.e., by merging equations (1)–(6)), we obtain a unified formulation that highlights its core inductive mechanism:
+In this section, we provide a theoretical analysis anchored in DIO’s optimization objective to assess the validity of the intuitively motivated methodology of causal-chain-based architecture optimization.
+V-A The Learning Objective of DIO
+By abstracting away the specific architecture of DIO (i.e., by collapsing Equations (1)–(4)), we obtain a unified formulation that highlights its core inductive mechanism:
 $S_{\alpha}$ $=\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})$ (7)
-By substituting Equation (7), DIO can instantiate the parameterized probability distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$:
+Using Equation (5) and (7), we recast the parameterized distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ into a form that explicitly embeds DIO’s inductive mechanism :
 $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (8)
-Notably, $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, which denotes the mutual information between the context matrix $\{x_{i}|i\in[1,8]\}$ and the correct option $x_{\alpha}$, admits the following variational lower bound:
-$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}P(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (11)
-where $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ represents the true probability that $x_{\alpha}$ can serve as the correct option for the given context matrix $\{x_{i}|i\in[1,8]\}$ and $K$ denotes the number of all potential incorrect options.
-If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a sufficient approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$, then substituting Equation (10) into Equation (11) yields a mutual information lower bound parameterized by DIO:
-$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right],$ (12)
-$\quad\quad\quad\quad\quad\quad\quad\text{where}\quad K=7$
-It can be observed that the parameterized lower bound within Equation (12) aligns with DIO’s optimization objective, as expressed by the following equation:
-$\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]=-\ell_{\text{DIO}}+\log K$ (13)
-Equation (13) reveals that DIO’s learning objective (i.e.
-maximizing $\ell_{\text{DIO}}$) is neither the human-like reasoning logic embedded in RPM problems nor even the suboptimal approach of directly maximizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather optimizing the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
-Furthermore, based on the properties of logarithmic functions, the following inequality holds:
-$-{\ell}_{\text{DIO}}>\mathbb{E}\left[\log\frac{\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\sum_{\hat{\alpha}=9}^{16}\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (14)
-The right-hand side of Inequality (14) is precisely the classical formulation for optimizing the probabilistic distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the Cross-Entropy (CE) loss function.
-From this, it can be inferred that for those past end-to-end RPM-solving models that parameterize $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ and employ the CE loss function, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
-Equation (13) and Inequality (14) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM-solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
-In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
-V-B Introducing Hypothetical Options to Derive a Tight Lower Bound for Mutual Information.
-This paper contends that by utilizing DIO to parameterize a tighter variational lower bound and adjusting its optimization objective to maximize this new lower bound, we can achieve a more effective maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, ultimately enhancing DIO’s performance.
-By analyzing Formula (12), we deduce that increasing the number $K$ of incorrect options encountered by DIO can guide its learning objective toward a tighter variational lower bound.
-However, RPM problem instances only offer a limited number (seven) of incorrect options.
-To this end, this paper proposes to introduce additional hypothetical incorrect options for DIO.
-Specifically, we attempt to introduce a set of hypothetical options $\{x_{i}|i\in[17,\beta]\}$ into the training pipeline of DIO.
-Consequently, the conditional distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ is reparameterized as:
+Accordingly, the loss function $\ell_{\text{DIO}}$ can be rewritten as:
+$\ell_{\text{DIO}}=\mathbb{E}\left[\log\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}\right]$ (9)
+For the sake of discussion, we have temporarily removed the safeguard $\delta$.
+Notably, $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, which denotes the mutual information between the context matrix $\{x_{i}|i\in[1,8]\}$ and the correct option $x_{\alpha}$, admits the following variational lower bound [30]:
+$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq-\mathbb{E}\left[\log\frac{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}P(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}{P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}\right]$ (10)
+Here, $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ represents the true probability that $x_{\alpha}$ is the correct option given the context matrix $\{x_{i}|i\in[1,8]\}$, i.e., that $x_{\alpha}$ can serve as the solution.
+$K$ denotes the total number of potential incorrect options.
+If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a reliable approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$, then substituting Equation (8) into Equation (10) yields a mutual information lower bound parameterized by DIO:
+$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq-\mathbb{E}\left[\log\frac{\frac{1}{7}\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}\right]$ (11)
+It can be observed that the parameterized lower bound in Equation (11) matches DIO’s optimization objective in Equation (9):
+$-\mathbb{E}\left[\log\frac{\frac{1}{7}\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}\right]=\log 7-\ell_{\text{DIO}}$ (12)
+Equation (12) reveals that DIO’s learning objective (i.e., maximizing $\ell_{\text{DIO}}$) is neither human-like reasoning logic embedded in RPM problems nor even directly optimizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather maximizing a variational lower bound on that mutual information.
+Furthermore, by the properties of logarithmic functions, the following inequality holds:
+$-\log\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)>{\ell}_{\text{DIO}}$ (13)
+The left-hand side of Inequality (13) is precisely the expression for the loss function when DIO is optimized using the cross-entropy (CE) loss.
+Notably, inequality (13) abstracts away the specific feed-forward architecture of DIO, primarily retaining its inductive mechanism.
+This implies that the same inequality holds for any model parameterizing the distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$.
+Consequently, it can be inferred that for those prior end-to-end RPM-solving models trained with the CE loss function, their learning objectives amount to optimizing a looser lower bound on $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
+Equation (12) and Inequality (13) jointly uncover a pivotal fact: whether training the DIO model using $\ell_{\text{DIO}}$ or previous end-to-end RPM solving models via the CE loss, their learning objectives are essentially maximizing lower bounds on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
+As a result, regardless of how sophisticated the model architectures are, the tightness of these bounds directly and significantly determines the performance of such models.
+This observation further implies that in RPM problems the model architecture optimization is intrinsically limited; the crux is directing the model’s objective to maximize a tighter lower bound on the mutual information.
+V-B Tightening the Lower Bound via Hypothetical Options
+This paper contends that if DIO parameterizes a tighter variational lower bound and its learning objective explicitly targets this bound, we can more effectively maximize the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, thereby further enhancing DIO’s performance.
+By analyzing Equation (10)–(12), we deduce that increasing the number $K$ of incorrect options presented to DIO can steer its learning objective toward a tighter variational lower bound.
+However, each RPM instance provides only seven such options.
+Given this, we propose supplying DIO with additional hypothetical incorrect options.
+Specifically, we introduce a set of hypothetical options, denoted as $\{x_{i}|i\in[17,\beta]\}$, into the original RPM instance $\{x_{i}|i\in[1,16]\}$, allowing DIO to recast the parameterized distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ from the form given in Equation (8) into the following:
 $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (14)
-Correspondingly, the optimization objective of DIO, the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, is updated as follows:
-$I\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{\beta-9}\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]$ (16)
-Intuitively, the looseness of the updated lower bound has been reduced.
-However, as shown in Formula (16), if these introduced hypothetical options fail to influence DIO’s decisions, this new lower bound, when used as an optimization objective, will not bring DIO any significant improvement.
-Specifically, when the probabilities that DIO assigns to the hypothetical options $\{x_{i}|i\in[17,\beta]\}$ approach zero, whereby these options cease to perturb the original distribution $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16]\}$, the variational lower bounds derived from Equation (16) and Equation (12) become equivalent when employed as loss functions.
+Correspondingly, the variational lower bound on mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ is updated as follows:
+$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq-\mathbb{E}\left[\log\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{(\beta-9)\cdot e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}\right]$ (15)
+Intuitively, this new lower bound is tighter. After incorporating the safeguard $\delta$, adopting it as the loss function for DIO should lead to improved performance.
+However, this remains merely superficial.
+As shown in Equation (15), if these introduced hypothetical options fail to influence DIO’s decisions, this new lower bound, when used as an optimization objective, will not bring DIO any significant improvement.
+Specifically, when the probabilities that DIO assigns to the hypothetical options $\{x_{i}|i\in[17,\beta]\}$ approach zero, whereby these options cease to perturb the original distribution $\hat{P}(x_{{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)$, the variational lower bounds derived from Equation (11) and (15) become equivalent when employed as loss functions.
 This demonstrates that unconstructive options, which are readily assigned low weights and low probabilities by DIO, contribute nothing to tightening the lower bound.
-Therefore, this paper proposes introducing extra constructive hypothetical options into DIO, options which can reinforce acceptance of the correct options and simultaneously strengthen repulsion of the incorrect ones, so as to steer its learning objective toward a tighter variational lower bound and, in turn, ultimately further enhance its performance.
-VI Brando: Tightening the Variational Lower Bound of Mutual Information by Constructing Hypothetical Options.
-This paper has put forward the idea that incorporating extra constructive hypothetical options into DIO’s training process offers a promising avenue for boosting its performance.
-To this end, this section introduces a training framework specifically designed to furnish DIO with such hypothetical options, thereby enhancing its overall effectiveness.
+In a nutshell, this paper proposes introducing extra constructive hypothetical options into DIO, such options that can reinforce acceptance of the correct options and strengthen repulsion of the incorrect ones, so as to steer its learning objective toward a tighter variational lower bound on mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ and, in turn, ultimately further enhance its performance.
+VI Brando: Tightening the Variational Lower Bound on Mutual Information by Constructing Hypothetical Options.
+This paper proposes incorporating additional constructive hypothetical options into DIO’s training process as a promising avenue for enhancing its performance.
+To this end, this section introduces a training framework, Brando (Bounded Random Options Mapping Method).
 VI-A Trainable Hypothetical Opions.
-The RPM problem expects the training process of the solving model to involve only a limited number of supervised signals, such as the index of correct options.
-Relying on manual design to create constructive options goes against this original intent of RPM problems.
-Therefore, this paper proposes training a set of constructive options instead.
-This methodology is referred to as Brando (Bounded Random Options Mapping Method).
-This paper establishes a set of trainable hypothetical options $\{x_{i}|i\in[17,\beta]\}$ for the context matrix $\{x_{i}|i\in[1,8]\}$.
+The RPM problem expects the solving model to be trained with only limited supervision, such as the index of the correct option.
+Relying on manual design to create constructive options contradicts this intent; we therefore propose learning a set of constructive options instead.
+We establish a set of trainable hypothetical options $\{x_{i}|i\in[17,\beta]\}$ for the context matrix $\{x_{i}|i\in[1,8]\}$.
 This paper is not concerned with the specific attributes of these options; rather, it focuses solely on how much these options can alter DIO’s evaluation of the original options $\{x_{i}|i\in[9,16]\}$.
 In other words, we are solely concerned with the constructiveness of these hypothetical options with respect to the context matrix $\{x_{i}|i\in[1,8]\}$ in DIO, rather than their specific configurations.
-To this end, this paper designs a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options for their constructiveness:
-$\ell_{\text{Brando}}=\ell_{\text{up}}+\ell_{down}$ (17)
-Where $\ell_{\text{up}}$:
-$\ell_{\text{up}}=-\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (18)
-and $\ell_{\text{down}}$:
-$\ell_{\text{down}}=-\frac{1}{\beta-17}\sum_{{\hat{\beta}}=17}^{\beta}\log\frac{e^{\text{DIO}(\{x_{i},x_{\hat{\beta}}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (19)
+To this end, we design a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options for constructiveness:
+$\ell_{\text{Brando}}=\ell_{\text{up}}+\frac{\ell_{down}}{\beta-17}$ (16)
+where the loss term $\ell_{\text{up}}$:
+$\ell_{\text{up}}=\log(\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}+\delta)-\log\delta$ (17)
+and the loss term $\ell_{\text{down}}$:
+$\ell_{\text{down}}=\sum_{{\hat{\beta}}=17}^{\beta}\log(\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\hat{\beta}}|i\in[1,8]\})}}+\delta)-\log\delta$ (18)
 In this novel loss function:
   1. 1.
-    It can be observed that $\ell_{\text{up}}$ requires the probability $\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ derived by DIO to be greater than all elements in the set $\{\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|{\hat{\alpha}}\in[9,\beta],{\hat{\alpha}}\neq\alpha\}$.
-In other words, $\ell_{\text{up}}$ demands that DIO be able to identify the correct option amidst the interference from the hypothetical options $\{x_{i}|i\in[17,\beta]\}$.
+    The loss term $\ell_{\text{up}}$ requires DIO to be able to identify the correct option from a pool that combines the original candidate options $\{x_{i}|i\in[9,16]\}$ with the hypothetical options $\{x_{i}|i\in[17,\beta]\}$.
+It serves the purpose of tightening the variational lower bound on mutual information.
   2. 2.
-    On the other hand, $\ell_{\text{down}}$ constrains that any element in the set $\{\hat{P}(x_{\hat{\beta}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\beta}\in[17,\beta]\}$ must be higher than all elements in the set $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16],\hat{\alpha}\neq\alpha\}$.
-In other words, $\ell_{\text{down}}$ ensures hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ outscore the original incorrect options $\{x_{i}\mid i\in[9,16],i\neq\alpha\}$ under DIO.
-In the loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive ones.
-By adopting $\ell_{\text{Brando}}$ as the training loss function, DIO not only obtains constructive hypothetical options but also parameterizes a tighter variational lower bound, thereby maximizing the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.
+    The loss term $\ell_{\text{down}}$ encourages hypothetical options $\{x_{i}|i\in[17,\beta]\}$ to attain higher probabilities than the original incorrect options $\{x_{i}|i\in[9,16],i\neq\alpha\}$ under DIO.
+It acts as a safeguard to prevent hypothetical options from degenerating into unconstructive ones.
+By adopting $\ell_{\text{Brando}}$ as the training loss function, DIO not only obtains constructive hypothetical options but also parameterizes a tighter variational lower bound on $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
 VI-B Map Out Hypothetical Options
@@
 $\{x_{i}|i\in[17,\beta]\}=\mathcal{F}(\{x_{i}|i\in[1,8]\})$ (19)
-Evidently, the mapping $\mathcal{F}(\cdot)$ is required to derive a variable number ($\beta-17$) and diverse hypothetical options for the context matrix $\{x_{i}|i\in[1,8]\}$.
-This indicates that conventional deep neural networks struggle to undertake the task of the mapping $\mathcal{F}(\cdot)$.
+Clearly, the mapping $\mathcal{F}(\cdot)$ needs to yield structurally diverse and numerically flexible ($\beta-17$) hypothetical options conditioned on the context matrix $\{x_{i}|i\in[1,8]\}$, implying an inherent many-to-many property.
+This indicates that conventional deep neural networks struggle to fulfill the role of mapping $\mathcal{F}(\cdot)$.
 Firstly, constructing a network with variable and dynamic output specifications is highly challenging.
-Additionally, from the perspective of neural network optimization, achieving a many-to-many mapping is infeasible.
-Given these challenges, we need to explore an alternative approach.
-Therefore, this paper proposes a many-to-one mapping approach:
-$x_{\hat{i}}=\hat{\mathcal{F}}(\{x_{i},o_{\hat{i}}|i\in[1,8]\})$ (21)
-$\text{Where}\quad o_{\hat{i}}\sim N(\mu,\sigma^{2})$
+Moreover, from the perspective of neural network optimization, achieving many-to-many mappings in a naive end-to-end fashion is infeasible.
+Given these challenges, we propose a many-to-one mapping as a substitute:
+$x_{\hat{i}}=\hat{\mathcal{F}}(\{x_{i},o_{\hat{i}}|i\in[1,8]\}),\,o_{\hat{i}}\sim\mathcal{N}(\mu,\sigma^{2})$ (20)
 […]
@@
 […]
+VI-C Instantiate Mapping $\hat{\mathcal{F}}$ As a Network.
 […]
-$z_{\hat{i}j}=\text{Brando}(\{z_{ij},o_{\hat{i}}|i\in[1,8]\})$ (23)
-$\text{Where}\quad o_{\hat{i}}\sim N(\mu,\sigma^{2})$
-Formula (23) shows that we move away from the option-mapping paradigm and instead adopt a representation-mapping approach, easing the learning pressure on Brando.
-According to Formula (23), we have designed the structure of the Brando network as shown in the Figure 5.
+$z_{\hat{i}j}=\text{Brando}(\{z_{ij},o_{\hat{i}}|i\in[1,8]\}),\,o_{\hat{i}}\sim\mathcal{N}(\mu,\sigma^{2})$ (22)
+Formula (22) shows that we abandon the option-mapping paradigm in favor of a representation-mapping approach, in order to reduce the computational load on Brando.
+The structure of the Brando network is shown in the Figure 5.
 Figure 5: The Brando network.
@@
 Except for the first layer, the input to all subsequent layers comes from the output of the preceding layer.
-For the first layer, $\{z_{ij}|i\in[1,8]\}$ are the features extracted by DIO from the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution $N(\mu,\sigma^{2})$ parameterized by mean $\mu$ and variance $\sigma^{2}$.
+For the first layer, $\{z_{ij}|i\in[1,8]\}$ are the features extracted by DIO from the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ parameterized by mean $\mu$ and variance $\sigma^{2}$.
 For the final layer, this paper treats its output $o_{\hat{i}}$ as the representation $z_{\hat{i}j}$ of a hypothetical option $x_{\hat{i}}$ provided by Brando with respect to the context matrix $\{x_{i}|i\in[1,8]\}$.
@@
   2. 2.
-    After these $Q$, $K$, and $V$ representations are appended positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$:
+    After these $Q$, $K$, and $V$ representations are appended with positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$:
   3. 3.
-    As indicated by the computation step 1 in the Figure 5, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input feature information stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the memory outputs $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$.
+    As indicated by computation step 1 in Figure 5, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input feature information stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the memory outputs $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$.
 The learning process can be interpreted as:
     $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}=\frac{\phi(Q)M_{n-1}}{\phi(Q)B_{n-1}}$ (23)
-    Where basis function $\phi(\cdot)$ is an element-wise ELU+1 operation.
+    where the basis function $\phi(\cdot)$ is an element-wise ELU+1 operation.
   4. 4.
-    As indicated by the computation steps 2 and 3 in the Figure 5, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the $K$ and $V$ representations derived from the current input features.
-The update process can be expressed as:
+    As indicated by computation steps 2 and 3 in Figure 5, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the $K$ and $V$ representations derived from the current input features, an update process inspired by infinite attention [31] and expressed as:
     $M_{n}$ $=M_{n-1}+\phi(K)^{T}\left(V-\frac{\phi(K)M_{n-1}}{\phi(K)B_{n-1}}\right)$ (24)
@@
     ${o}_{\hat{i}}\leftarrow{o}_{\mu}+e^{\frac{o_{\ln\sigma}}{2}}\cdot\epsilon$ (30)
-    The $\epsilon$ represents a random sample drawn from $N(0,1)$ and MLP()($\cdot$) refers to a Multi-Layer Perceptron capable of doubling the dimension of $^{o}_{\hat{i}}$.
+    The $\epsilon$ represents a random sample drawn from $\mathcal{N}(0,1)$ and MLP()($\cdot$) refers to a Multi-Layer Perceptron capable of doubling the dimension of $^{o}_{\hat{i}}$.
 As indicated in Formulas (24)-(25), in each logical layer, the memory bank $M_{n}$ and the buffer $B_{n}$ are dynamically updated as each new batch of context matrices is input.
@@
 This satisfies the requirement to map diverse hypothetical options for the context matrix.
-VI-C The Combination of Brando and DIO.
+VI-D The Combination of Brando and DIO.
 Prior to integrating the Brando network with DIO, DIO must be pre-trained on the original option pool $\{x_{i}|i\in[9,16]\}$ using the loss function $\ell_{\text{DIO}}$.
@@
 Another key point is that during the joint training process, DIO’s parameters excluding those in the image feature extraction module will be frozen every other iteration.
-VII WORLD: Tightening the Variational Lower Bound of Mutual Information Through Deconstructing the Problem Space
+VII WORLD: Tightening the Variational Lower Bound on Mutual Information Through Deconstructing the Problem Space
 It is readily apparent that the architecture of the Brando network leaves substantial room for theoretical refinement.
@@
 Moreover, this paper also points out that the introduced additional incorrect options must avoid being unconstructive.
-The Brando network essentially functions as a distribution transformation mechanism, whose task is to map samples $o_{\hat{i}}$ drawn from a Gaussian distribution $N(\mu,\sigma^{2})$ into features $z_{\hat{i}j}$ of incorrect options customized for a given context matrix.
+The Brando network essentially functions as a distribution transformation mechanism, whose task is to map samples $o_{\hat{i}}$ drawn from a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ into features $z_{\hat{i}j}$ of incorrect options customized for a given context matrix.
 This paper argues that instead of implicitly estimating the distribution of incorrect option features for each given context matrix, it is more effective to explicitly estimate the distribution of all RPM image features $z_{ij}$.
@@
 Given these advantages, we shift our focus to transforming the underlying distribution of RPM image features into a GMM.
-VII-B Model the Distribution of RPM Image Features as a Gaussian Mixture Model.
+VII-B Model the Distribution of RPM Image Features as a GMM.
 To model the distribution of features encoded by DIO for RPM images as a GMM, this paper proposes the WORLD method.
@@
 Subsequently, by projecting the image features $z_{ij}$ encoded by DIO onto this constructed GMM framework, we can model the distributional characteristics of $z_{ij}$.
-In order to enable the features $z_{ij}$ to be more easily projected onto the constructed GMM $\{U_{m}|m\in[1,M]\}$, the way DIO encodes features for RPM images requires minor modifications.
+In order to enable the features $z_{ij}$ to be projected more easily onto the constructed GMM $\{U_{m}|m\in[1,M]\}$, the way DIO encodes features for RPM images requires minor modifications.
 The modifications are shown in Figure 7.
 Figure 7: The modified Image Feature Extraction Module.
-Specifically, doubling the overall dimension of the ViT within DIO’s image feature extraction module can achieve the desired effect, causing the dimension of the latent features encoded for RPM images to double accordingly.
-After doubling the dimension, the resulting latent features are evenly split into two parts, denoted as $u_{ij}$ and $b_{ij}$ respectively.
-Then, the sum of $u_{ij}$ and $b_{ij}$ is taken as the new feature representation $z_{ij}$ for the image $x_{i}$.
-This process can be expressed by the following formula:
-$\{u_{ij},b_{ij}\}=\text{ViT}(x_{i}),\,z_{ij}=u_{ij}+b_{ij}$ (34)
+Specifically, doubling the overall width of the ViT within DIO’s image feature extraction module can achieve the desired effect, causing the dimensions of the latent features encoded for RPM images to double accordingly.
+These doubled features are evenly split into two parts, denoted as $u_{ij}$ and $b_{ij}$.
+Then, their sum is taken as the new feature representation $z_{ij}$ for the image $x_{i}$.
+This updated module can be expressed by the following formula:
+$\{u_{ij},b_{ij}\}=\text{WORLD-FEM}(x_{i}),\,z_{ij}=u_{ij}+b_{ij}$ (33)
 The new derivation method for $z_{ij}$ ensures that, when $b_{ij}$ follows a standard Gaussian distribution $\mathcal{N}(0,1)$, the distribution of $z_{ij}$ explicitly adheres to a Gaussian distribution $\mathcal{N}(u_{ij},1)$.
-Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the constructed GMM $\{U_{m}|m\in[1,M]\}$ and make the latent features $b_{ij}$ follow the distribution $N(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM with trainable means and known variances.
+Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the constructed GMM $\{U_{m}|m\in[1,M]\}$ and make the latent features $b_{ij}$ follow the distribution $\mathcal{N}(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM with trainable means and known variances.
 The aforementioned method for modeling the distribution of $z_{ij}$ can be implemented by introducing five additional loss function terms:
@@
     A loss term $\ell_{1}$ aims to encourage the mean of the distribution of $z_{ij}$ to approximate the constructed GMM $\{U_{m}|m\in[1,M]\}$, and its mathematical expression can be formulated as:
-    $\ell_{1}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|u_{ij}-\text{gs}[{U}_{\hat{m}}]\|_{2}^{2}$ (35)
+    $\ell_{1}=\frac{1}{16\cdot N}\sum_{i=1}^{16}\sum_{j=1}^{N}\|u_{ij}-\text{gs}[{U}_{\hat{m}}]\|_{2}^{2}$ (34)
     $\text{where}\quad\hat{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{ij}-{U}_{m}\|_{2}^{2}$
@@
     A loss term $\ell_{2}$ is introduced to enforce $b_{ij}$ to follow a standard normal distribution, expressed as:
-    $\ell_{2}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|b_{ij}-\epsilon_{ij}]\|_{2}^{2}$ (36)
-    $\text{where}\quad\epsilon_{ij}\sim N(0,1)$
+    $\ell_{2}=\frac{1}{16\cdot N}\sum_{i=1}^{16}\sum_{j=1}^{N}\|b_{ij}-\epsilon_{ij}]\|_{2}^{2}$ (35)
+    $\text{where}\quad\epsilon_{ij}\sim\mathcal{N}(0,1)$
   3. 3.
     A loss term $\ell_{3}$, which is utilized to train the components within the constructed GMM $\{U_{m}|m\in[1,M]\}$, can be expressed as:
-    $\ell_{2}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|{U}_{\hat{m}}-\text{gs}[u_{ij}]\|_{2}^{2}$ (37)
+    $\ell_{3}=\frac{1}{16\cdot N}\sum_{i=1}^{16}\sum_{j=1}^{N}\|{U}_{\hat{m}}-\text{gs}[u_{ij}]\|_{2}^{2}$ (36)
     $\text{where}\quad\hat{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{ij}-{U}_{m}\|_{2}^{2}$
@@
     A mutual information loss term $\ell_{4}$, which facilitates the training of components within the constructed GMM $\{U_{m}|m\in[1,M]\}$, is expressed as:
-    $\ell_{4}$ $=\sum_{i=1}^{16}\|x_{i}-\hat{x}_{i}\|_{2}^{2}$ (38)
+    $\ell_{4}$ $=\frac{1}{16}\sum_{i=1}^{16}\|x_{i}-\hat{x}_{i}\|_{2}^{2}$ (37)
     where $\hat{x}_{i}=\mathcal{D}\left(\left\{\hat{u}_{ij}+b_{ij}\mid j\in[1,N]\right\}\right),$
@@
     A mode collapse prevention loss term $\ell_{5}$ for the latent features $u_{ij}$, defined as:
-    $\ell_{5}$ $=\sum_{i=1}^{16}\|x_{i}-\tilde{x}_{i}\|_{2}^{2}$ (39)
+    $\ell_{5}$ $=\frac{1}{16}\sum_{i=1}^{16}\|x_{i}-\tilde{x}_{i}\|_{2}^{2}$ (38)
     where $\tilde{x}_{i}=\mathcal{D}\left(\left\{\hat{u}_{ij}+\epsilon_{ij}\mid j\in[1,N]\right\}\right),$
-    $\epsilon_{ij}\sim N(0,1)$
+    $\epsilon_{ij}\sim\mathcal{N}(0,1)$
     $\hat{u}_{ij}=U_{\hat{m}}-\text{sg}[u_{ij}]+u_{ij},$
@@
 After performing a weighted summation of these five loss terms, we obtain the complete loss function $\ell_{\text{WORLD}}$:
-$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+\lambda\cdot\ell_{3}+\ell_{4}+\ell_{5}$ (40)
-The coefficient $\lambda$ for $\ell_{3}$ is set to 0.25 in order to balance its relative magnitude compared to $\ell_{1}$.
-By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss functions, we can train a DIO model where the distribution of image features $z_{ij}$ is modeled as a GMM.
+$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+\ell_{3}+\lambda_{0}\cdot\ell_{4}+\lambda_{1}\cdot\ell_{5}$ (39)
+Here, $\lambda_{0}$ and $\lambda_{1}$ are set to $5$ to balance the inductive difficulties between these terms.
+By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss function, we can train a DIO model in which the distribution of image features $z_{ij}$ is modeled as a GMM.
 Notably, when training the DIO using $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, the feedforward mechanism undergoes the transformation illustrated in Figure 8.
 Figure 8: Feedforward Process of DIO Model Incorporating $\ell_{\text{WORLD}}$ as an Additional Loss Term.
-In Figure 8, the latent variable fed into the progressive pattern induction module within DIO is no longer $z_{ij}$, but rather $u_{ij}+\epsilon$.
+As shown in Figure 8, the latent variable fed into the progressive pattern induction module within DIO is no longer $z_{ij}$, but rather $u_{ij}+\epsilon_{ij}$.
 In other words, the feedforward process indicated by Equation (IV-C) is updated as follows:
-$\{P_{jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}_{i=\alpha}$
+$\{P_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$
 $=\text{PPIM}(\{u_{ij}+\epsilon_{ij},u_{\alpha j}+\epsilon_{\alpha j}|i\in[1,8]\})|_{j=1}^{N}$ (40)
@@
 This design enables DIO to interact with a fully modeled GMM rather than deterministic image features, thereby training the DIO to make decisions based on the complete probabilistic structure of the GMM instead of specific feature instantiations.
-VII-C Leveraging the Modeled GMM to Provide Incorrect Options for Context Matrices
+VII-C Targeted Sampling From the Learned GMM.
 To steer DIO’s learning objective towards a tighter variational lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, we also need to sample sufficient and diverse incorrect option features from the modeled GMM $\{U_{m}|m\in[1,M]\}$ for each given context matrix.
 Thanks to the inherent properties of GMM, targeted sampling within the set $\{U_{m}|m\in[1,M]\}$ is relatively straightforward.
-For a given context matrix $\{x_{i}|i\in[1,8]\}$, after identifying the components in the set $\{U_{m}|m\in[1,M]\}$ that correspond to the provided correct option $x_{\alpha}$, samples randomly drawn from the remaining components can be considered, to a certain extent, as features of incorrect options.
-Identifying the components corresponding to the correct option, targeted sampling of incorrect option features, and steering the learning objective of DIO can all be achieved by introducing an additional loss function term:
-${\ell_{\text{Weighted}}}=$ $-\sum_{j=1}^{N}\log\frac{{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{\tilde{m}}+\epsilon_{\tilde{m}}|i\in[1,8]\})}{|\tau|}}}}}{\sum_{m=1,m\neq\tilde{m}}^{M}{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{m}+\epsilon_{m}|i\in[1,8]\})}{|\tau|}}}}}$ (42)
-where $\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2},\,\epsilon_{m}\sim N(0,1)$
+For a given context matrix $\{x_{i}|i\in[1,8]\}$, after identifying the components in the set $\{U_{m}|m\in[1,M]\}$ that correspond to the provided correct option $x_{\alpha}$, samples randomly drawn from the remaining components can be considered, to a certain extent, as features of the incorrect options.
+Identifying the components corresponding to the correct option, targeted sampling of the incorrect option features, and steering the learning objective of DIO can all be achieved by introducing an additional loss function term:
+${\ell_{\text{Weighted}}}$
+$=\sum_{j=1}^{N}\log(\frac{\sum_{\begin{subarray}{c}m=1\\ m\neq\tilde{m}\end{subarray}}^{M}{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{m}+\epsilon_{m}|i\in[1,8]\})}{|\tau|}}}}}{{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{\tilde{m}}+\epsilon_{\tilde{m}}|i\in[1,8]\})}{|\tau|}}}}}+\delta)+\log\delta$ (41)
+$\text{where}\quad\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2},\,\,\epsilon_{m}\sim\mathcal{N}(0,1)$
 $\tilde{\text{DIO}}$ refers to the DIO without the image feature extraction module, which is capable of processing image features into scores.
@@
 The role of $\ell_{\text{Weighted}}$ is to fine-tune the DIO model.
-Specifically, when DIO approaches optimal performance under the constraint of the loss function $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, incorporating the additional loss term $\ell_{\text{Weighted}}$ to update the loss function to $\ell_{\text{WORLD}}+\ell_{\text{DIO}}+\ell_{\text{Weighted}}$ creates additional learning capacity and performance improvement potential.
+Specifically, when DIO approaches optimal performance under the constraint of the loss function $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, incorporating the additional loss term $\ell_{\text{Weighted}}$ to update the loss function to $\ell_{\text{WORLD}}+\ell_{\text{DIO}}+\ell_{\text{Weighted}}$ creates additional learning capacity and potential of performance improvement.
 This adjustment enables the introduction of supplementary incorrect options during DIO’s late-stage training, thereby steering its learning objective toward a tighter variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, ultimately leading to further performance enhancement.
@@
 VIII-A Target Reasoning Rather Than Solving Problems
-This paper realizes that setting maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal.
-Although this statistical measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic [35], yielding a model optimized for high accuracy rather than for faithful causal reasoning.
+This paper realizes that setting the maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal.
+Although this statistical measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic [33], yielding a model optimized for high accuracy rather than faithful causal reasoning.
 Thus, the performance gains brought by mutual information may be superficial.
@@
 Given that mutual information fails to account for causality [33], the semantic alignment awaits empirical verification.
-In order to verify semantic alignment, this paper takes the “Distribute 9” dataset from the RPM-style RAVEN benchmark [18] as the research object.
+In order to verify semantic alignment, this paper takes the $\times3$ Grid” sub-task from the RPM-style RAVEN benchmark [15] as the research object.
 We conduct training on this dataset using DIO, as well as DIO supplemented with Brando and WORLD.
-Subsequently, we extract 30 instances that share strictly identical progressive patterns and compute the pairwise similarities among the progressive pattern representations that DIO encodes for these 30 instances.
+Subsequently, we extract 30 instances that share strictly identical progressive patterns and compute the pairwise similarities among the progressive-pattern representations that DIO encodes for these 30 instances.
 The calculated $30\times 30$ similarity matrices are visualized as heatmaps in Figure 9.
-Figure 9: Heatmap of Progressive Pattern Representation Extracted from the Distribute 9 under RAVEN Problems.
+Figure 9: Heatmap of Progressive-Pattern Representation Extracted from the $\times3$ Grid under RAVEN Problems.
 If DIO accurately models the causal chain embedded in RAVEN problems, then once DIO encodes 30 instances sharing the same progressive patterns into pattern representations, those representations should exhibit high cosine similarity among themselves.
@@
 This corresponds to a misalignment between the outputs $P_{\alpha{jkl}}$ produced by Formula (IV-C) and human semantics.
-Therefore, this paper proposes rectifying the progressive pattern representations $P_{\alpha{jkl}}$ encoded by DIO as an effective remedy.
+Therefore, this paper proposes rectifying the progressive-pattern representations $P_{\alpha{jkl}}$ encoded by DIO as an effective remedy.
 This paper notes that each RPM instance is accompanied by metadata describing its progressive patterns that precisely match the semantics predefined for $P_{\alpha{jkl}}$ in the DIO architecture.
@@
 DIEGO constructs a set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$.
-By assigning each vector a distinct metadata modality, this entire set can be considered as a comprehensive metadata dictionary.
-In addition, DIEGO introduces an auxiliary loss term $\ell_{\text{DIEGO}}$ that serves to align each progressive pattern representation $P_{\alpha{jkl}}$ extracted by the DIO with its corresponding reference vector in $\{V_{f}|f\in[1,F]\}$ based on the accompanying metadata.
+By assigning each vector a distinct metadata modality, this entire set can be considered a comprehensive metadata dictionary.
+An auxiliary loss term $\ell_{\text{DIEGO}}$ is introduced to align every progressive pattern representation $P_{\alpha jkl}$ extracted by DIO with its corresponding reference vector in $\{V_{f}\mid f\in[1,F]\}$ according to the accompanying metadata.
 This term is formulated as follows:
@@
 Here, $\text{Meta}_{\hat{f}l}$ is a binary indicator that flags whether the $l$-th metadata entry matches the modality associated with the $\hat{f}-$th vector within the reference dictionary $\{V_{f}|f\in[1,F]\}$.
-Meanwhile $\tau$ is a trainable scaling coefficient.
+Meanwhile, $\tau$ is a trainable scaling coefficient.
 The detailed formulation of $\ell_{\text{DIEGO}}$ elucidates our alignment mechanism.
 Guided by the metadata-derived indicators, this loss term employs logarithmic smoothing and exponential scaling to regulate the cosine similarities between each $P_{\alpha_{jkl}}$ and all vectors in the reference dictionary.
-Evidently, our objective has shifted from statistical quantity control to vector similarity control, distinguishing DIEGO from Brando and WORLD: it is no longer mutual-information-driven but progressive-pattern-driven.
+Evidently, our objective has shifted from statistical quantity control to vector similarity control, distinguishing DIEGO from Brando and WORLD in that it is no longer mutual-information-driven but progressive-pattern-driven.
 When the DIO model is trained with the composite loss $\ell_{\text{DIO}}+\ell_{\text{DIEGO}}$ and the trainable reference dictionary $\{V_{f}|f\in[1,F]\}$, the forward pass of the relevant modules proceeds as illustrated in Figure 10.
@@
 After training, a solving model that captures causal chains in RPM problems is obtained.
-It is widely accepted that introducing more supervisory signals should improve model performance.
-Indeed, previous studies have attempted to leverage metadata as additional supervision, aiming to boost both interpretability and reasoning accuracy simultaneously.
-Recent work [23, 26, 25, 29] has collectively shown, however, that this practice can backfire: directly embedding metadata into proposed models actually degrades reasoning performance.
-This paper argues that the DIO model, constructed upon causal chains, together with the DIEGO method that strategically rectifies key segments of those chains, offers a viable way to overcome this limitation.
 IX Solving Generative RPM Problems
 Generative RPM problems represent an advanced variant of RPM tasks.
-In these problems, the solving models are required to generate at least one solution that conforms to the progressive patterns within the given context matrix, rather than selecting from a predefined bounded option pool, thereby more accurately reflecting the model’s comprehension of the reasoning logic.
-As solution identification remains progressive-pattern-driven, generative RPM problems inherit the multi-solution characteristic from conventional RPM problems.
-Consequently, models are expected to produce multiple solutions that are structurally diverse, which poses significant challenges for most traditional neural networks.
+In these problems, the solving models are required to generate at least one solution that conforms to the progressive patterns within the given context matrix, rather than selecting from a predefined and bounded option pool, thereby more accurately reflecting the model’s comprehension of the reasoning logic.
+As solution identification remains progressive-pattern-driven, generative RPM problems inherit the multi-solution characteristic from conventional RPM tasks.
+Consequently, solving models are expected to produce multiple solutions that are structurally diverse, which poses significant challenges for most traditional neural networks.
+However, WORLD endows DIO with the capability to effectively tackle generative RPM problems.
 IX-A WORLD Guides DIO on Generative RPM problems
 Beyond enabling DIO to solve discriminative RPM tasks, the WORLD method also equips it with the capability to tackle generative RPM problems.
-This method constructs a complete generation pipeline for DIO: “estimating solution distribution $\rightarrow$ sampling solution distribution $\rightarrow$ decoding samples”.
+This method constructs a complete generation pipeline for DIO: “estimating the solution distribution $\rightarrow$ sampling the solution distribution $\rightarrow$ decoding the sampling results”.
 The WORLD method extends DIO by incorporating an external GMM $\{U_{m}|m\in[1,M]\}$, an additional loss term $\ell_{\text{Weighted}}$, and a decoder $\mathcal{D}(\cdot)$.
-Once trained in this framework, DIO can not only discriminate among the options for each context matrix but also directly generate the corresponding solutions.
-Specifically, as shown in Formula (42), the loss term $\ell_{\text{Weighted}}$ prompts DIO to assign a context-dependent weight to each GMM component $U_{m}$, where the weight reflects the likelihood that $U_{m}$ contains the correct options for the given context matrix.
+Once trained in this framework, DIO can not only discriminate among the options for each context matrix but also generate the corresponding solutions.
+Specifically, as shown in Formula (VII-C), the loss term $\ell_{\text{Weighted}}$ prompts DIO to assign a context-dependent weight to each GMM component $U_{m}$, where the weight reflects the likelihood that $U_{m}$ contains the correct options for the given context matrix.
 Within this loss term, the weight for each $U_{m}$ given the context matrix features $\{z_{ij}|i\in[1,8]\}$ is calculated as:
 $S_{mj}=\tilde{\text{DIO}}(\{z_{ij},U_{m}|i\in[1,8]\}).$ (43)
-Since these weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$ indicate the likelihood that the components cover the correct options, we not only gain the chance to sample diverse incorrect options from low-weight components to tighten the lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, but also have the opportunity to combine high-weight components into a solution distribution.
+Since these weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$ indicate the likelihood that the components cover the correct options, we not only gain the chance to sample diverse incorrect options from low-weight components to tighten the lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, but also have the opportunity to combine high-weight components into a solution distribution to provide promising options.
 To this end, we adopt the Gumbel-Max trick [34] to sample $N$ high-weight components as the solution distribution that DIO estimates for the context matrix $\{x_{i}|i\in[1,8]\}$.
@@
 $\hat{x}=\mathcal{D}(\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\})$ (47)
-$\text{where}\quad\epsilon_{ij}\sim N(0,1)$
-In a nutshell, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve the generative RPM problems.
-Notably, WORLD is well-aligned with the multi-solution characteristic of generative RPM problems.
-Formulas (45)–(47) indicate that the Gumbel-Max trick ensures the estimated solution distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$ exhibits diversity.
-Formula (48) further reveals that the solutions decoded from samples $\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\}$ of the estimated distribution have the potential to incorporate the sampling diversity $\epsilon_{ij}$.
-By iterating the “estimating $\rightarrow$ sampling $\rightarrow$ decoding” pipeline, which can generate diverse solutions, WORLD systematically explores multiple potential solutions and fully embraces the multi-solution logic of generative RPM tasks.
-IX-B Evaluating the Generated Solutions
-Verifying the solution accuracy of generative RPM problems has long been challenging.
-Generative RPM problems possess the multi-solution characteristic, which means that the generative solving models cannot be evaluated simply by comparing each generated solution with the provided correct option.
-This also makes it challenging for RPM experts to devise a perfect verification algorithm.
-Most prior work on generative RPM problems has used self-verification, i.e., using their own discriminative models to evaluate their own generative models.
-However, this paper argues that this self-verification paradigm is inadequate.
-In generative RPM problems, the multi-solution characteristic stems from the progressive-pattern-driven solution evaluation approach.
-This implies that the algorithms or models used to validate the accuracy of solutions for such problems should also be progressive-pattern-driven rather than mutual-information-driven.
-The DIEGO approach meets this expectation, as the DIO model trained solely with the loss term $\ell_{\text{DIEGO}}$ will be a progressive-pattern-driven solving model.
-More importantly, WORLD is a strong mutual-information-driven inductive model, yet it cannot escape the inherent drawback of neglecting reasoning logic.
-Using the progressive-pattern-driven model DIEGO to verify WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also examines the extent to which WORLD’s performance gains are merely superficial.
+$\text{where}\quad\epsilon_{ij}\sim\mathcal{N}(0,1)$
+In a nutshell, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve generative RPM problems.
+Commendably, WORLD matches the multi-solution characteristic of generative RPM problems.
+Equations (44)–(46) show that the Gumbel-Max trick transfers the diversity inherent in the Gumbel noise $G_{mj}$ to the solution index $m_{j}$, so that the estimated distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$ ultimately inherits this diversity.
+Equation (47) then demonstrates that decoding the samples $\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\}$ has the potential to embed sampling diversity $\epsilon_{ij}$ directly into the resulting solutions $\hat{x}$.
+Iterating the estimate $\rightarrow$ sample $\rightarrow$ decode pipeline, WORLD systematically explores multiple valid solutions and fully embraces the multi-solution logic of generative RPM tasks.
+IX-B Hierarchical Interaction and EMA-based Optimization.
+WORLD minimizes the loss function $\ell_{\text{WORLD}}$ to embed the RPM image features $u_{ij}+b_{ij}$ encoded by DIO into its trainable GMM $\{U_{m}|m\in[1,M]\}$.
+Evidently, the component count $M$ governs WORLD’s effectiveness: an insufficient $M$ forces DIO to compress intricate image features into a low-capacity GMM, diluting the expressivity of these features, degrading DIO’s performance, and obscuring WORLD’s true potential.
+Although increasing $M$ alleviates these issues, it also incurs computational overhead that scales linearly with $M$ and quickly drives up the cost of optimizing $\ell_{\text{WORLD}}$.
+This places the setting of $M$ in a quandary when WORLD confronts demanding generative RPM problems.
+To this end, we propose two dedicated improvements for the two loss terms $\ell_{1}$ (i.e., Formula (34)) and $\ell_{3}$ (i.e., Formula (36)) within $\ell_{\text{WORLD}}$: hierarchical interaction and exponential-moving-average (EMA)-based component optimization.
+Specifically, we adopt a G-head hierarchical interaction mechanism that splits the image feature $u_{ij}$ into $G$ independent heads $\{u^{(g)}_{ij}|g\in[1,G]\}$, partitions the $M$ trainable vectors $\{U_{m}|m\in[1,M]\}$ into $G$ groups $\{U^{(g)}_{m}|m\in[1,M],g\in[1,G]\}$, and then performs head-wise hierarchical optimization with the rewritten term $\ell_{1}$ given by:
+$\ell_{1}=\frac{1}{16N}\sum_{i=1}^{16}\sum_{j=1}^{N}\Bigl\|u_{ij}-\operatorname{concat}\bigl[U^{(g)}_{\hat{m}_{g}}\bigr]_{g=1}^{G}\Bigr\|_{2}^{2}$ (48)
+$\quad\quad\text{where}\quad\hat{m}_{g}=\underset{m\in[1,M]}{\text{argmin}}\bigl\|u^{(g)}_{ij}-U^{(g)}_{m}\bigr\|_{2}^{2}.$
+This enables the same $M$ trainable vectors to represent a GMM with $M^{G}$ components, whereby the capacity grows exponentially in $G$ instead of linearly in $M$, eliminating the need to drastically increase $M$ for greater expressivity.
+However, symmetrically extending the loss term $\ell_{3}$ in the same fashion would significantly increase optimization overhead.
+Therefore, rather than supervising the hierarchical training of components $\{U^{(g)}_{m}|m\in[1,M],g\in[1,G]\}$ with $\ell_{3}$, we switch to an EMA-based update strategy that enables these GMM components to evolve without the gradient back-propagation overhead.
+The EMA-based update for $U_{m}^{(g)}$ is given by:
+$\hat{m}_{ij}^{(g)}$ $=\underset{m\in[1,M]}{\mathrm{argmin}}\|u_{ij}^{(g)}-U_{m}^{(g)}\|_{2}^{2}$ (49)
+$n_{m}^{(g)}$ $=\sum_{i,j}\mathcal{I}[m=\hat{m}_{ij}^{(g)}],s_{m}^{(g)}=\sum_{i,j}\mathcal{I}[m=\hat{m}_{ij}^{(g)}]\frac{u_{ij}^{(g)}}{n_{m}^{(g)}}$ (50)
+$N_{m}^{(g)}$ $\leftarrow\eta N_{m}^{(g)}+\gamma\,n_{m}^{(g)},\,S_{m}^{(g)}\leftarrow\eta S_{m}^{(g)}+\gamma\,s_{m}^{(g)}$ (51)
+$U_{m}^{(g)}$ $\leftarrow\dfrac{S_{m}^{(g)}}{N_{m}^{(g)}+\varepsilon}\Bigg/\sum_{k=1}^{M}\dfrac{N_{k}^{(g)}}{N_{k}^{(g)}+\varepsilon}$ (52)
+Here, $\eta$ is 0.99; $\gamma$ is 0.02; $\varepsilon$ is $10^{-5}$; $N_{m}^{(g)}$ and $S_{m}^{(g)}$ are initialized to 0; $\mathcal{I}[\cdot]$ denotes the indicator function.
+Equations (49) and (50) compute the count $n_{m}^{(g)}$ and the mean $s_{m}^{(g)}$ of the representations $u_{ij}^{(g)}$ assigned to the GMM component $U_{m}^{(g)}$, while Equations (51) and (52) perform the momentum updates of $n_{m}^{(g)}$, $s_{m}^{(g)}$, and $U_{m}^{(g)}$.
+These updates are executed per gradient step on the $\ell_{1}$ loss term, ensuring that the GMM evolves in lockstep with DIO.
+IX-C EMA-Based Efficient Estimation of Solution Distribution
+The G-head hierarchical interaction mechanism enables $M$ vectors to support a GMM with $M^{G}$ components; however, as indicated by Equations (43)–(46), it causes a sharp increase in the number of times $\tilde{\text{DIO}}$ must weigh the components, imposing a heavy computational burden.
+To this end, a EMA-Based approach is adopted: cluster the $M^{G}$ components into $M$ centroids, assign weights to these centroids using the procedure in Equation (43), and perform Gumbel-MAX sampling on the $M$ weights as indicated by Equation (46); then iteratively repeat clustering, weight assignment, and sampling within the selected cluster until only $M$ components remain.
+Finally, use Gumbel-MAX to select the component $U_{\hat{m}}$ that can serve as the solution distribution from these $M$ components.
+This reduces the number of weighting operations that $\tilde{\text{DIO}}$ must perform from $M^{G}$ to $M\times G$.
+This recursive scheme demands that $\tilde{\text{DIO}}$ be Lipschitz continuous (i.e., it must assign similar weights to similar components).
+We therefore apply spectral normalization [35] to every parameter matrix in $\tilde{\text{DIO}}$ and apply Lipschitz normalization [36] to the self-attention mechanism within each block, ensuring the required Lipschitz property.
+IX-D Validation of Generative RPM problems
+Notably, the validation of generative RPM has long been hindered by its multi-solution nature: any answer that satisfies the progressive pattern can be correct, rendering traditional point-to-point comparisons ineffective.
+Validation must therefore adopt the same progressive-pattern-driven principle, one that is already embedded in the DIO model trained with the $\ell_{\text{DIEGO}}$ loss, enabling it to perform verification effectively.
 X Experiment
-All experiments conducted in this paper were programmed in Python based on the PyTorch[37] framework. This paper proposes four distinct configurations for RPM problems:
+All experiments in this paper were conducted on four A100-80G GPUs using Python and the PyTorch[37] framework. This paper proposes four configurations for RPM problems:
 X-1 Solo DIO
@@
 X-3 WORLD-augmented DIO
-In the WORLD framework, the hyperparameter $M$ is set to 512, equipping the GMM employed for DIO with 512 components.
-This configuration enables the generation of $M^{N}=512^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options.
-It is worth noting that when WORLD is used to assist DIO in solving generative RPM problems, $M$ is set to 2048;
+In the WORLD framework, when solving discriminative RPM problems, $M$ is set to $2^{11}$, and a random window of size 256 is defined to select the components used for computing the loss term $\ell_{\text{Weighted}}$.
+If the $G$-head hierarchical interaction is activated, $G$ is set to 2, $M$ remains $2^{11}$, and each head employs a window of size 16 to compute the $\ell_{\text{Weighted}}$ loss.
+When the framework is applied to generative RPM tasks, $G$ is set to 2, and $M$ is set to $2^{14}$; the random window for $\ell_{\text{Weighted}}$ inside each head is kept at 16.
 X-4 DIEGO-guided DIO
 In the DIEGO framework, the parameter $F$ is determined by the number of metadata modalities in the targeted RPM problem.
-This paper conducted experiments on the RAVEN [18], I-RAVEN[32] and PGM [19] datasets, centered around these four configurations.
-To ensure fair comparison, we uniformly align the optimizer hyperparameters, dataset scales, and batch sizes for optimizing the aforementioned four configurations with those adopted by the current state-of-the-art RS-Tran model [29].
+This paper conducts experiments on the RAVEN [15], I-RAVEN [28], and PGM [16] datasets, centered around these four configurations.
+To ensure a fair comparison, we use the same optimizer hyperparameters, dataset scale, and batch size as those of the current SOTA RS-Tran model [26].
 X-A Experiment on RAVEN and I-RAVEN Problems
-The performance of our four configurations alongside previous models on the RAVEN and I-RAVEN datasets is summarized in Table I.
+Table I summarizes the performance of our four configurations in comparison with previous models on the RAVEN and I-RAVEN datasets, where “WORLDG” denotes the activation of the $G$-head hierarchical interaction.
 TABLE I: Reasoning Accuracies on RAVEN and I-RAVEN.
 Test Accuracy(%)
-Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
+Model Average Center $\times2$ Grid $\times3$ Grid L-R U-D O-IC O-IG
 CoPINet[18] 52.96/22.84 49.45/24.50 61.55/31.10 52.15/25.35 68.10/20.60 65.40/19.85 39.55/19.00 34.55/19.45
@@
 RS-TRAN[26] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
-Triple-CFN[31] 98.9/99.1 100.0/100.0 99.7/99.8 96.2/97.5 99.8/99.9 99.8/99.9 99.9/99.9 97.0/97.3
-DIO 98.8/99.3 100.0/100.0 99.6/99.7 96.1/97.4 99.8/99.9 99.8/99.9 99.8/99.8 96.7/97.9
-DIO+Brando 99.4/99.5 100.0/100.0 99.8/99.8 98.4/99.0 100.0/100.0 99.9/100.0 100.0/100.0 98.6/98.9
-DIO+WORLD 99.5/99.7 100.0/100.0 99.8/99.8 98.8/99.5 99.9/100.0 100.0/100.0 100.0/100.0 99.0/99.5
+DIO 98.8/99.2 100.0/100.0 99.6/99.7 96.3/97.4 99.8/99.9 99.8/99.9 99.8/99.8 96.7/97.9
+DIO+Brando 99.2/99.5 100.0/100.0 99.8/99.8 97.4/98.6 99.9/99.9 99.9/100.0 100.0/100.0 97.8/98.8
+DIO+WORLD 99.4/99.7 100.0/100.0 99.8/99.8 98.5/99.2 99.9/100.0 100.0/100.0 100.0/100.0 98.4/99.3
+DIO+WORLDG 99.6/99.7 100.0/100.0 99.8/99.8 99.2/99.5 100.0/100.0 100.0/100.0 100.0/100.0 99.3/99.5
 DIO+DIEGO 99.9/99.9 100.0/100.0 99.9/99.9 99.7/99.8 100.0/100.00 100.0/100.0 100.0/100.0 99.8/99.8
-These results indicate that all four model configurations proposed in this study outperform existing baselines on both the RAVEN and I-RAVEN datasets.
-Notably, the DIO+DIEGO configuration achieves a peak accuracy of 99.9%, delivering significant performance gains over prior state-of-the-art models (Triple-CFN[31] and RS-TRAN[29]).
+These results indicate that all four configurations outperform the existing baselines on both the RAVEN and I-RAVEN datasets.
+Notably, the DIO+DIEGO configuration achieves a peak accuracy of 99.9%, delivering significant performance gains over the prior SOTA model (RS-TRAN[26]).
 Particularly in complex reasoning tasks such as $\times3$ Grid and O-IG, the DIO series models demonstrate distinct advantages, validating the generalizability and robustness of their architectural design for abstract visual reasoning challenges.
-More critically, the performance achieved by the DIEGO method has overturned a key experimental conclusion in prior studies[31, 23, 29], which posited that the progressive pattern supervision directly implemented from metadata would inevitably undermine the model’s reasoning capabilities.
-This logically flawed assertion is obsolete.
-Furthermore, our experiments demonstrate that when no additional supervision signals are incorporated, the strength of mutual information inferred by the solving model serves as the primary determinant of its problem-solving performance, offering critical architectural insights for future abstract reasoning systems.
-When progressive pattern descriptions are employed as supervisory signals, the supervision utilization strategy combining causal chain modeling with correction of key chain segments proves highly effective.
+Moreover, our experiments demonstrate that when no additional causal supervisory signals are introduced, the strength of the mutual information inferred by the reasoning model becomes the primary determinant of its problem-solving performance, offering crucial architectural insights for future abstract-reasoning systems.
+When progressive pattern descriptions are employed as supervisory signals, a supervision strategy that combines causal-chain modeling with the correction of key chain segments proves to be highly effective.
+This overturns the long-standing conclusion of prior studies [20, 23, 22, 26] that direct progressive-pattern supervision inevitably undermines model performance.
 X-B Experiment on PGM Problem
-The performance comparison among the four DIO configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset is summarized in Table II.
-TABLE II: Reasoning Accuracies on PGM-neutral.
-Model Test Accuracy(%)
-CoPINet[21] 56.4
-SAVIR-T [25] 91.2
-MRNet [23] 94.5
-RS-CNN[29] 82.8
-RS-TRAN[29] 97.5
-Triple-CFN[31] 97.8
-DIO 98.1
-DIO+Brando 99.0
-DIO+WORLD 99.3
-DIO+DIEGO 99.6
-As shown in Table II, DIO achieves state-of-the-art performance on the PGM-neutral subset of the i.i.d.
-RPM dataset, a benchmark where prior models exhibited relatively low accuracy.
-Notably, the integration of mutual information lower bound tightening methods (Brando and WORLD) yields substantial performance improvements for DIO.
-Furthermore, the DIO+DIEGO attains a peak accuracy of 99.6%, thereby challenging the prevailing hypothesis that metadata supervision necessarily impairs relational reasoning capabilities.
-The paper also conducted experiments on the o.o.d. sub-task of PGM, with the results documented in Table III.
-TABLE III: Accuracies on PGM Generalization Tasks.
-Dataset
-Model/
-Task
-Interpolation Extrapolation
+Column 1 of Table II summarizes the performance comparison among the four DIO configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset.
+TABLE II: Accuracies on PGM Generalization Tasks.
+Model Dataset
+Neutral Interpolation Extrapolation
 Held-out
@@
 Pairs
-MRNet[23]
-68.1 19.2 16.9 30.1 25.9 55.3 38.4
-RS-Tran[29]
-77.2 19.2 12.9 24.7 22.2 43.6 28.4
-Triple-CFN[31]
-80.4 18.4 12.6 25.2 22.0 44.5 29.2
-DIO
-82.0 18.8 13.6 26.5 23.3 45.5 29.9
-DIO+Brando
-83.2 18.2 14.2 26.7 23.6 45.8 30.5
-DIO+WORLD
-87.8 18.9 15.0 27.6 23.8 46.2 30.9
-DIO+DIEGO
-93.2 13.8 14.4 27.2 28.8 98.3 98.3
-The results recorded in Table III reveal that while mutual information tightening and causal chain modeling yield comparable performance gains for DIO on i.i.d.
-RPM tasks, the latter demonstrates significantly superior generalization capabilities under o.o.d.
-RPM conditions.
-To a certain extent, these results suggest that mutual information tightening remains a highly effective approach for i.i.d.
-reasoning tasks, despite mutual information being a statistical measure that does not inherently encode causal relationships crucial for reasoning.
-However, when addressing o.o.d.
-reasoning challenges, adopting causal chain modeling as the methodological cornerstone proves to be a more robust choice.
+CoPINet[18] 56.4 - - - - - - -
+SAVIR-T [22] 91.2 - - - - - - -
+RS-CNN[26] 82.8 - - - - - - -
+MRNet[20] 94.5 68.1 19.2 16.9 30.1 25.9 55.3 38.4
+RS-Tran[26] 97.5 77.2 19.2 12.9 24.7 22.2 43.6 28.4
+DIO 98.1 82.0 18.8 13.6 26.5 23.3 45.5 29.9
+DIO+Brando 99.0 83.2 18.2 14.2 26.7 23.6 46.8 32.8
+DIO+WORLD 99.3 85.4 18.9 15.0 27.6 23.8 55.5 39.0
+DIO+WORLDG 99.5 87.8 18.9 15.0 27.6 23.8 56.2 39.5
+DIO+DIEGO 99.6 93.2 13.8 14.4 27.2 28.8 98.3 98.3
+As shown in this column, DIO attains SOTA performance on PGM-neutral, a benchmark on which earlier models achieved relatively low accuracy.
+Notably, incorporating mutual information lower bound tightening methods (Brando and WORLD) yields substantial improvements for DIO.
+Moreover, DIO+DIEGO reaches a peak accuracy of 99.6%, challenging the prevailing conclusion that direct metadata supervision necessarily impairs relational reasoning.
+Columns 2 onward in Table II record the experimental results on the o.o.d.
+sub-tasks of PGM.
+All results in Table II jointly show that mutual-information tightening and causal-chain modeling yield comparable gains for DIO on i.i.d.
+RPM tasks, while the latter delivers significantly better generalization under o.o.d.
+conditions.
+This suggests that, although mutual information as a statistical measure does not inherently encode the causal relationships vital for reasoning, it remains highly effective for i.i.d.
+tasks; however, when tackling o.o.d.
+challenges, adopting causal-chain modeling as the core methodology is the more robust choice.
+We suggest that this occurs because, under i.i.d.
+conditions, the correlations between contexts and solutions coincide with their causal relationships, allowing mutual-information-driven models to remain effective; in contrast, o.o.d.
+reasoning tasks expose a divergence between correlation and causality, making causal-chain-driven approaches essential.
 X-C Experiment on Generative RAVEN and PGM Problems
-While DIO+WORLD can solve discriminative RPM problems with high precision, it is also capable of solving generative RPM problems.
-Specifically, DIO+WORLD can generate solutions for the context matrices instead of making a selection from a predefined bound option pool.
-This paper proposes that it is reasonable to use a progressive-pattern-driven discriminative model to verify the accuracy of the generative model in solving generative RPM problems.
-Therefore, we employ DIO+DIEGO to validate the generative accuracy of DIO+WORLD.
-Prior to this, we must report DIEGO’s discriminative accuracy for progressive patterns under the supervision of RPM metadata to confirm that DIEGO can effectively verify the solution accuracy for generative RPM problems.
-DIEGO’s reasoning accuracy for progressive patterns on the RAVEN, I-RAVEN, and PGM datasets is recorded in Table IV.
-TABLE IV: DIEGO’s Reasoning Accuracy for Progressive Patterns on the RAVEN, I-RAVEN, and PGM Datasets.
+We argue that a progressive-pattern-driven discriminative model is a reliable way to verify solutions from generative RPM tasks.
+Therefore, we use DIO+DIEGO to evaluate the generation quality of DIO+WORLDG.
+Before running this pipeline, we must report DIEGO’s discriminative performance for progressive patterns under RPM metadata supervision to confirm its reliability as a solution validator.
+Table III lists DIEGO’s reasoning accuracy for progressive patterns on RAVEN, I-RAVEN, and PGM.
+TABLE III: DIEGO’s Reasoning Accuracy for Progressive Patterns on the RAVEN, I-RAVEN, and PGM Datasets.
 Dataset Sub-task Accuracy(%)
@@
 L-R left:99.9 right:99.9
-U-D left:99.9 right:99.9
+U-D up:99.9 down:99.9
 OI-C in:99.9 out:99.9
@@
 L-R left:99.9 right:99.9
-U-D left:99.9 right:99.9
+U-D up:99.9 down:99.9
 OI-C in:99.9 out:99.9
@@
 PGM neutral shape:99.7 line:99.9
-Table IV indicates that the DIO+DIEGO demonstrates remarkably high accuracy in progressive pattern reasoning, making it a suitable validation model for generative RPM tasks.
-Consequently, we evaluate the DIO+WORLD model using DIO+DIEGO and report its accuracy on generative RAVEN, I-RAVEN and PGM-neutral problems in Table V.
-TABLE V: Generation Accuracies of DIO+WORLD
+Table III shows that DIO+DIEGO achieves extremely high accuracy on progressive-pattern reasoning, making it a suitable validator for generative RPM tasks.
+Consequently, we use DIO+DIEGO to evaluate DIO+WORLDG and report its accuracy on generative RAVEN, I-RAVEN, and PGM-neutral problems in Table IV.
+TABLE IV: Generation Accuracies of DIO+WORLDG
 Dataset Sub-task Accuracy (%) Average (%)
-RAVEN/I-RAVEN
+RAVEN/
+I-RAVEN
 Center 98.0/98.2
@@
 PGM Neutral 80.5 80.5
-Table V shows that without extra supervisory signals, DIO+WORLD has attained highly competitive results in open-ended generative RPM problems, strongly confirming the significant research value of the WORLD method.
-Moreover, these results indicate that though mutual information doesn’t directly reveal causal relationships, it’s still highly valuable for building reasoning capabilities and can withstand the tests by progressive-pattern-driven models.
-Notably, the results in Table V still have the potential to improve with further training.
-In the WORLD method, the ability to tighten mutual information emerges in the early training epochs, whereas the generation capability matures only after the loss term $\ell_{\text{Weighted}}$ stabilizes at a low value, necessitating additional training epochs.
-To avoid excessive resource consumption, training was halted once the results adequately demonstrated WORLD’s effectiveness.
-XI Ablation Study
-This section presents a series of ablation experiments focusing on key parameter settings and network architecture.
-XI-A Ablation Study on DIO
-The ablation study on DIO centers on the design of the network’s feed-forward structure.
-Within its progressive-pattern induction module, DIO employs a permutation and combination strategy ($C_{n}^{r}$) to extract progressive patterns from row and column features, aiming to model the causal-chain details inherent in RPM problems.
-For ablation, we remove the ($C_{n}^{r}$) operation and instead feed all six row-column features directly into the progressive-pattern extractor.
-By masking the structure that aligns with the causal chain, we assess the importance of causal-chain modeling in network design.
-The resulting accuracies on key sub-tasks of RAVEN, I-RAVEN and PGM are reported in Table VI.
-TABLE VI: Reasoning Accuracies of masked DIO.
-Dataset Sub-task Accuracy (%)
-RAVEN/I-RAVEN
-$\times3$ Grid 95.6/96.8
-O-IG 95.5/95.6
-PGM Neutral 97.2
-Table VI shows that building RPM-solving models that explicitly incorporate causal chains is crucial for superior reasoning performance.
-Moreover, across all RAVEN and PGM sub-tasks, DIO uniformly splits every image into $4\times 4$ patches.
-This design deliberately bypasses the grid-like entity layouts in RAVEN’s $2\times 2$ Grid and $3\times 3$ Grid sub-tasks, preventing unintended inductive biases and prior injection while embracing the causal chain rationale of RPM problems: multiple entities must be analyzed holistically.
-We therefore conducted ablation experiments on the hyperparameter $N$ to examine how aligning the patch grid with RAVEN’s $2\times 2$ Grid and $3\times 3$ Grid sub-tasks affects performance, and the results are shown in Table VII.
-The results indicate that favoring priors over the causal chain may undermine model performance.
-TABLE VII: The reasoning accuracy of DIO when its patch grid is aligned with entity positions in the “2×2 Grid” and “3×3 Grid” sub-tasks of RAVEN and I-RAVEN.
-Accuracy(%)
-Patch grid 2 $\times$ 2 Grid 3 $\times$ 3 Grid
-Aligned with entity 95.0/98.3 92.8/94.2
-$4\times 4$ Grid 99.6/99.7 96.1/97.4
-XI-B Ablation Study on Brando
-The ablation study on Brando focuses on the hyperparameter $\beta$.
-$\beta$ denotes the number ($\beta-17$) of hypothetical options the Brando network provides to DIO.
-As this paper shows that the count of constructive incorrect options for DIO directly influences the tightness of its derived mutual information variational lower bound, which in turn directly impacts DIO’s reasoning potential.
-Therefore, this paper conducts ablation experiments on the DIO + Brando model with different $\beta$ values across key sub-tasks of RAVEN, I-RAVEN, and PGM, and records the results in the Table VIII.
-TABLE VIII: Accuracies of DIO+Brando under Different $\beta$.
-Dataset Sub-task Accuracy(%)
-$\beta$=24(8) $\beta$=22(6) $\beta$=20(4) $\beta$=18(2) Solo DIO
-RAVEN/
-I-RAVEN
-$\times3$ Grid 98.4 / 99.0 98.0 / 98.5 97.2 / 98.0 96.8 / 97.8 96.1 / 97.4
-O-IG 98.6 / 98.9 98.2 / 98.6 97.6 / 98.3 97.5 / 98.3 96.7 / 97.9
-$\times3$ Grid 97.9 / 98.4 97.6 / 98.0 96.9 / 97.8 96.4 / 97.7 96.1 / 97.4
-O-IG 98.0 / 98.6 97.5 / 98.3 97.2 / 98.2 97.0 / 98.2 96.7 / 97.9
-PGM Neutral 99.0 98.8 98.8 98.5 98.1
-Neutral 98.7 98.4 98.4 98.3 98.1
-Table VIII indicates that even when only two constructive hypothetical options are introduced into DIO, significant performance improvements can be achieved.
-The asterisk (*) on a sub-task indicates that, when the Brando network is applied to this task, the process of learning historical image features defined by Equation (27) is deactivated, facilitating ablation and analysis of this mechanism.
-The asterisk-marked results demonstrate that learning historical image features is effective in mapping out constructive hypothetical options.
-XI-C Ablation Study on WORLD
-The ablation study on WORLD focuses on the hyperparameter $M$.
-$M$ indicates the number of components in the GMM constructed by World for RPM image features.
-The value of $M$ determines the maximum diversity of solution distributions that WORLD can generate for DIO, expressed as $M^{N}$.
-Therefore, ablation experiments centered around the value of $M$ are warranted.
-The reasoning accuracies of DIO+WORLD on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $M$ values are recorded in Table IX.
-TABLE IX: Accuracies of DIO+WORLD under Different $M$.
-Dataset Sub-task Accuracy(%)
-MM=512 MM=256 MM=128 MM=32 Solo DIO
-RAVEN/
-I-RAVEN
-$\times3$ Grid 98.8 / 99.5 98.5 / 99.1 97.9 / 98.5 96.5 / 97.9 96.1 / 97.4
-O-IG 99.0 / 99.5 98.6 / 99.0 98.0 / 98.7 97.1 / 98.0 96.7 / 97.9
-PGM Neutral 99.3 99.0 98.5 98.0 98.1
-The number of components directly affects the complexity of the GMM.
-WORLD proposes an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM.
-As a highly versatile distribution model, the GMM can theoretically ensure WORLD’s broad applicability.
-However, as shown in Table IX, when the number of components is relatively small ($M=32$), WORLD constructs a less complex GMM, which negatively impacts DIO’s performance in the PGM task.
-This paper contends that this phenomenon arises because a 32-component GMM fails to capture the complexity of the underlying distribution of PGM image features.
-If DIO attempts to fit these more complex features into such a simplified GMM, a performance decline can be anticipated.
-Therefore, this study establishes a critical prerequisite for WORLD’s applicability: the GMM modeled by WORLD must be able to adapt to the inherent complexity of the target problem.
-XII Conclusion
-This study has constructed a complete abstract reasoning solution through systematic innovations, with four core innovations forming a progressive technical system that achieves significant progress in both theoretical breakthroughs and empirical performance:
-XII-1 DIO
-We propose that structuring neural network architectures around the causal chains can lead to enhanced reasoning performance.
-We also suggest that the learning objectives of end-to-end reasoning models, including DIO, are essentially variational lower bounds of the mutual information between the contexts and the correct options.
-Consequently, both performance and learning efficacy are directly impacted by the looseness of these lower bounds.
-XII-2 Brando
+The relevant generation results are provided in Figure 11, where, for RAVEN, we display only the most typical $\times3$ Grid and O-IG sub-task outcomes.
+Figure 11: Generated RPM images.
+Table IV shows that, without extra supervisory signals, DIO+WORLDG has attained competitive results in open-ended generative RPM problems, strongly confirming its significant research value.
+X-D Semantic Disentanglement via WORLD Encoding.
+Under the WORLD framework, as illustrated in Figure 7, each RPM image $x_{i}$ is encoded into two latent features: $u_{ij}$ and $b_{ij}\sim\mathcal{N}(0,1)$, and the downstream reasoning operates on $u_{ij}+\epsilon_{ij}$.
+This may give rise to a latent division of labor: $u_{ij}$ encapsulates the core semantics required for reasoning, while the decoder $\mathcal{D}(\cdot)$, trained to maximize the mutual information between $u_{ij}+b_{ij}$ and $x_{i}$, pushes $b_{ij}$ to carry task-irrelevant perturbations.
+To verify this hypothesis, we compare the reconstructed image $\mathcal{D}(u_{ij}+\epsilon_{ij})$ with the original $x_{i}$ in terms of semantics, revealing how random variations in $b_{ij}$ drive semantic drift.
+The comparisons are shown in Figure 12.
+Figure 12: Verification on Image Semantics.
+As can be seen in Figure 12, DIO+WORLDG begins to disentangle perturbing semantics, such as the occlusion relationships among background lines in PGM and the rotation angles of entities in RAVEN, from $u_{ij}$ to some extent.
+XI Conclusion
+This study has constructed a comprehensive abstract reasoning solution through systematic innovations, with four core innovations forming a progressive technical system that achieves significant advancements in both theoretical breakthroughs and empirical performance:
+XI-1 DIO
+We propose that structuring networks around causal chains can enhance reasoning performance.
+We further suggest that the learning objectives of end-to-end reasoning models, including DIO, are essentially variational lower bounds on the mutual information between contexts and correct options.
+Consequently, however elaborate the network, performance hinges on the tightness of this bound.
+This points out that the key to machine intelligence lies in mutual-information tightening or causal-chain modeling.
+XI-2 Brando
 We have integrated the concept of mutual information tightening into abstract reasoning frameworks, showing that training with more constructive incorrect options serves as a key mechanism for bounding mutual information.
@@
 Concurrently, we have developed a constructive hypothetical incorrect option mapping mechanism to analytically constrain the variational lower bound of mutual information.
-XII-3 WORLD
-We constructed a GMM for the distribution of RPM image features, thereby enabling targeted sampling of infinite and semantically rich options.
-This approach not only assists the DIO model in tightening mutual information but also equips it with the capability to solve open-ended generative RPM problems.
-XII-4 DIEGO
-We have introduced external metadata to construct supervision signals, precisely correcting semantic biases in the causal chain modeled by DIO.
-We have established mappings between model reasoning chains and ground-truth causal chains, aligning internal representations with human cognitive frameworks.
-Combining these four innovations yields a complete technical closed loop of “Architecture Design $\rightarrow$ Training Optimization $\rightarrow$ Data Augmentation $\rightarrow$ Causal Rectification”.
+XI-3 WORLD
+We construct a GMM for the distribution of RPM image features, enabling targeted sampling of infinite and semantically rich options.
+This approach not only helps the DIO model tighten mutual information and, to some extent, disentangle perturbing semantics, but also equips it with the capability to solve open-ended generative RPM problems.
+XI-4 DIEGO
+We introduce a non-statistical constraint that precisely corrects semantic biases in the causal chain modeled by DIO, and we establish mappings between the model’s reasoning chains and the ground-truth causal chains, aligning internal representations with human cognitive frameworks.
+This work unites four causal-information innovations that together yield a complete abstract-reasoning stack.